### v1

# maximum a posteriori

Abstract
We introduce BOLT-GAN, a simple yet effective modification of the WGAN framework inspired by the Bayes optimal learning threshold (BOLT). We show that with a Lipschitz continuous discriminator, BOLT-GAN implicitly minimizes a different metric distance than the Earth-Mover (Wasserstein) distance, and also has better training stability. Empirical evaluations on four standard image-generation benchmarks, (CIFAR-10, CelebA-64, LSUN Bedroom-64, and LSUN Church‐64) show that BOLT-GAN consistently outperforms WGAN, achieving 10-60% lower Fréchet Inception Distance (FID). Our results suggest that BOLT is a broadly applicable principle for enhancing GAN training. The software is available in our anonymous repository.
1 Introduction
The Bayes optimal learning threshold (BOLT) (Naeini et al.,, 2025) is a recently proposed framework for discriminative learning. In its basic form, BOLT develops a computational approach to estimate the Bayes error rate, i.e., the lowest classification error achievable by a representation, by sampling an arbitrary learning model. This introduces the BOLT loss, whose minimization drives the model toward the Bayes optimal classifier.
This work adapts the BOLT framework to generative modeling by training the discriminator of a generative adversarial network (GAN) via the BOLT loss. We treat the GAN discriminator as a binary classifier and train it using the BOLT framework. Under this framework, the BOLT objective recovers the well-known Wasserstein GAN (WGAN) (Arjovsky et al.,, 2017) formulation, up to a minor adjustment: the discriminator outputs are constrained to a bounded range. This perspective provides both a Bayesian interpretation of adversarial learning and a pathway to improve training stability. Building on this foundation, we propose the BOLT-GAN, which inherits WGAN-style convergence guarantees while driving discriminator performance toward Bayes optimality.
1.1 Background and Related Work
Bayes Error Estimation
The Bayes error rate is a central concept in statistical learning, characterizing the minimum error of a classification task (Bishop,, 2006). Classical approaches estimate this error rate via distance-based measures such as Bhattacharyya or Mahalanobis distances (Bhattacharyya,, 1943; Mahalanobis,, 1936). These methods often fail in high dimensions (Devroye et al.,, 1996). More recent methods employ nonparametric or model-based estimators, such as $k$-nearest neighbors (Moon et al.,, 2017) or $\epsilon$-ball techniques (Noshad et al.,, 2019). The BOLT framework (Naeini et al.,, 2025) provides a unifying alternative, deriving a bound on Bayes error from classifier outputs and introducing a loss that aligns learning with approximation of the Bayes optimal classifier.
GAN
GAN (Goodfellow et al.,, 2014) consists of a generator, which synthesizes data samples from a latent variable, and a discriminator, which distinguishes between real and generated samples. It is trained via a min-max game between the generator and discriminator. While powerful, the original formulation based on Jensen–Shannon divergence is unstable (Goodfellow et al.,, 2014; Arjovsky et al.,, 2017). Numerous alternatives have been proposed: Wasserstein GAN (WGAN) (Arjovsky et al.,, 2017) leverages the Lipschitz continuity of the discriminator to improve gradient backpropagation. This is implemented by approaches such as WGAN-GP (Gulrajani et al.,, 2017) which enforce the Lipschitz constraint via gradient penalty. Other variants include $f$-GAN (Nowozin et al.,, 2016), which generalizes the objective to any $f$-divergence, LS-GAN (Mao et al.,, 2017), BEGAN (Berthelot et al.,, 2017), and techniques such as spectral normalization (Miyato et al.,, 2018) or progressive growing (Karras et al.,, 2018). These lines of work aim to stabilize training, but do not try to achieve statistical optimality.
Statistical and Bayesian Perspectives
The GAN has been studied from many statistical or Bayesian viewpoints. Bayesian GANs (Saatchi and Wilson,, 2017; Ye and Zhu,, 2018) introduce uncertainty quantification by estimating posterior distribution. Others, such as (Zhang et al.,, 2021; Arora et al.,, 2017), analyze GAN training through the lens of generalization and statistical distances. Our paper adds to this line of study by describing the GAN discriminator explicitly as a model-free Bayes-aligned classifier. In doing so, we provide a Bayesian perspective for adversarial learning and introduce a learning objective function, the bolt loss, for learning a representation that minimize the classification error.
1.2 Main Contributions
The key contributions of this work are as follows: ($i$) We introduce BOLT-GAN, an adversarial framework in which the discriminator is trained with the BOLT loss. The generator is implicitly trained to maximize the Bayes error of discrimination task. The inner maximization identifies the Bayes-aligned discriminator with the largest real-fake gap, and the outer minimization drives the generator to reduce this gap, effectively pushing the Bayes error toward random guessing. ($ii$) We characterize convergence of the BOLT-GAN. We show that, under balanced class priors, the BOLT objective is directly linked to total variation (TV) between the data and generator distributions. This finding shed lights on the numerical instability of the plain BOLT-GAN training algorithm. (iiiiii) To address this instability issue, we propose Lipschitz BOLT-GAN, which enforces a 1-Lipschitz constraint on the discriminator. We show that under this constraint, the learning objective is bounded above Wasserstein distance. This indicates that this framework guides the generator toward the data distribution using a weaker notion of distance. This leads to stabilization of the gradient backpropagation. ($iv$) We validate our results empirically. On CIFAR-10, CelebA-64, and LSUN, Lipschitz BOLT-GAN showing significant improvements.
Notation
Throughout the paper, we adapt the following notation: expectations of of a random variable $X$ with respect to a distribution $P$ is denoted by $\mathbb{E}_{X\sim P}[\cdot]$. The distribution $P=P_{C}$ is defined as the conditional data distribution given the class C. The set $\left\{1,\ldots,n\right\}$ is abbreviated as $[n]$.
2 Background and Preliminaries
We restrict our analysis to a binary classification task with classes $C_{1}$ and $C_{2}$, although extensions to multi-class problems are possible. Consider a binary classification task with classes $C_{1}$ and $C_{2}$ occurring with prior probabilities $q_{1}$ and $q_{2}=1-q_{1}$, respectively. The data samples are drawn from the space $\mathcal{X}\subseteq\mathbb{R}^{d}$, for some integer $d$, with class-conditional distributions $P_{1}$ and $P_{2}$ for the classes $C_{1}$ and $C_{2}$, respectively. For this setting, the Bayes error rate (BER), denoted by $\varepsilon_{\mathrm{Bayes}}$, is defined as the minimum error probability achievable by a measurable classifier $f:\mathcal{X}\mapsto\{C_{1},C_{2}\}$. The BER is determined from the data distribution as (Bishop,, 2006)
$\varepsilon_{\mathrm{Bayes}}$ $=1-\int_{\mathcal{X}}\max_{\lambda}\left[q_{\lambda}P_{\lambda}(x)\right]\,\mathrm{d}x.$ (1)
and is achieved by the maximum a posteriori (MAP) classifier, which assigns a sample $x$ to the class $\lambda\in\{1,2\}$ that maximizes the joint distribution $q_{\lambda}P_{\lambda}(x)$.
2.1 BOLT Loss for Classification
An explicit computation of the BER(1) requires full knowledge of the data distribution, i.e.,  both the class prior $q_{\lambda}$ and the class conditional feature distribution $P_{\lambda}$ , As these are unavailable, the BER must be bounded and/or estimated ; see (Devroye et al.,, 1996; Fukunaga,, 2013; Noshad et al.,, 2019) for some sample studies. The recent study by Naeini et al., (2025) develops a new framework to estimate an upper bound on the Bayes error. The bound is universal in the sense that it can be computed by sampling from an arbitrary measurable and bounded function, e.g., a neural network. It is further tight, as it can achieve the Bayes error when optimized over a set of measurable functions. Motivated by these observations, Naeini et al., (2025) introduces the BOLT loss, which utilizes the proposed bound to approximate the optimal Bayes classifier via a parameterized learning model.
Theorem 1 (Binary Bayes error (Naeini et al.,, 2025)).
Let $h:\mathcal{X}\to[-1,0]$ be a measurable function. Then, the Bayes error of the binary classification satisfies
$\varepsilon_{\mathrm{Bayes}}\leq q_{2}+q_{2}\mathbb{E}_{X\sim P_{2}}\left[h(X)\right]-\hskip-1.42262ptq_{1}\mathbb{E}_{X\sim P_{1}}\left[h(X)\right].$ (2)
Naeini et al., (2025) show that the bound (2) is tight when optimized over $h$. This enables a computational means to estimate $\varepsilon_{\mathrm{Bayes}}$ by sampling from output of classifier. In particular if $h$ is a parameterized predictor model $h_{\theta}(\cdot)$, e.g., a neural network with weights $\theta$, the bound can be estimated and minimized. This transforms the upper bound in Theorem 1 into a training objective referred to as the Bayes optimal learning threshold (BOLT). While Naeini et al., (2025) establish the BOLT loss for a general multi-class problem, in this work we focus on the binary case, relevant to the discrimination component of the GANs.
Training via BOLT
Consider a training dataset $\mathcal{D}=\left\{(x_{i},\lambda_{i}):i\in[n]\right\}$, where $\lambda_{i}\in\{1,2\}$ denotes the class index from which $x_{i}$ is drawn. Assume that the samples in $\mathcal{D}$ are drawn i.i.d.  from the data distribution. BOLT trains a parameterized model $h_{\theta}:\mathcal{X}\rightarrow[-1,0]$ on the dataset $\mathcal{D}$ to estimate the Bayes error rate by optimizing BER empirical loss over $\theta$.
Remark 1.
Note that the output of the model $h_{\theta}$ is restricted to be within the interval [1,0][-1,0]. This is not a significant restriction, as it corresponds to a simple negation of the softmax output of the classifier.
For a given sample $(x,\lambda)$$\in\mathcal{D}$, the BOLT loss is defined as
$\ell_{\rm{BOLT}}(h_{\theta}(x),\lambda)=(-1)^{\lambda}h_{\theta}(x).$ (3)
To train the model $h_{\theta}$ on $\mathcal{D}$, where the empirical risk
$\hat{R}(\theta)=\frac{1}{n}\sum_{i\in[n]}\ell_{\text{BOLT}}(h_{\theta}(x_{i}),\lambda_{i}),$ (4)
This minimization can be done by mini-batch stochastic gradient descent or other standard training method.
The connection, between the BOLT-trained model and the Bayes error rate, can be understood using the results of Naeini et al., (2025): from Theorem 1, we can write
$\varepsilon_{\rm{Bayes}}\leq q_{2}+\mathbb{E}\left[\ell_{\rm{BOLT}}(h_{\theta}(X),\lambda)\right],$ (5)
where the expectation is taken with respect to the data distribution. Note that the prior $q_{2}$ is efficiently estimated from the data. By universal approximation (Cybenko,, 1989) and the tightness result of Naeini et al., (2025), one can conclude that the trained model in this case efficiently estimates the Bayes error.
2.2 Generative Adversarial Networks
A GAN consists of two learning models: a generator $g_{\phi}$ that generates data samples from the latent variable [^1] Typically, $Z$ is drawn from a multivariate Gaussian distribution. $Z\sim P_{Z}$ to generate samples of an estimator of $\mathcal{X}$, and a discriminator $h_{\theta}$ that distinguishes the generated sample $g_{\phi}(Z)$ from the true data sample $X\sim P_{\rm data}$. To keep the notation tractable, we denote the induced distribution, i.e., distribution of $g_{\phi}(Z)$, by $P_{\rm g}$, which is determined by $P_{Z}$ and the generator model $g_{\phi}$.
From a learning perspective, the discriminator solves a binary classification problem: it classifies the input as $C=$ real or $C=$ fake, corresponding to $\lambda=1$ or $\lambda=2$. The conventional approach is to train $h_{\theta}$ such that it estimates the posterior class probability. This leads to the vanilla GAN design proposed by Goodfellow et al., (2014). An alternative approach is to use the Kantorovich-Rubinstein duality (Villani et al.,, 2009) to implicitly minimize the Wasserstein distance between the data and induced distributions. This approach leads to the Wasserstein GAN design (Arjovsky et al.,, 2017). We briefly explain these two approaches in the sequel.
Vanilla GAN
The vanilla training approach, proposed in the initial study by Goodfellow et al., (2014), treats the discrimination as a classic binary classification problem. In this sense, it sets $h_{\theta}(x)=P(C=\texttt{real}|x)$ and uses the cross-entropy loss to train it. [^2] In practice, $h_{\theta}$ is a network with sigmoid-activated output. This approach implicitly implements the maximum likelihood training of the generator (Arjovsky et al.,, 2017). The risk function, i.e., the expected loss used for training, in this case is given by
$R_{\rm ML}(\phi,\theta)=\mathbb{E}\left[\mathrm{CE}(Y,\lambda)\right],$ (6)
where $\mathrm{CE}$ denotes the cross-entropy loss function, $Y$ is the output of the discriminator, i.e., $Y=h_{\theta}(X)$, for an input $X\in\mathcal{X}$ to the discriminator, and $\lambda\in\left\{1,2\right\}$ is the label of $X$ with $\lambda=1$ and $\lambda=2$ denoting real and fake inputs respectively. Where $\lambda$ is sampled from a prior $\pi=\Pr\left\{\lambda=1\right\}$ for some $\pi\in(0,1)$, [^3] Typically, $\pi=0.5$ is considered. and $X$ is sampled conditionally from $P_{1}=P_{\rm data}$ when $\lambda=1$ and $P_{2}=P_{\rm g}$ when $\lambda=2$.
Noting that $P_{2}=P_{\rm g}$ is the distribution of $g_{\phi}(Z)$ computed from the latent sample $Z\sim P_{Z}$, we can use the definition of cross-entropy to expand the risk (6) as
$R_{\rm ML}(\phi,\theta)=$ $-\pi\mathbb{E}_{X\sim P_{\rm{data}}}\left[\log h_{\theta}(X)\right]$ (7)
$-(1-\pi)\mathbb{E}_{Z\sim P_{Z}}\left[\log(1-h_{\theta}(g_{\phi}(Z)))\right].$
Defining $\mathcal{L}_{\rm ML}(\phi,\theta)=-R_{\rm ML}(\phi,\theta)$, the generator and discriminator are jointly trained by the following min-max game
$\min_{\phi}\max_{\theta}\mathcal{L}_{\rm ML}(\phi,\theta),$ (8)
which is equivalent to the worst-case risk minimization, i.e., fooling the best discriminator. It is shown that the solution to this min-max game, when solved for any measurable generator $g$ and discriminator $h$, guarantees the convergence of $P_{\rm g}$ to $P_{\rm data}$ as measured by the convergence of the Jensen–Shannon (JS) divergence to zero(Goodfellow et al.,, 2014, Theorem 1).
Wasserstein GAN
Arjovsky et al., (2017) propose to train the generator-discriminator pair by minimizing the Wasserstein-$1$ distance, also known as Earth Mover (EM) distance, between $P_{\rm{data}}$ and $P_{\rm g}$. Using the Kantorovich-Rubinstein duality, it is shown that the EM distance minimization is equivalent to solving a min-max problem with the following objective (Arjovsky et al.,, 2017, Theorem 3)
$\mathcal{L}_{\rm EM}(\phi,\theta)=$ $\;\mathbb{E}_{X\sim P_{\rm{data}}}[h_{\theta}(X)]$
$-\mathbb{E}_{Z\sim P_{Z}}[h_{\theta}(g_{\phi}(Z))].$ (9)
The training is carried out by solving the same min-max game as in (8) with $\mathcal{L}_{\rm EM}(\phi,\theta)$ as the objective. While the WGAN still guarantees the convergence of $P_{\rm g}$ to $P_{\rm data}$, the notion of convergence is weaker than vanilla GAN, in the sense that it only minimizes the EM distance between the two distributions instead of an information divergence. Arjovsky et al., (2017) show that the WGAN has better training stability, as compared with the vanilla GAN approach.
Remark 2.
Note that in the WGAN, the discriminator model $h_{\theta}$ is not restricted to compute probabilities. Intuitively, The discriminator in this case is a real-valued critic rather than a probabilistic classifier. The model $h_{\theta}$ is however constrained to be 1-Lipschitz. In practice, this constraint can be enforced by approaches such as weight clipping, gradient penalty, or spectral normalization. See the supplementary file for additional discussion.
3 BOLT-GAN Framework
Here we develop the proposed BOLT-GAN framework for training a GAN model. This scheme invokes the BOLT loss to train the generator to fool the Bayes optimal discriminator. The convergence in BOLT-GAN is in the sense of convergence in total variation, i.e., stronger than vanilla GAN, for a Lipschitz discriminator it converges in a weaker sense than WGAN. This implies a trade-off between the strength of convergence and stability.
3.1 MAP Estimator via BOLT
To understand the BOLT-GAN framework, we first present a result which relates a BOLT-trained model to the Bayesian optimal classifier: consider $h_{\theta}:\mathcal{X}\rightarrow[-1,0]$ that is trained for a binary classification task using the BOLT loss. Theorem 2 shows how an estimator of the Bayesian optimal classifier, i.e., the MAP estimator, is computed from $h_{\theta}$.
Theorem 2 (BOLT Estimator of MAP Classifier).
Let $h^{\star}$ be the minimizer of the upper bound in Theorem 1. Assume that the model $h_{\theta}$ approximates $h^{\star}$ in the sense that $\|h^{\star}(x)-h_{\theta}(x)\|_{\infty}\leq\epsilon$. Then, the plug-in classifier
$\widehat{C}_{\theta}(x)=\begin{cases}C_{1}&h_{\theta}(x)\geq-0.5,\\[4.0pt] C_{2}&h_{\theta}(x)<-0.5.\end{cases}$ (10)
is a plug-in estimator of the MAP classifier, i.e.
$\widehat{C}(x)=\arg\max_{\lambda\in\{1,2\}}q_{\lambda}P_{\lambda}(x)$ (11)
Proof.
The proof is given in the supplementary file. ∎
Remark 3 (BOLT Bias and Variance).
Let $U(x)$ denote the likelihood ratio $U(x)=P_{1}(x)/P_{2}(x)$ and let $\widehat{U}$ be its plug-in estimator (see the supplementary file for construction details). Assume $\mathbb{E}[\,|\widehat{U}(X)-U(X)|\,]\leq\varepsilon_{0}$. Then the BOLT plug-in Bayes-error estimator satisfies
$\bigl|\mathbb{E}\big[\widehat{\varepsilon}_{\mathrm{BOLT}}\big]-\varepsilon_{\mathrm{Bayes}}\bigr|$ $\;\leq\;q_{1}\,\varepsilon_{0}+\mathcal{O}\!\big(M^{-1/2}\big),$ (12)
$\mathrm{Var}\!\big(\widehat{\varepsilon}_{\mathrm{BOLT}}\big)$ $\;=\;\mathcal{O}\!\big(M^{-1}\big),$ (13)
where $M$ is the number of samples used in the plug-in estimator. Additional details are given in the supplementary file.
Theorem 2 gives an intuitive interpretation to the BOLT loss. It indicates that using BOLT, the model learns to mimic the Bayesian optimal classifier. From the computational point of view, using BOLT leads to gradients that guide the model to approximate the MAP estimator. This finding enables us to use the BOLT loss to train the discriminator of a GAN model. We illustrate this approach in the sequel.
BOLT-GAN
Let $s_{\theta}:\mathcal{X}\to\mathbb{R}$ denote a model which computes a raw score for input $x$, and let $h_{\theta}$ be a negative sigmoid activation of its output, i.e.
$h_{\theta}(x)=-\sigma(s_{\theta}(x)).$ (14)
From Theorem 2, one can conclude that in this case, the model $h_{\theta}$ estimates the Bayesian optimal discriminator.
Following a similar approach as in vanilla GAN, we can consider sampling from data and generator distributions conditionally with with $P_{1}=P_{\rm data}$ and $P_{2}=P_{\rm g}$ using a prior $\pi=P(\lambda=1)$ for some $\pi\in(0,1)$, where $\lambda=1$ and $\lambda=2$ correspond to real and fake samples, respectively. Noting that the generator samples are given by $g_{\phi}(Z)$ for latent samples $Z\sim P_{Z}$, the expected BOLT loss of the discriminator is given by
$R_{\text{BOLT}}(\phi,\theta)=-\mathcal{L}_{\text{BOLT}}(\phi,\theta),$ (15)
with $\mathcal{L}_{\text{BOLT}}(\phi,\theta)$ being
$\mathcal{L}_{\text{BOLT}}(\phi,\theta)=\,$ $\pi\mathbb{E}_{X\sim P_{\text{data}}}[h_{\theta}(X)]$
$(1-\pi)\mathbb{E}_{Z\sim P_{Z}}[h_{\theta}(g_{\phi}(Z))].$ (16)
Given this risk function, the generator and discriminator are jointly trained by maximizing the minimum loss. This is equivalent to solving the following min-max game:
$\min_{\phi}\max_{\theta}\mathcal{L}_{\text{BOLT}}(\phi,\theta).$ (17)
Considering Theorem 1, one can see that the inner optimization minimizes the Bayes error upper bound. [^4] Note that $\mathcal{L}_{\text{BOLT}}=-R_{\text{BOLT}}$.
Stability of BOLT-GAN
The learning objective in BOLT-GAN is similar to that of WGAN, with two main differences: ($i$) BOLT-GAN considers an arbitrary sampling prior $\pi$, as in vanilla GAN, and ($ii$) it does not impose any Lipschitz continuity restriction on the discriminator model. The discussions in (Arjovsky et al.,, 2017; Gulrajani et al.,, 2017; Miyato et al.,, 2018) suggest that the latter difference can lead to lack of stability, when we use BOLT-GAN in this vanilla form. Our numerical experiments confirm the validity of this conjecture. To understand the reason behind this behavior, we study the convergence of the vanilla BOLT-GAN in the sequel. Our result indicates that BOLT-GAN implicitly minimizes the total variation (TV), which is a strict notion of distance. This finding enables us to modify the BOLT-GAN learning objective to a stable form.
3.2 Convergence of BOLT-GAN
Consider the general learning objective
$\mathcal{L}_{\text{BG}}^{(\pi)}(g,h)=\,$ $\pi\mathbb{E}_{X\sim P_{\text{data}}}[h(X)]$ (18)
$-(1-\pi)\mathbb{E}_{Z\sim P_{Z}}[h(g(Z))],$
which is defined for any measurable $h:\mathcal{X}\mapsto[0,1]$ and measurable $g:\mathcal{Z}\mapsto\mathcal{X}$ that maps that latent space $\mathcal{Z}$ to the data space. $\mathcal{L}_{\text{BG}}^{(\pi)}$ reduces to $\mathcal{L}_{\text{BOLT}}$, when we replace $g$ and $h$ are parameterized models $g_{\theta}$ and $h_{\theta}$.
Our goal is to understand the metric, under which the generator distribution $P_{\rm g}$ converges to $P_{\rm data}$, when the min-max game with objective $\mathcal{L}_{\text{BG}}^{(\pi)}$ is solved. To this end, let us first define total variation (TV).
Definition 1 (Total variation norm).
The total variation between two distributions $P$ and $Q$ defined on a measurable $\mathcal{X}$ is defined as
$TV(P,Q)=\sup_{A\subseteq\mathcal{X}}|P(A)-Q(A)|,$ (19)
with the supremum taken over all events $A$ in the $\sigma$-field over $\mathcal{X}$.
The convergence of TV to zero implies for convergence of other distances, such as including JS and EM. The following lemma gives an alternative representation for TV.
Lemma 1 (Dudley, 2002a ).
The total variation between $P$ and $Q$ is equivalently given by
$TV(P,Q)=\sup_{f:\mathcal{X}\to[0,1]}\int f\,\mathrm{d}(P-Q),$ (20)
where the difference PQP-Q is a signed measure on $\mathcal{X}$.
The following theorem characterizes the connection between the BOLT-GAN learning objective $\mathcal{L}_{\text{BG}}^{(\pi)}$ and the TV between the generator and the data distributions.
Theorem 3 (BOLT vs TV).
Fix the prior $\pi\in(0,1)$. For generator $g$, let $P_{\rm g}$ denote the distribution of $g(Z)$, when $Z\sim P_{Z}$. Define $D^{(\pi)}$ as
$D^{(\pi)}(g)$ $\triangleq\max_{h}\mathcal{L}_{\text{BG}}^{(\pi)}(g,h),$ (21)
with the maximum taken over all discriminators, i.e., all $h:\mathcal{X}\mapsto[0,1]$. Then, $D^{(\pi)}$ satisfies the following property
$D^{(\pi)}(g)+D^{(1-\pi)}(g)$ $\;\geq\;\operatorname{TV}(P_{\mathrm{data}},P_{\rm g}).$ (22)
with equality when $\pi=0.5$.
Proof.
From the definition (18), we have
$\mathcal{L}_{\text{BG}}^{(\pi)}(g,h)+\mathcal{L}_{\text{BG}}^{(1-\pi)}(g,h)=$ $\mathbb{E}_{X\sim P_{\text{data}}}[h(X)]$
$-\mathbb{E}_{Z\sim P_{Z}}[h(g(Z))],$ (23)
which noting that $g(Z)$ is distributed as $P_{\rm g}$ can be written as
$\mathcal{L}_{\text{BG}}^{(\pi)}(g,h)+\mathcal{L}_{\text{BG}}^{(1-\pi)}(g,h)=$ $\mathbb{E}_{X\sim P_{\text{data}}}[h(X)]$
$-\mathbb{E}_{X\sim P_{\rm g}}[h(X)].$ (24)
Setting the signed metric $P_{\text{data}}-P_{\rm g}$ to be the difference of densities, we can conclude that
$\Sigma(g)$ $\triangleq\max_{h:\mathcal{X}\mapsto[0,1]}\left[\mathcal{L}_{\text{BG}}^{(\pi)}(g,h)+\mathcal{L}_{\text{BG}}^{(1-\pi)}(g,h)\right]$ (25)
$=\max_{h:\mathcal{X}\mapsto[0,1]}\left[\int h(X)\mathrm{d}(P_{\text{data}}-P_{\rm g})\right]$ (26)
$\stackrel{{\scriptstyle\text{(a)}}}{{=}}\mathrm{TV}(P_{\rm data},P_{\rm g}),$ (27)
where (a) follows from Lemma 1. We next use the fact that
$\Sigma(g)\leq\max_{h}\left[\mathcal{L}_{\text{BG}}^{(\pi)}(g,h)\right]+\max_{h}\left[\mathcal{L}_{\text{BG}}^{(1-\pi)}(g,h)\right],$ (28)
and using the definition of $D^{(\pi)}$ to conclude that
$\mathrm{TV}(P_{\rm data},P_{\rm g})$ $=\Sigma(g)$ (29)
$\leq D^{(\pi)}(g)+D^{(1-\pi)}(g)$ (30)
which proves the inequality. We further note that with $\pi=0.5$
$\mathrm{TV}(P_{\rm data},P_{\rm g})$ $=\Sigma(g)$
$=\max_{h}\left[\mathcal{L}_{\text{BG}}^{(0.5)}(g,h)+\mathcal{L}_{\text{BG}}^{(0.5)}(g,h)\right]$
$=2D^{(0.5)}(g)$ (31)
which indicates that equality is achieved with $\pi=0.5$. ∎
The above result indicates that using BOLT-GAN, the generator distribution converges to the data distribution at least as strong as minimizing the TV between the two distributions.
Corollary 1 (Half–TV guarantee).
Let $D^{(\pi)}$ be as in Theorem 3. Assume for a given generator $g$, we have
$\max\left\{D^{(\pi)}(g),D^{(1-\pi)}(g)\right\}\leq\varepsilon$ (32)
Then, the TV between the generator and data distributions is bounded as
$\operatorname{TV}\!\bigl(P_{\mathrm{data}},P_{\rm g}\bigr)\leq 2\varepsilon.$ (33)
Proof.
Using Theorem 3, we have
$2\max\left\{D^{(\pi)}(g),D^{(1-\pi)}(g)\right\}$ $\geq D^{(\pi)}(g)+D^{(1-\pi)}(g)$ (34)
$\;\geq\;\operatorname{TV}(P_{\mathrm{data}},P_{\rm g}).$ (35)
We can hence write
$\varepsilon$ $\geq\max\left\{D^{(\pi)}(g),D^{(1-\pi)}(g)\right\}$ (36)
$\;\geq\;\frac{1}{2}\operatorname{TV}\!\bigl(P_{\mathrm{data}},P_{\rm g}\bigr).$ (37)
which concludes the proof. ∎
The above result suggests that the BOLT-GAN objective encourages the generated distribution $P_{g}$ to approach the true data distribution $P_{\mathrm{data}}$ under the total variation (TV) metric.
3.3 Regularized BOLT-GAN via Lipschitz
Lipschitz critics are known to stabilize adversarial training and provide informative, non-saturating gradients Arjovsky et al., (2017); Gulrajani et al., (2017); Miyato et al., (2018). Motivated by this, we enforce $1$-Lipschitz constraint on the discriminator. Intuitively, this can lead to better gradients while training the generator. This can reduce gradient explosion or vanishing, as well as the variance of stochastic updates flowing through $\nabla_{\phi}\,h_{\theta}(g_{\phi}(z))$. In this section, we show that under this modification, the BOLT-GAN implies a weaker notion of convergence and better stability.
Lipschitz BOLT-GAN
The Lipschitz BOLT-GAN trains the generator and discriminator by solving the min-max game (17) , subject to Lipschitz continuity of the discriminator. More precisely, it solves
$\min_{\phi}\max_{h\in\mathcal{H}_{\mathrm{Lip}}}\mathcal{L}_{\text{BOLT}}(\phi,\theta),$ (38)
where $\mathcal{H}_{\mathrm{Lip}}$ is the set of functions $h_{\theta}:\mathcal{X}\mapsto[-1,0]$ and
$\mathcal{H}_{\mathrm{Lip}}=\{h:\mathcal{X}\mapsto[0,1]:\,$ $|h(x)-h(y)|\leq\|x-y\|$ (39)
$\forall\,x,y\in\mathcal{X}\},$
The Lipschitz continuity further imposes a smoothness regularization on the BOLT-GAN discriminator.
For simplicity, we drop the subscript $\theta$ in $h_{\theta}$, stating the following theorem. 4.
Theorem 4 (Convergence of Lipchitz BOLT-GAN).
Fix $0<\pi\leq 0.5$, and define the bounded 1-Lipschitz class of discriminator models as
$\mathcal{H}_{\mathrm{Lip}}=\{h:\mathcal{X}\mapsto[0,1]:\,$ $|h(x)-h(y)|\leq\|x-y\|$ (40)
$\forall\,x,y\in\mathcal{X}\},$
for some norm $\|\cdot\|$ defined on data space $\mathcal{X}$. For the generator model $g$, define
$D^{(\pi)}_{\mathrm{Lip}}(g)\triangleq\max_{h\in\mathcal{H}_{\mathrm{Lip}}}L^{(\pi)}_{\mathrm{BG}}(g,h),$ (41)
with $L^{(\pi)}_{\mathrm{BG}}(g,h)$ being defined in (18). Then, for any prior $0<\pi\leq 0.5$, we have
$D^{(\pi)}_{\mathrm{Lip}}(g)\leq W_{1}\big(P_{\mathrm{data}},P_{\rm g}\big).$ (42)
where $W1$ is the Lipchitz Wasserstein‑1 measure defined in (46) below.
Proof.
Starting from (3.2) in the proof of Theorem 3, we can write
$\Sigma_{\mathrm{Lip}}(g)\triangleq\max_{h\in\mathcal{H}_{\mathrm{Lip}}}\left[\mathcal{L}_{\text{BG}}^{(\pi)}(g,h)+\mathcal{L}_{\text{BG}}^{(1-\pi)}(g,h)\right]$ (43)
$=\max_{h\in\mathcal{H}_{\mathrm{Lip}}}\Bigl\{\mathbb{E}_{X\sim P_{\mathrm{data}}}\!\big[h(X)\big]-\mathbb{E}_{X\sim P_{\rm g}}\!\big[h(X)\big]\Bigr\}.$ (44)
In the supplementary file, we show that for $0<\pi\leq 0.5$,
$D^{(\pi)}_{\mathrm{Lip}}(g)\leq\max_{h\in\mathcal{H}_{\mathrm{Lip}}}\left[\mathcal{L}_{\text{BG}}^{(\pi)}(g,h)+\mathcal{L}_{\text{BG}}^{(1-\pi)}(g,h)\right]$ (45)
and therefore
$D^{(\pi)}_{\mathrm{Lip}}(g)\leq\max_{h\in\mathcal{H}_{\mathrm{Lip}}}\Bigl\{\mathbb{E}_{X\sim P_{\mathrm{data}}}\!\big[h(X)\big]-\mathbb{E}_{X\sim P_{\rm g}}\!\big[h(X)\big]\Bigr\}.$
The Kantorovich-Rubinstein duality (Villani et al.,, 2009) gives
$W_{1}(P_{\mathrm{data}},P_{g})=$ (46)
$\sup_{h\in\mathcal{H}_{\mathrm{Lip}}^{+}}\Bigl\{\mathbb{E}_{X\sim P_{\mathrm{data}}}\!\big[h(X)\big]-\mathbb{E}_{X\sim P_{\rm g}}\!\big[h(X)\big]\Bigr\},$
where $\mathcal{H}_{\mathrm{Lip}}^{+}$ denotes the set of all 1-Lipschitz functions on $\mathcal{X}$
$\mathcal{H}_{\mathrm{Lip}}^{+}=\{h:|h(x)-h(y)|\leq\|x-y\|\hskip 2.84526pt\forall\,x,y\in\mathcal{X}\}.$ (47)
Noting that $\mathcal{H}_{\mathrm{Lip}}\subset\mathcal{H}_{\mathrm{Lip}}^{+}$, we can write
$W_{1}(P_{\mathrm{data}},P_{\rm g})$
$\geq\sup_{h\in\mathcal{H}_{\mathrm{Lip}}}\Bigl\{\mathbb{E}_{X\sim P_{\mathrm{data}}}\!\big[h(X)\big]-\mathbb{E}_{X\sim P_{\rm g}}\!\big[h(X)\big]\Bigr\},$ (48)
$\geq D^{(\pi)}_{\mathrm{Lip}}(g).$ (49)
This concludes the proof. ∎
The above result indicates an interesting dichotomy. While the WGAN implicitly minimizes the Earth-Mover (Wasserstein-1) distance between $P_{\mathrm{data}}$ and $P_{g}$, the BOLT-GAN objective can be viewed as implicitly minimizing the total variation (TV) distance. The Lipschitz BOLT-GAN, on the other hand, bridges these two perspectives by enforcing Lipschitz continuity, thereby promoting a smoother optimization landscape. This intuition may be suggestive of the relative advantages of these three methods.
Lipschitz enforcement. We impose Lipschitzness on $h_{\theta}$ by adding a gradient penalty (GP, $\lambda_{\mathrm{GP}}{=}10$) (Gulrajani et al.,, 2017): for $\hat{x}=\epsilon x+(1-\epsilon)g_{\phi}(z)$ with $\epsilon\!\sim\!\mathcal{U}(0,1)$, we add $\lambda_{\mathrm{GP}}\,\mathbb{E}_{\hat{x}}\!\big(\|\nabla_{\hat{x}}h_{\theta}(\hat{x})\|_{2}-1\big)^{2}$ to the loss. Alternatives (spectral normalization, weight clipping) and implementation details are discussed in the supplement.
4 Experiments and Results
To establish the benefits of the lipchitz BOLT-GAN design, we conduct experiments on standard image generation tasks using the MNIST (LeCun et al.,, 1998) dataset. and CelebA (Liu et al.,, 2015). We use the Fréchet inception distance (FID) (Heusel et al.,, 2017) to assess the quality of the generated images over training epochs and provide comparisons to the WGAN.
4.1 Experimental Setup
We adopt a residual DCGAN backbone (Gulrajani et al.,, 2017): the generator upsamples a set of 128-D Gaussian latent images to $64{\times}64$ RGB via residual blocks; the critic symmetrically downsamples these images with global-sum pooling and a final linear head for classification. BatchNorm/ReLU activation is used in the generator. LeakyReLU in the critic, with a Lipschitz constraint enforced through a gradient penalty ($\lambda_{GD}=10$). We evaluate performance on CIFAR-10, CelebA-64, and LSUN Bedroom/Church-64. For training we use Adam ($\alpha=2{\times}10^{-4},\beta_{1}=0.5,\beta_{2}=0.999$), with batch size 64, for 20 epochs. FID (Heusel et al.,, 2017) is computed every 5 epochs on 10,000 samples using torch-fidelity (Seitzer,, 2020), averaged over three seeds. Additional architectural and hardware details appear in the supplementary.
4.2 Quantitative and Convergence Results
Table 3 shows that BOLT–GAN consistently surpasses WGAN, achieving 10–60% lower FID across all benchmarks. As illustrated in Figure 1, it also converges faster and exhibits smoother training dynamics with reduced variance and fewer oscillations. Complementary qualitative results in Figure 2 highlight sharper textures and fewer artifacts in samples generated by BOLT–GAN after 100 epochs on CIFAR-10. We also trained BOLT–GAN without Lipschitz regularization. This TV-based variant quickly diverges (FID >>300–400), confirming that the gradient-penalty term is essential for stability.
Table 1: FID (↓) after 20 epochs for the non-Lipschitz BOLT–GAN ($\lambda_{\mathrm{GP}}=0$) compared with the Lipschitz version.
Dataset Non-Lipschitz ($\lambda_{\mathrm{GP}}=0$) Lipschitz ($\lambda_{\mathrm{GP}}=10$)
CIFAR-10 $\approx$350 (unstable) 44.2 ± 1.2
CelebA-64 $\approx$300 9.2 ± 0.4
LSUN Bedroom-64 Diverged (>>400) 40.3 ± 2.3
4.3 Ablation Studies
[…]
Table 2: FID (↓) of WGAN-GP and BOLT-GAN-GP on CIFAR-10 across different values of $\lambda_{\mathrm{GP}}$.
Method $\lambda_{\mathrm{GP}}=1$ 5 10 20
WGAN-GP 60.8 59.6 60.0 61.1
BOLT-GAN-GP 44.3 43.6 44.2 46.7
Table 3: FID ($\downarrow$) after 20 epochs.
Dataset WGAN BOLT GAN Relative $\Delta$ (%)
CIFAR-10 $60.0\!\pm\!1.8$ $44.2\!\pm\!1.2$ 26.3-26.3
CelebA-64 $10.3\!\pm\!0.5$ $9.2\!\pm\!0.4$ 10.7-10.7
LSUN Bedroom-64 $102.5\!\pm\!2.9$ $40.3\!\pm\!2.3$ 60.7-60.7
LSUN Church-64 $43.5\!\pm\!1.4$ $14.8\!\pm\!0.6$ 65.9-65.9
Table 4: Best FID ($\downarrow$) after 20 epochs for different $\lambda_{\text{GP}}$. Bold marks the default used in all other experiments.
Dataset $\lambda_{\text{GP}}$
1 5 10 20
[…]
CelebA-64 $10.5\!\pm\!0.5$ $9.4\!\pm\!0.4$ $\mathbf{9.2\!\pm\!0.4}$ $10.0\!\pm\!0.5$
Table 5: CIFAR‑10 FID at 20 ep with balanced vs. imbalanced prior.
Configuration FID $\Delta$ vs. $\pi=0.50$
Balanced ($\pi=0.50$) $44.2$
Imbalanced ($\pi=0.45$) $\mathbf{43.1}$  $2.4\%$
020204040606080801001005050100100150150EpochFID (lower is better)WGAN (mean & 95% CI)BOLT GAN(mean & 95% CI)
[…]
Figure 2: CIFAR-10 samples at 100 epochs.
[…]
5 Conclusion
This paper has presented BOLT–GAN, a streamlined modification of the WGAN framework inspired by the Bayes optimal learning threshold (BOLT). Through extensive experiments on four diverse image‐generation benchmarks, namely CIFAR-10, CelebA-64, LSUN Bedroom-64, and LSUN Church‐64, we have demonstrated that BOLT–GAN consistently produces smoother training dynamics and lowers FID scores by 10-60% compared to the standard WGAN. These gains hold across both low‐ and high‐resolution settings, underscoring the robustness and generality of our approach. The current study introduces two extreme cases of BOLT-GAN, namely the vanilla and Lipschitz BOLT-GAN. While the former minimizes the total variation, the latter restricts a metric that is weaker than the Wasserstein distance. Exploring constraints under which BOLT-GAN framework restricts alternative measures remains an interesting direction for future research.
Appendix A Definitions and Preliminaries
A.1 Metrics
We recall the total variation (TV) distance and the Wasserstein–$p$ distance, together with basic equivalences used in our analysis.
Definition 2 (Total variation).
Let $P$ and $Q$ be probability measures on a measurable space $(\mathcal{X},\mathcal{F})$. The total variation distance is
$\mathrm{TV}(P,Q)\;\vcentcolon=\;\sup_{A\in\mathcal{F}}\big|P(A)-Q(A)\big|.$ (50)
The total variation can be computed by equivalent forms. The following lemma gives two equivalent expressions for it.
Lemma 2 (Equivalent forms of total variation (Dudley, 2002b, )).
For probability measures $P$ and $Q$ on $(\mathcal{X},\mathcal{F})$,
$\mathrm{TV}(P,Q)$ $=\sup_{f:\,\mathcal{X}\to[0,1]}\Big(\mathbb{E}_{X\sim P}[f(X)]-\mathbb{E}_{X\sim Q}[f(X)]\Big)$ (51)
$=\tfrac{1}{2}\sup_{\|g\|_{\infty}\leq 1}\Big|\mathbb{E}_{X\sim P}[g(X)]-\mathbb{E}_{X\sim Q}[g(X)]\Big|,$ (52)
and if $P$ and $Q$ are absolutely continuous w.r.t. a $\sigma$-finite measure $\mu$ with densities $P$ and $Q$,
$\mathrm{TV}(P,Q)\;=\;\tfrac{1}{2}\int_{\mathcal{X}}|p(x)-q(x)|\,\mathrm{d}\mu(x).$ (53)
Next, we define the Wasserstein distance.
Definition 3 (Wasserstein–$p$ distance).
Let $(\mathcal{X},d)$ be a metric space and $p\in[1,\infty)$. For probability measures $P$ and $Q$ on $\mathcal{X}$ with finite $p$-th moments, the Wasserstein–$p$ distance is
$W_{p}(P,Q)\;\vcentcolon=\;\Bigg(\inf_{\gamma\in\Gamma(P,Q)}\int_{\mathcal{X}\times\mathcal{X}}d(x,y)^{p}\,\mathrm{d}\gamma(x,y)\Bigg)^{\!1/p},$ (54)
where $\Gamma(P,Q)$ is the set of couplings with marginals $P$ and $Q$.
The Kantorovich–Rubinstein duality result gives an alternative expression for the Wasserstein-1 distance.
Definition 4 (Lipschitz continuity).
A function $f:(\mathcal{X},\|\cdot\|)\to\mathbb{R}$ is said to be $L$-Lipschitz if
$|f(x)-f(y)|\leq L\,\|x-y\|\qquad\forall\,x,y\in\mathcal{X}.$
Its global Lipschitz constant is denoted
$\mathrm{Lip}(f)\;\vcentcolon=\;\sup_{x\neq y}\frac{|f(x)-f(y)|}{\|x-y\|}.$
Hence $\mathrm{Lip}(f)\leq 1$ exactly when $f$ is $1$-Lipschitz.
Lemma 3 (Kantorovich–Rubinstein duality for $W_{1}$(Villani et al.,, 2009)).
On a Polish metric space $(\mathcal{X},d)$, for probability measures $P,Q$ with finite first moments,
$W_{1}(P,Q)\;=\;\sup_{\mathrm{Lip}(f)\leq 1}\Big(\mathbb{E}_{X\sim P}[f(X)]-\mathbb{E}_{Y\sim Q}[f(Y)]\Big),$ (55)
where $\mathrm{Lip}(f)\leq 1$ means $|f(x)-f(y)|\leq d(x,y)$ for all $x,y\in\mathcal{X}$.
In general, the total variation is a stronger notion of distance in the probability space. This can be observed from the following result.
Lemma 4 (Bounding $W_{1}$ by TV on bounded spaces (Villani et al.,, 2009)).
If $(\mathcal{X},d)$ has finite diameter $D\vcentcolon=\sup_{x,y}d(x,y)<\infty$, then for all probability measures $P$ and $Q$,
$W_{1}(P,Q)\;\leq\;D\,\mathrm{TV}(P,Q).$ (56)
Sketch. By Lemma 3, scale any 1‑Lipschitz $f$ to $\tilde{f}\in[0,D]$ and apply (51).
A.2 BOLT and BOLT–GAN
Throughout, we use the following functional objectives. These objectives match the definitions given in the main paper (see Sec. 3).
Definition 5 (BOLT functional).
For a measurable $h:\mathcal{X}\to[-1,0]$,
$\mathcal{L}_{\mathrm{BOLT}}(h)\;=\;q_{2}-\Big(q_{1}\,\mathbb{E}_{X\sim P_{1}}[h(X)]-q_{2}\,\mathbb{E}_{X\sim P_{2}}[h(X)]\Big).$ (57)
The universal bound of Naeini et al. (2025, Thm. 1) implies that for any such $h$, $\varepsilon_{\mathrm{Bayes}}\leq\mathcal{L}_{\mathrm{BOLT}}(h)$.
Risk and per-example BOLT loss
Let $Y\in\{1,2\}$, $X\mid(Y=i)\sim P_{i}$, and let $h:\mathcal{X}\to[-1,0]$ be measurable. Define $s(y)\vcentcolon=\mathbbm{1}\{y=2\}-\mathbbm{1}\{y=1\}\in\{-1,+1\}$ (or equivalently, $s(y)=(-1)^{y}$), . We further define the per-example BOLT loss and the population risk:
$\ell_{\mathrm{BOLT}}(z,y)$ $\;\vcentcolon=\;s(y)\,z,$ (58)
$R(h)$ $\;\vcentcolon=\;\mathbb{E}\!\big[\ell_{\mathrm{BOLT}}(h(X),Y)\big]\;=\;-\,q_{1}\,\mathbb{E}_{X\sim P_{1}}[h(X)]\;+\;q_{2}\,\mathbb{E}_{X\sim P_{2}}[h(X)].$ (59)
With the BOLT functional from Def. 5,
$\mathcal{L}_{\mathrm{BOLT}}(h)$ $=q_{2}\;-\;\Big(q_{1}\,\mathbb{E}_{X\sim P_{1}}[h(X)]\;-\;q_{2}\,\mathbb{E}_{X\sim P_{2}}[h(X)]\Big)$ (60)
$=q_{2}\;+\;R(h),$ (61)
so the minimizers of $R(h)$ coincide with the minimizers of $\mathcal{L}_{\mathrm{BOLT}}(h)$.
Definition 6 (Prior‑weighted GAN/BOLT–GAN functional).
Let $g$ be a generator inducing $P_{g}$ ,and let $h:\mathcal{X}\to[0,1]$ be a bounded critic. For prior $\pi\in(0,1)$,
$\mathcal{L}^{(\pi)}_{\mathrm{BG}}(g,h)\;\vcentcolon=\;\pi\,\mathbb{E}_{X\sim P_{\mathrm{data}}}[h(X)]\;-\;(1-\pi)\,\mathbb{E}_{X\sim P_{g}}[h(X)].$ (62)
It is worth mentioning that $\mathcal{L}^{(\pi)}_{\mathrm{BG}}(g,h)$ satisfies the complementary‑sum identity
$\mathcal{L}^{(\pi)}_{\mathrm{BG}}(g,h)+\mathcal{L}^{(1-\pi)}_{\mathrm{BG}}(g,h)\;=\;\mathbb{E}_{X\sim P_{\mathrm{data}}}[h(X)]-\mathbb{E}_{X\sim P_{g}}[h(X)].$ (63)
This is used repeatedly in our analysis to connect the BOLT-GAN objective to the TV and Wasserstein distance (see Secs. 3.2–3.3 of the main paper).
Appendix B Proofs and Further Results on BOLT
In this section we restate the result on the BOLT-assisted estimator of MAP classifier given in the main paper and provide a complete proof.
Theorem 5 (BOLT Estimator of MAP Classifier (Theorem 2 from main paper)).
Let $(X,Y)$ describe a binary classification problem with $Y\in\{1,2\}$, priors $q_{1}=\Pr\{Y=1\}$ and $q_{2}=\Pr\{Y=2\}$, class‑conditional laws $P_{1}$ and $P_{2}$, and posterior
$\eta(x)\;\vcentcolon=\;\Pr\{Y=1\mid X=x\}.$ (64)
Let $h^{\star}\in\arg\min_{h:\,\mathcal{X}\to[-1,0]}\mathcal{L}_{\mathrm{BOLT}}(h)$ (or equivalently, $h^{\star}\in\arg\min R(h)$). Then, $h^{\star}$ is pointwise optimal almost surely and takes only the endpoint values:
$h^{\star}(x)\;=\;\begin{cases}0,&\eta(x)>\tfrac{1}{2},\\[2.0pt] -1,&\eta(x)<\tfrac{1}{2},\\[2.0pt] \text{any }z\in[-1,0],&\eta(x)=\tfrac{1}{2}.\end{cases}$ (65)
Consequently, the plug‑in classifier at threshold 0.5-0.5, i.e.
$\widehat{C}(x)\;=\;1\{\,h^{\star}(x)\geq-0.5\,\},$ (66)
coincides almost everywhere with the MAP decision rule $C_{\mathrm{MAP}}(x)=1\{\eta(x)\geq\tfrac{1}{2}\}$. Moreover, if a model $h_{\theta}$ satisfies $\|h_{\theta}-h^{\star}\|_{\infty}\leq\epsilon$, then the plug‑in decision
$\widehat{C}_{\theta}(x)\;=\;\begin{cases}C_{1},&h_{\theta}(x)\geq-0.5,\\[2.0pt] C_{2},&h_{\theta}(x)<-0.5,\end{cases}$ (67)
is a plug‑in estimator of the MAP classifier.
Proof.
Fix $x\in\mathcal{X}$ and abbreviate $\eta=\eta(x)$. The conditional risk of predicting a value $z\in[-1,0]$ is
$r_{\eta}(z)$ $\;\vcentcolon=\;\mathbb{E}\!\big[\ell_{\mathrm{BOLT}}(z,Y)\mid X=x\big]\;$ (68)
$=\;\eta\,\ell_{\mathrm{BOLT}}(z,1)+(1-\eta)\,\ell_{\mathrm{BOLT}}(z,2)=$ (69)
$\;\eta(-z)+(1-\eta)z\;=\;(1-2\eta)\,z.$ (70)
To minimize $r_{\eta}(z)$ over $z\in[-1,0]$, we can say
  1. 1.
    If $\eta>\tfrac{1}{2}$, then $1-2\eta<0$. Thus, $r_{\eta}(z)$ is strictly decreasing in $z$. The minimum over [1,0][-1,0] in this case is attained at the maximal $z=0$.
  2. 2.
    If $\eta<\tfrac{1}{2}$, then $1-2\eta>0$. Thus, $r_{\eta}(z)$ is strictly increasing in $z$. The minimum therefore occurs at the minimal z=1z=-1.
  3. 3.
    If $\eta=\tfrac{1}{2}$, then $r_{\eta}(z)\equiv 0$ on [1,0][-1,0], and hence any $z\in[-1,0]$ is optimal.
This yields the pointwise characterization (65). Since $\widehat{C}(x)=1\{h^{\star}(x)\geq-0.5\}$ equals $1\{\eta(x)\geq\tfrac{1}{2}\}$ by (65), the plug‑in rule (66) coincides with MAP almost everywhere. The final claim for $h_{\theta}$ follows the fact that $\|h_{\theta}-h^{\star}\|_{\infty}\leq\epsilon$ leaves the decision boundary at 0.5-0.5 unchanged whenever $h^{\star}\in\{-1,0\}$. ∎
Remark 4 (Alternative normalization).
If one parametrizes $h:\mathcal{X}\to[0,1]$ (e.g., the bounded score after sigmoid activation), the same argument applied to the loss $\tilde{\ell}(z,y)=(-1)^{y}(2z-1)$ yields $h^{\star}(x)=1\{\eta(x)>\tfrac{1}{2}\}$ and a threshold at +0.5+0.5. The two normalizations are related by an affine map, which adjusts the threshold accordingly.
Corollary 2 (Bayes consistency of the plug‑in classifier).
Assume $\Pr\!\big(\eta(X)=\tfrac{1}{2}\big)=0$ under the data distribution of $(X,Y)$. Let $\widehat{h}_{n}:\mathcal{X}\to[-1,0]$ be risk consistent for $\ell_{\mathrm{BOLT}}$, i.e.
$R(\widehat{h}_{n})\;\xrightarrow[n\to\infty]{}\;\inf_{h}R(h)\;=\;R(h^{\star}),$ (71)
with $R$ as in (59). Define $\widehat{C}_{n}(x)=1\{\widehat{h}_{n}(x)\geq-0.5\}$. Then the 011 risk converges to the Bayes risk:
$\Pr\!\big(\widehat{C}_{n}(X)\neq Y\big)\;\xrightarrow[n\to\infty]{}\;\Pr\!\big(C_{\mathrm{MAP}}(X)\neq Y\big).$ (72)
Proof.
Write the conditional excess risk at $x$ as
$r_{\eta(x)}\big(\widehat{h}_{n}(x)\big)-r_{\eta(x)}\big(h^{\star}(x)\big)\;=\;|1-2\eta(x)|\cdot\big|\,\widehat{h}_{n}(x)-h^{\star}(x)\,\big|.$ (73)
Then, by (71), we can write
$R(\widehat{h}_{n})-R(h^{\star})=\mathbb{E}\!\big[|1-2\eta(X)|\,|\widehat{h}_{n}(X)-h^{\star}(X)|\big]\to 0.$ (74)
Fix any $\delta\in(0,1)$ and split the misclassification event as follows:
$\mathcal{E}_{n}\;\vcentcolon=\;\{\widehat{C}_{n}(X)\neq C_{\mathrm{MAP}}(X)\}\;=\;\big(\mathcal{E}_{n}\cap\{|1-2\eta(X)|\geq\delta\}\big)\;\cup\;\{|1-2\eta(X)|<\delta\}.$ (75)
For the case $|1-2\eta(X)|\geq\delta$, a decision mismatch implies $|\widehat{h}_{n}(X)-h^{\star}(X)|\geq\tfrac{1}{2}$ because $h^{\star}\in\{-1,0\}$ and the plug‑in threshold is 0.5-0.5. Therefore,
$R(\widehat{h}_{n})-R(h^{\star})$ $\;\geq\;\mathbb{E}\!\big[\,|1-2\eta(X)|\,\tfrac{1}{2}\,\mathbf{1}_{\mathcal{E}_{n}\cap\{|1-2\eta(X)|\geq\delta\}}\,\big]$ (76)
$\;\geq\;\tfrac{\delta}{2}\,\Pr\!\big(\mathcal{E}_{n}\cap\{|1-2\eta(X)|\geq\delta\}\big).$ (77)
Letting $n\to\infty$ gives $\Pr(\mathcal{E}_{n}\cap\{|1-2\eta(X)|\geq\delta\})\to 0$. Moreover, by assumption, $\Pr(|1-2\eta(X)|<\delta)\to 0$ as $\delta\downarrow 0$. We hence first send $n\to\infty$, and then $\delta\downarrow 0$ in (75) to obtain (72). ∎
B.1 Bias and Variance of BOLT
This subsection provides the full statements and proofs corresponding to Remark 3 in the main paper. We analyze the BOLT loss via plug-in estimators of the likelihood ratio together with a hinge map. The key structural fact is that the population optimizer of the BOLT bound is a hinge of the likelihood ratio; hence optimizing $h(\cdot)$ amounts to estimating $U(x)$, and errors in $h$ mirror errors in $U$.
Setup and Notation
Let $(X,C)$ be binary with classes $C\in\{C_{1},C_{2}\}$, priors $q_{i}=\Pr\{C=C_{i}\}>0$ with $q_{1}+q_{2}=1$, and conditionals $X\mid(C=C_{i})\sim P_{i}$ (densities $p_{i}$). Define the likelihood ratio $U(x)$ and the prior threshold $\tau$ as
$U(x)\;\vcentcolon=\;\frac{p_{1}(x)}{p_{2}(x)},\qquad\tau\;\vcentcolon=\;\frac{q_{2}}{q_{1}}.$ (78)
Theorem 6 (Optimal bounding function and tightness).
Let $c(x)\vcentcolon=q_{1}p_{1}(x)-q_{2}p_{2}(x)$. A pointwise maximizer of c(x)h(x)c(x)\,h(x) over $h(x)\in[-1,0]$ is
$h^{\star}(x)\;=\;\begin{cases}-1,&U(x)<\tau,\\[2.0pt] 0,&U(x)\geq\tau,\end{cases}$ (79)
and $\mathcal{L}_{\mathrm{BOLT}}(h^{\star})=\varepsilon_{\mathrm{Bayes}}$.
Proof.
We start by showing the endpoint optimality. For fixed $x$, the map $h\mapsto c(x)h$ is affine on [1,0][-1,0], hence maximized at an endpoint chosen by the sign of $c(x)$, i.e., $h^{\star}(x)=-1$ if $c(x)<0$ and $h^{\star}(x)=0$ otherwise. Since $c(x)<0$ holds if and only if $U(x)<\tau$, we can conclude (79).
We next show the tightness of the result. Using the definition of $\mathcal{L}_{\mathrm{BOLT}}$ (Def. 5, Sec. S1) and substituting $h^{\star}$, we have
$\mathcal{L}_{\mathrm{BOLT}}(h^{\star})$ $=q_{2}\;-\;\Big(q_{1}\,\mathbb{E}_{X\sim P_{1}}[h^{\star}(X)]-q_{2}\,\mathbb{E}_{X\sim P_{2}}[h^{\star}(X)]\Big)$ (80)
$=q_{2}\;-\;\int\big(q_{1}p_{1}(x)-q_{2}p_{2}(x)\big)\,h^{\star}(x)\,dx$ (81)
$=q_{2}\;-\;\int\big(q_{2}p_{2}(x)-q_{1}p_{1}(x)\big)\,1\{U(x)<\tau\}\,dx$ (82)
$=q_{2}\;-\;\int\big[\,q_{2}p_{2}(x)-q_{1}p_{1}(x)\,\big]_{+}\,dx$ (83)
$=1\;-\;\int\max\{\,q_{1}p_{1}(x),\,q_{2}p_{2}(x)\,\}\,dx\;=\;\varepsilon_{\mathrm{Bayes}}.$ (84)
Here, we have used the fact that a $1\{a>0\}=[a]_{+} a$nd $q_{2}=\int q_{2}p_{2}$. This completes the proof. ∎
The above result gives an interesting connection between the Bayes error and the likelihood ratio.
Corollary 3 (Bayes error as a hinge of likelihood ratio).
With $h^{\star}$ in (79), define the hinge map as
$t_{0}(u)\;\vcentcolon=\;[\,q_{2}-q_{1}u\,]_{+}.$ (85)
The Bayes error rate is then given by
$\varepsilon_{\mathrm{Bayes}}=q_{2}-\mathbb{E}_{X\sim P_{2}}\!\big[t_{0}\big(U(X)\big)\big].$ (86)
Proof.
Starting from (81) and using $h^{\star}=-1\{U<\tau\}$,
$q_{1}\mathbb{E}_{P_{1}}[h^{\star}]-q_{2}\mathbb{E}_{P_{2}}[h^{\star}]$ $=\int\!\big(q_{1}p_{1}-q_{2}p_{2}\big)h^{\star}\,dx=\int\!\big(q_{2}p_{2}-q_{1}p_{1}\big)\,1\{U<\tau\}\,dxlabel{eq:S3-swap}$ (87)
$=\int\!\big[\,q_{2}p_{2}-q_{1}p_{1}\,\big]_{+}\,dx=\mathbb{E}_{X\sim P_{2}}\!\big[\,q_{2}-q_{1}U(X)\,\big]_{+},$ (88)
where the last equality is obtain from dividing by as $p_{2}(x)$ on $\{p_{2}>0\}$ (the complement contributes 0). Plugging into the expression for $\mathcal{L}_{\mathrm{BOLT}}(h^{\star})$, we obtain (86). ∎
Plug-in estimator (empirical context).
Suppose we observe $M_{1}$ samples from class $C_{1}$ and $M_{2}$ from class $C_{2}$ (with total $M=M_{1}+M_{2}$), and set empirical priors $\widehat{q}_{i}\vcentcolon=M_{i}/M$. Let $\widehat{U}$ estimate $U$, and let $X^{(2)}_{1},\ldots,X^{(2)}_{M_{2}}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}P_{2}$ be the $M_{2}$ samples of class-$C_{2}$ used to estimate the expectation over $P_{2}$ in (86). Define
$\widehat{t}_{0}(u)\;\vcentcolon=\;[\,\widehat{q}_{2}-\widehat{q}_{1}\,u\,]_{+},\qquad\widehat{\varepsilon}_{\mathrm{BOLT}}\;\vcentcolon=\;\widehat{q}_{2}\;-\;\frac{1}{M_{2}}\sum_{i=1}^{M_{2}}\widehat{t}_{0}\!\big(\widehat{U}(X^{(2)}_{i})\big)$. (89)
Note $0\leq t_{0}(u)\leq q_{2}\leq 1$ and $0\leq\widehat{t}_{0}(u)\leq\widehat{q}_{2}\leq 1$, so all summands are bounded in $[0,1]$.
Lemma 5 (Lipschitz property of the hinge).
The map $t_{0}(u)=[\,q_{2}-q_{1}u\,]_{+}$ is globally $q_{1}$-Lipschitz:
$\big|\,t_{0}(u)-t_{0}(v)\,\big|\;\leq\;q_{1}\,|u-v|\qquad\forall\,u,v\in\mathbb{R}.$ (90)
Proof.
Let $\tau=q_{2}/q_{1}$. If $u,v\leq\tau$, then $t_{0}$ is affine with slope $q_{1}$, giving equality in (90). If $u,v\geq\tau$, both values are 0. If $u\leq\tau\leq v$ (or vice versa), then $t_{0}(u)-t_{0}(v)=t_{0}(u)=q_{2}-q_{1}u\leq q_{1}(\tau-u)\leq q_{1}|u-v|$. ∎
Theorem 7 (Bias and variance of the binary BOLT plug-in estimator).
Assume i.i.d. draws and bounded outputs in $[0,1]$. Then
$\big|\,\mathbb{E}[\widehat{\varepsilon}_{\mathrm{BOLT}}]-\varepsilon_{\mathrm{Bayes}}\,\big|$ $\;\leq\;q_{1}\,\mathbb{E}\!\big[\,|\widehat{U}(X)-U(X)|\,\big]\;+\;O(M^{-1/2}),$ (91)
$\mathrm{Var}\!\big(\widehat{\varepsilon}_{\mathrm{BOLT}}\big)$ $\;=\;O(M^{-1}).$ (92)
In particular, if $\mathbb{E}[\,|\widehat{U}(X)-U(X)|\,]\leq\varepsilon_{0}$, then
$\big|\,\mathbb{E}[\widehat{\varepsilon}_{\mathrm{BOLT}}]-\varepsilon_{\mathrm{Bayes}}\,\big|\;\leq\;q_{1}\,\varepsilon_{0}\;+\;O(M^{-1/2}).$ (93)
Proof.
Bias. By (86)-(89) and adding and subtracting $\mathbb{E}_{X\sim P_{2}}\!\big[t_{0}(\widehat{U}(X))\big]$, we have:
$\mathbb{E}[\widehat{\varepsilon}_{\mathrm{BOLT}}]-\varepsilon_{\mathrm{Bayes}}=\underbrace{\mathbb{E}_{X\sim P_{2}}\!\big[t_{0}(U(X))-t_{0}(\widehat{U}(X))\big]}_{\text{plug-in error}}+\underbrace{\mathbb{E}\!\Big[t_{0}(\widehat{U}(X))-\tfrac{1}{M_{2}}\sum_{i=1}^{M_{2}}\widehat{t}_{0}\big(\widehat{U}(X_{i})\big)\Big]}_{\text{sampling/priors}}.$ (94)
By Lemma 5, $t_{0}$ is $q_{1}$-Lipschitz, so $\big|\mathbb{E}_{P_{2}}[t_{0}(U)-t_{0}(\widehat{U})]\big|\leq q_{1}\,\mathbb{E}_{P_{2}}[|\widehat{U}-U|]$. For the sampling term, write $Z_{i}=\widehat{t}_{0}(\widehat{U}(X_{i}))\in[0,1]$; Hoeffding gives $\mathbb{E}\big|\tfrac{1}{M_{2}}\sum Z_{i}-\mathbb{E}Z_{1}\big|=O(M_{2}^{-1/2})$. The empirical priors $\widehat{q}_{i}=M_{i}/M$ add a zero-mean fluctuation with variance $O(M^{-1})$, hence $O(M^{-1/2})$ in expectation. With fixed class proportions ($M_{2}=\Theta(M)$), these contributions are of the same order, yielding (91).
Variance.  Using $\widehat{\varepsilon}_{\mathrm{BOLT}}=\widehat{q}_{2}-\tfrac{1}{M_{2}}\sum_{i=1}^{M_{2}}Z_{i}$ with bounded $Z_{i}$, $\mathrm{Var}\!\big(\tfrac{1}{M_{2}}\sum_{i}Z_{i}\big)=O(M_{2}^{-1})$ and $\mathrm{Var}(\widehat{q}_{2})=O(M^{-1})$. Any covariance between these two bounded averages is $O(M^{-1})$ and is dominated by the same order; hence (92). ∎
Remark 5 (Typical rates for $\varepsilon_{0}$).
We bundle approximation and estimation errors into $\varepsilon_{0}\vcentcolon=\varepsilon_{\mathrm{approx}}+\varepsilon_{\mathrm{est}}$. Under standard smoothness/capacity assumptions,
$\varepsilon_{0}\;=\;O\!\big(W^{-\gamma}+N^{-1/2}\big),$
where $W$ is a width/capacity proxy, $N$ is the number of training samples used to learn $\widehat{U}$, and $\gamma>0$ depends on the target smoothness and the approximating family (Lu et al.,, 2020; Shalev-Shwartz and Ben-David,, 2014).
Appendix C Further Results on BOLT-GAN
We present additional properties of the BOLT–GAN objective and its connections to standard distances. Throughout, we use the prior-weighted functional $\mathcal{\mathcal{}}{L}_{\mathrm{BG}}^{(\pi)}(g,h)$ as defined in the main text (Sec. 3). We do not restate it here to avoid duplication.
Basic bounds.
We first preclude degenerate values when the critic is bounded.
Lemma 6 (Lower and upper bounds).
For any generator $g$ and any bounded critic $h:\mathcal{X}\to[0,1]$,
$\pi-1\;\leq\;\mathcal{\mathcal{}}{L}_{\mathrm{BG}}^{(\pi)}(g,h)\;\leq\;\pi.$ (95)
Both bounds are tight in the sense that there exists a pair $(g,h)$ attaining them.
Proof.
Since $0\leq\mathbb{E}_{P_{\mathrm{data}}}[h]\leq 1$ and $0\leq\mathbb{E}_{P_{g}}[h]\leq 1$, $\mathcal{\mathcal{}}{L}_{\mathrm{BG}}^{(\pi)}(g,h)=\pi\,\mathbb{E}_{P_{\mathrm{data}}}[h]-(1-\pi)\,\mathbb{E}_{P_{g}}[h]\in[\pi-1,\;\pi]$. Tightness follows by choosing distributions and a measurable $h$ that separate supports: if $h=1$ on $\operatorname{supp}(P_{\mathrm{data}})$ and $h=0$ on $\operatorname{supp}(P_{g})$, the value is $\pi$; if $h=0$ on $\operatorname{supp}(P_{\mathrm{data}})$ and $h=1$ on $\operatorname{supp}(P_{g})$, it is $\pi-1$. ∎
Monotonicity in the prior.
We next show that the maximized gap grows with $\pi$.
Lemma 7 (Monotonicity in $\pi$).
Fix a generator $g$ with density $q$ and write $p$ for the data density w.r.t. a common dominating measure. Let
$D^{(\pi)}(g)\;\vcentcolon=\;\max_{h:\,\mathcal{X}\to[0,1]}\;\mathcal{\mathcal{}}{L}_{\mathrm{BG}}^{(\pi)}(g,h)\;=\;\int_{\mathcal{X}}\!\big[\pi\,p(x)-(1-\pi)\,q(x)\big]_{+}\,dx.$ (96)
Then $D^{(\pi)}(g)$ is nondecreasing in $\pi\in[0,1]$.
Proof.
For each fixed $x$, the map $\pi\mapsto[\,\pi p(x)-(1-\pi)q(x)\,]_{+}$ is nondecreasing: its derivative is $p(x)+q(x)>0$ wherever the bracket is positive and 0 elsewhere. We obtain the desired result by integrating over $\mathcal{X}$. ∎
Balanced vs. prior‑weighted gap (Lipschitz critics).
We now relate the prior‑weighted objective to its “balanced” counterpart on a 1‑Lipschitz class. Let:
$\mathcal{H}_{\mathrm{Lip}}$ $\;\vcentcolon=\;\big\{\,h:\mathcal{X}\!\to\![0,1]\;:\;|h(x)-h(y)|\leq\|x-y\|\ \forall\,x,y\in\mathcal{X}\,\big\},$ (97)
$D^{(\pi)}_{\mathrm{Lip}}(g)$ $\;\vcentcolon=\;\sup_{h\in\mathcal{H}_{\mathrm{Lip}}}\mathbb{E}_{X\sim P_{\mathrm{data}}}\!\big[h(X)\big]-(1-\pi)\,\mathbb{E}_{X\sim P_{g}}\!\big[h(X)\big],$ (98)
$\Sigma_{\mathrm{Lip}}(g)$ $\;\vcentcolon=\;\sup_{h\in\mathcal{H}_{\mathrm{Lip}}}\Big(\mathbb{E}_{X\sim P_{\mathrm{data}}}\!\big[h(X)\big]-\mathbb{E}_{X\sim P_{g}}\!\big[h(X)\big]\Big).$ (99)
The next inequality is the pointwise scalar comparison we use repeatedly.
Lemma 8 (Pointwise dominance).
For $0<\pi\leq\tfrac{1}{2}$ and any $a,b\in[0,1]$,
$\max\{\,a-b,\;b-a\,\}\;\geq\;\pi\,a-(1-\pi)\,b.$ (100)
Proof.
If $a\geq b$, then
$(a-b)-\big(\pi a-(1-\pi)b\big)=(1-\pi)a-\pi b\;\geq\;(1-\pi)b-\pi b=(1-2\pi)b\geq 0.$
If $a<b$, the claim is equivalent to
[…]
which holds since $a\leq b$ and $(2-\pi)/(1+\pi)\geq 1$ for $\pi\leq\tfrac{1}{2}$. ∎
Proposition 1 (From pointwise to functional dominance).
For $0<\pi\leq\tfrac{1}{2}$,
$D^{(\pi)}_{\mathrm{Lip}}(g)\;\leq\;\Sigma_{\mathrm{Lip}}(g).$ (101)
Proof.
For any $h\in\mathcal{H}_{\mathrm{Lip}}$, write $a=\mathbb{E}_{P_{\mathrm{data}}}[h]$ and $b=\mathbb{E}_{P_{g}}[h]$. By Lemma 8, $\max\{a-b,b-a\}\geq\pi a-(1-\pi)b$. Since $h\mapsto 1-h$ preserves $\mathcal{H}_{\mathrm{Lip}}$ and flips aba-b to bab-a,
$\sup_{h\in\mathcal{H}_{\mathrm{Lip}}}\max\{a-b,b-a\}=\sup_{h\in\mathcal{H}_{\mathrm{Lip}}}(a-b)=\Sigma_{\mathrm{Lip}}(g).$
Taking suprema on both sides over $h\in\mathcal{H}_{\mathrm{Lip}}$ yields the claim. ∎
Bridge to $W_{1}$ (pointer).
Combining Proposition 1 with the complementary‑sum identity from the main paper and the Kantorovich–Rubinstein dual (sup over all $1$-Lipschitz functions without range constraints) gives the bound $D^{(\pi)}_{\mathrm{Lip}}(g)\leq W_{1}(P_{\mathrm{data}},P_{g})$ used in the main text. We omit the re‑statement here to avoid duplication.
Role of the basic bound (Lemma 6).
Together with the balanced‑sum identity in the paper (Sec. 3), Lemma 6 ensures that each prior‑weighted term is uniformly bounded when $h\in[0,1]$, preventing pathological cancellations when one passes to the balanced form in intermediate maximizations.
Symmetry for $\pi>\tfrac{1}{2}$.
The case $\pi>\tfrac{1}{2}$ follows by exchanging labels (replacing $\pi$ with $1-\pi$, i.e., swaping $P_{\mathrm{data}}$ and $P_{g}$). Thus the conclusions above extend to all $\pi\in(0,1]$ by symmetry.
In summary, for $0<\pi\leq 1$,
$D^{(\pi)}_{\mathrm{Lip}}(g)\;\leq\;\Sigma_{\mathrm{Lip}}(g)\;\leq\;W_{1}(P_{\mathrm{data}},P_{g}),$
This is precisely the inequality chain invoked in the proof of the Lipschitz BOLT–GAN result in the main paper.
Appendix D Implementation of Lipschitz BOLT-GAN
This section records the implementation details for our Lipschitz BOLT–GAN experiments. We explain the properties we rely on and the techniques used to enforce the $1$-Lipschitz constraint on the discriminator (critic). As discussed in the main paper, Lipschitz continuity is key both (i) to recover meaningful dual objectives (e.g., the KR dual for $W_{1}$) and (ii) to stabilize adversarial training; see the TV/Wasserstein connections in Secs. 3.2–3.3. In our code, the gradient penalty is applied to the raw critic output $\tilde{h}_{\theta}$ (pre‑activation), and the bounded score $h_{\theta}=\sigma(\tilde{h}_{\theta})\in[0,1]$ is used inside the BOLT objective, as summarized in Algorithm 1. (Gulrajani et al.,, 2017; Miyato et al.,, 2018, Sec. 3).
D.1 Preliminaries on Lipschitz continuity
Lemma 9 (Basic Lipschitz calculus (Miyato et al.,, 2018)).
Let $f$, $g$, $f_{1}$, and $f_{2}$ be Lipschitz. Then
$\mathrm{Lip}(g\circ f)$ $\leq\mathrm{Lip}(g)\,\mathrm{Lip}(f),$ (102)
$\mathrm{Lip}(f_{1}+f_{2})$ $\leq\mathrm{Lip}(f_{1})+\mathrm{Lip}(f_{2}).$ (103)
Lemma 10 (Layerwise Lipschitz constants in common architectures (Miyato et al.,, 2018; Anil et al.,, 2019)).
Assume all vector spaces are equipped with the Euclidean norm $\|\cdot\|_{2}$, and for a function $f$ we write $\mathrm{Lip}(f)$ for its global Lipschitz constant with respect to $\|\cdot\|_{2}$.
  1. (i)
    Linear / convolutional layers. For an affine map $T(x)=Wx+b$ (fully connected) or a convolutional layer viewed as a linear map $x\mapsto Wx$,
    $\mathrm{Lip}(T)=\|W\|_{2},$
    i.e., the operator (spectral) norm of $W$. Biases do not affect $\mathrm{Lip}$.
  2. (ii)
    Pointwise activations (elementwise maps). If $\phi$ is applied elementwise, then $\mathrm{Lip}(\phi)=\sup_{t\in\mathbb{R}}|\phi^{\prime}(t)|$ (when the derivative exists a.e.).
  3. (iii)
    Residual blocks. For a residual block of the form $x\mapsto Sx+F(x)$ with a linear skip $S$ (identity or $1{\times}1$ projection),
    $\mathrm{Lip}(x\mapsto Sx+F(x))\;\leq\;\|S\|_{2}+\mathrm{Lip}(F).$
    In particular, if $S=I$, then $\mathrm{Lip}\leq 1+\mathrm{Lip}(F)$.
  4. (iv)
    Resampling operators. If $S$ is a (linear) down/up‑sampling operator used before or after a block $F$, then
    $\mathrm{Lip}(F\!\circ\!S)\leq\mathrm{Lip}(F)\,\|S\|_{2},\qquad\mathrm{Lip}(S\!\circ\!F)\leq\|S\|_{2}\,\mathrm{Lip}(F).$
    Common strided convolutions, pooling, and interpolation are linear maps with finite $\|S\|_{2}$ that must be accounted for in per‑layer bounds.
Remark 6 (Standard elementwise activations and their $\mathrm{Lip}$).
The following are explicit formulas for common activations used in (ii), together with their Lipschitz constants:
$\mathrm{ReLU}(t)=\max\{0,t\},\quad\mathrm{Lip}=1;\qquad\tanh(t)=\frac{e^{t}-e^{-t}}{e^{t}+e^{-t}},\quad\mathrm{Lip}=1;$
$\mathrm{LeakyReLU}_{\alpha}(t)=\begin{cases}t,&t\geq 0\\ \alpha t,&t<0\end{cases},\quad\mathrm{Lip}=\max\{1,\alpha\};\qquad\sigma(t)=\frac{1}{1+e^{-t}},\quad\sigma^{\prime}(t)=\sigma(t)\big(1-\sigma(t)\big)\leq\tfrac{1}{4},\ \mathrm{Lip}=\tfrac{1}{4}.$
D.2 Spectral Normalization (SN)
Method. Each linear layer $W$ is rescaled by an estimate of its top singular value $\sigma(W)$ via power iteration with persistent vectors $(u,v)$: $\bar{W}\leftarrow W/\sigma(W)$. For convolutions, $\sigma(W)$ is the operator norm of the induced linear map and can be estimated by alternating conv/transpose‑conv passes. (Miyato et al.,, 2018).
Guarantee. If every linear/convolutional layer is normalized so that $\|W\|_{2}\leq 1$ and all activations are $1$-Lipschitz, then by Lemmas 910 the whole network is $1$-Lipschitz (up to residual additions; see “Residual blocks” below). This yields strong stability at low computational cost and is widely adopted in GAN critics. (Miyato et al.,, 2018).
Residual blocks and skips. Because $\mathrm{Lip}(x+F(x))\leq 1+\mathrm{Lip}(F)$, residual connections can increase the global constant above $1$. Two common remedies are: (i) apply SN to the skip $1{\times}1$ projection; (ii) scale residual branches by $c<1$ (“residual scaling”) to keep $\mathrm{Lip}(F)\leq c$, so the block is $\leq 1+c$.
D.3 Gradient Penalty (GP)
Method. Gradient penalty encourages $\|\nabla_{x}\tilde{h}_{\theta}(x)\|_{2}\approx 1$ on a chosen sampling distribution. The most common choice is the interpolation penalty (Gulrajani et al.,, 2017): draw $(x_{\mathrm{real}},x_{\mathrm{fake}})$, form $\hat{x}=\alpha x_{\mathrm{real}}+(1-\alpha)x_{\mathrm{fake}}$ with $\alpha\sim\mathrm{Unif}[0,1]$, and add
$\lambda\,\mathbb{E}_{\hat{x}}\!\big(\,\|\nabla_{\hat{x}}\tilde{h}_{\theta}(\hat{x})\|_{2}-1\,\big)^{2}$ (104)
to the critic loss. We penalize pre‑activation $\tilde{h}_{\theta}$, not $\sigma(\tilde{h}_{\theta})$, to avoid shrinking gradients through the $1/4$-Lipschitz sigmoid.
Guarantee (local, not global). The penalty (104) is a soft constraint: it is zero iff the gradient norm equals $1$ on the support where it is evaluated (here, segments between real/fake samples). It does not by itself prove a global $1$-Lipschitz bound, but it aligns with the KR optimal‑critic condition and markedly improves training stability. (Gulrajani et al.,, 2017).
Implementation notes.
  • Backprop through the norm. Ensure gradients flow through $\|\nabla_{\hat{x}}\tilde{h}_{\theta}(\hat{x})\|_{2}$ (autodiff: create graph for higher‑order grads).
  • Normalization. Avoid BatchNorm in the critic; prefer layer/instance norm or none.
  • Lazy regularization. Apply the penalty every $k$ steps and scale $\lambda\leftarrow k\lambda$ to keep the expected contribution unchanged.
  • Tuning $\lambda$. Start with $\lambda\in[1,10]$ and monitor the penalty/critic ratio (see diagnostics below).
D.4 Weight clipping
Method. Clip each parameter after every update: $w\leftarrow\mathrm{clip}(w,\,-\mathcal{\mathcal{}}{L}_{\mathrm{wc}},\,\mathcal{\mathcal{}}{L}_{\mathrm{wc}})$ (Arjovsky et al.,, 2017). This loosely bounds per‑layer operator norms via element‑wise bounds.
Guarantee (coarse). If a layer $W\in\mathbb{R}^{m\times n}$ is element‑wise clipped as $|W_{ij}|\leq\mathcal{\mathcal{}}{L}_{\mathrm{wc}}$, then
$\|W\|_{2}\;\leq\;\|W\|_{\mathrm{F}}\;\leq\;\sqrt{mn}\,\mathcal{\mathcal{}}{L}_{\mathrm{wc}},$ (105)
so a depth‑$L$ network with 11-Lipschitz activations satisfies $\mathrm{Lip}(f)\leq\prod_{\ell=1}^{L}\sqrt{m_{\ell}n_{\ell}}\,\mathcal{\mathcal{}}{L}_{\mathrm{wc}}$. This yields at best a coarse global bound and often harms capacity. In practice we prefer SN or GP. (Arjovsky et al.,, 2017; Miyato et al.,, 2018).
D.5 Other gradient penalties and related methods
  • One‑sided GP (Lipschitz penalty). Penalize only $\|\nabla\tilde{h}_{\theta}\|_{2}>1$: $\lambda\,\mathbb{E}[\max(0,\|\nabla\tilde{h}_{\theta}\|-1)^{2}]$, avoiding a push toward <1<1 gradients that can reduce capacity (Petzka et al.,, 2018).
  • Zero‑centered penalties $R_{1}/R_{2}$. Apply $\lambda\,\mathbb{E}_{x\sim P_{\mathrm{data}}}\|\nabla_{x}\tilde{h}_{\theta}(x)\|_{2}^{2}$ ($R_{1}$) or $\lambda\,\mathbb{E}_{x\sim P_{g}}\|\nabla_{x}\tilde{h}_{\theta}(x)\|_{2}^{2}$ ($R_{2}$) to improve convergence empirically—even though they do not strictly enforce $1$-Lipschitzness (Mescheder et al.,, 2018).
  • DRAGAN (local penalties). Sample $\hat{x}=x_{\mathrm{real}}+\delta$ with small noise and penalize $\|\nabla_{\hat{x}}\tilde{h}_{\theta}(\hat{x})\|_{2}$ near the data manifold to promote local smoothness (Kodali et al.,, 2017).
  • Orthogonal/Parseval constraints. Constrain $W^{\top}W\approx I$ (Parseval/Björck) to shrink $\|W\|_{2}$ toward $1$ at higher compute cost (Cissé et al.,, 2017).
  • Lipschitz activations. Using activations with known Lipschitz constants (e.g., GroupSort) can yield provably $1$-Lipschitz networks when combined with spectral control of linear layers (Anil et al.,, 2019).
  • Optimizer‑level gradient clipping. Stabilizes updates but does not impose a function‑level Lipschitz bound; . We emphasize that this is different from SN or GP.
These alternatives differ in how strictly they enforce the Lipschitz constraint versus how much computational cost or flexibility they introduce, and we mention them here for completeness.
D.6 Diagnostics and failure modes
We use the following techniques to keep the critic near the $1$-Lipschitz regime while preserving capacity and useful gradients to the generator.
  • Monitor gradient norms. Track the empirical distribution of $\|\nabla_{\hat{x}}\tilde{h}_{\theta}(\hat{x})\|_{2}$ (for $\hat{x}$ used in the penalty). For WGAN‑GP, it should concentrate near $1$. Heavy mass $\gg 1$: increase $\lambda$ or apply the penalty more frequently. Heavy mass $\ll 1$: reduce $\lambda$ or switch to one‑sided GP.
  • Penalty/critic ratio. Log $\mathbb{E}\big[(\|\nabla\|-1)^{2}\big]/\mathbb{E}[\text{critic term}]$. Ratios $\ll 10^{-2}$ indicates an ineffective penalty; $\gg 10$ indicates capacity strangling.
  • Check residual branches. With SN, ensure skip paths are constrained (SN on $1{\times}1$ skips or residual scaling); otherwise the global constant can inflate.
  • Avoid BatchNorm in critics. Batch‑dependent statistics interact poorly with gradient penalties and can inject noise into $\|\nabla_{x}\tilde{h}_{\theta}\|$; use layer/instance norm or none.
Appendix E Additional Experimental Results
E.1 CelebA–64 $\lambda_{\mathrm{GP}}$ Sensitivity for BOLT–GAN
For completeness, Figure 4 presents the CelebA–64 $\lambda_{\mathrm{GP}}$ sweep mirroring the CIFAR-10 analysis in the main text. The stability/performance trends are consistent: $\lambda_{\mathrm{GP}}\!\in\![5,10]$ is best, with variations within $\pm 1$ FID.
[…]
E.2 Qualitative Results
We provide qualitative samples to complement the quantitative results in the main paper: each panel below shows an uncurated $4\times 4$ grid of random images generated by BOLT–GAN–GP after 20 training epochs under the same setup as our primary experiments (same architectures, optimizer, and $\lambda_{\mathrm{GP}}$).
Figure 5: Qualitative samples (first experiment of the main paper). BOLT-GAN-GP after 20 epochs with $\lambda_{\mathrm{GP}}{=}10$. Top-left: CIFAR-10; top-right: CelebA-64; bottom-left: LSUN Bedroom-64; bottom-right: LSUN Church-64. Each panel shows the same $4\times 4$ grid of generated samples at a reduced scale. Under the identical setup used in our main experiments, Lipschitz BOLT–GAN produces sharper textures, fewer checkerboard artifacts, and more stable color statistics than the WGAN–GP baseline.
Appendix F Implementation Details
This section records everything needed to reproduce our results, together with the rationale for each design choice and how the pieces interact (objective, Lipschitz control, optimization, and evaluation). The discriminator produces a raw scalar score $\tilde{h}_{\theta}(x)\in\mathbb{R}$ and a bounded score $h_{\theta}(x)=\sigma(\tilde{h}_{\theta}(x))\in[0,1]$ that is used in the BOLT objective. All gradient penalties are applied to the raw score $\tilde{h}_{\theta}$.
F.1 Codebase, Environment, and Hardware
We use Python 3.10+ and the CUDA-enabled build of PyTorch. To improve throughput and reduce memory footprint without affecting FID, we enable automatic mixed precision (AMP) for both networks. When we report “deterministic” runs, we enable determinism switches so that snapshots and sample grids can be reproduced exactly, and we average metrics over multiple random seeds. All experiments were run on NVIDIA A100 GPUs.
F.2 Datasets and Preprocessing
We intentionally keep preprocessing minimal in order to maintain comparability of FID with prior work. Unless otherwise stated, we do not use data augmentation. For CIFAR‑10, we train at the native $32{\times}32$ resolution and normalize images to [1,1][-1,1]; for visualization only, we upsample to $64{\times}64$ using nearest-neighbor interpolation. For CelebA‑64, we center‑crop faces, resize to $64{\times}64$, and normalize to [1,1][-1,1]. For LSUN Bedroom/Church‑64, we apply a square center‑crop, resize to $64{\times}64$, and normalize to [1,1][-1,1].
F.3 Architectures
We adopt a residual DCGAN‑style backbone because it offers a good balance between stability and capacity at the $64{\times}64$ resolution. The discriminator (critic) does not use BatchNorm in order to avoid interactions with gradient penalties. when spectral methods are used, we treat residual paths carefully.
Generator $g_{\varphi}$ (latent $\dim z=64$).
A linear layer projects $z\sim\mathcal{N}(0,I_{64})$ to a $4{\times}4{\times}(4\cdot\mathrm{GF})$ feature map. We then apply a stack of upsampling residual blocks whose depth depends on the target resolution: four blocks for $64{\times}64$ (CelebA and LSUN) and three blocks for $32{\times}32$ (CIFAR‑10).
Discriminator (critic) $\tilde{h}_{\theta}$.
The critic receives images at dataset resolution and applies a stack of downsampling residual blocks mirroring the generator. For $64{\times}64$, we use four downsampling residual blocks with stride‑2 $3{\times}3$ convolutions (or the standard ResBlockD average‑pool projection), two $3{\times}3$ convolutions per block, and LeakyReLU activations with negative slope $0.2$; the skip path uses a stride‑2 $1{\times}1$ projection. For $32{\times}32$, we use three downsampling residual blocks of the same design. A global sum‑pool feeds a linear head that outputs the raw scalar $\tilde{h}_{\theta}(x)$. Unless otherwise noted, the critic uses no BatchNorm (i.e., we only consider instance or layer normalization in explicit ablations).
F.4 Objectives and Priors
[…]
$\mathcal{L}_{\mathrm{BOLT}}(\varphi,\theta)=\pi\,\frac{1}{B}\sum_{j=1}^{B}h_{j}\;-\;(1-\pi)\,\frac{1}{B}\sum_{j=1}^{B}\hat{h}_{j}.$
For reference, we also evaluate a non‑Lipschitz baseline by setting the gradient‑penalty coefficient to zero $(\lambda_{\mathrm{GP}}=0)$. As expected, this variant is numerically unstable and yields high FID; see Sec. S5 for details.
F.5 Lipschitz Enforcement
We enforce approximate $1$-Lipschitz behavior of the critic using the WGAN‑GP gradient penalty applied to the raw score $\tilde{h}_{\theta}$. Concretely, we sample interpolants $\tilde{x}=\alpha x+(1-\alpha)\hat{x}$ with $\alpha\sim\mathrm{Unif}[0,1]$ and add
$\mathrm{GP}\;=\;\frac{\lambda_{\mathrm{GP}}}{B}\sum_{j=1}^{B}\Big(\,\|\nabla_{\tilde{x}_{j}}\tilde{h}_{\theta}(\tilde{x}_{j})\|_{2}-1\Big)^{2}$
to the critic objective. We set $\lambda_{\mathrm{GP}}=10$ as our default because this value is the de facto choice introduced by WGAN‑GP and widely adopted in subsequent implementations; it offers a robust trade‑off between stability and capacity in practice (Gulrajani et al.,, 2017; Mescheder et al.,, 2018). In Sec. S5 we also report a small sweep over $\lambda_{\mathrm{GP}}\in\{1,5,10,20\}$ that confirms the default is near‑optimal under our settings.
F.6 Optimization and Training Schedule
We optimize both networks with Adam using $\beta_{1}=0.5$, $\beta_{2}=0.999$, and learning rates $\eta_{\mathrm{d}}=\eta_{\mathrm{g}}=2\times 10^{-4}$. Each iteration performs $n_{\mathrm{d}}=5$ critic updates followed by one generator update, a ratio that we found to be consistently reliable across datasets. Unless otherwise stated, the learning rate is held constant. We optionally maintain an exponential moving average (EMA) of the generator parameters with decay $0.999$ for evaluation, but we leave EMA off by default to match baseline comparisons. The batch size is $B=64$; when memory is limited, we use gradient accumulation to preserve the effective batch size. Unless otherwise specified, figures correspond to the 20‑epoch checkpoint.
F.7 Training Algorithm (Overview and Pseudocode)
[…]
Algorithm 1 BOLT–GAN training with gradient penalty (notation matched to paper)
1:
  • Discriminator (raw score) $\tilde{h}_{\theta}$, generator $g_{\varphi}$
  • Data distribution $P_{\mathrm{data}}$, latent distribution $P_{Z}$
  • Prior $\pi\in(0,1]$
  • Batch size $B$, critic updates per generator update $n_{\mathrm{d}}$
  • Learning rates $\eta_{\mathrm{d}},\eta_{\mathrm{g}}$, iterations $T$
  • Gradient-penalty coefficient $\lambda_{\mathrm{gp}}$
2:trained parameters $\theta,\varphi$
3:for iteration =1=1 to $T$ do
4:  for $k=1$ to $n_{\mathrm{d}}$ do
5:   Sample $\{x_{j}\}_{j=1}^{B}\sim P_{\mathrm{data}}$;  sample $\{z_{j}\}_{j=1}^{B}\sim P_{Z}$;  $\hat{x}_{j}\leftarrow g_{\varphi}(z_{j})$
6:   Raw scores: $f_{j}\leftarrow\tilde{h}_{\theta}(x_{j})$,   $\hat{f}_{j}\leftarrow\tilde{h}_{\theta}(\hat{x}_{j})$;  bounded: $h_{j}\leftarrow\sigma(f_{j})$,   $\hat{h}_{j}\leftarrow\sigma(\hat{f}_{j})$
7:   (a) Batch estimator of $\mathcal{L}_{\mathrm{BOLT}}$ for fixed $\theta$:
${\mathcal{L}_{\mathrm{BOLT}}}(\varphi,\theta)\;=\;\pi\cdot\frac{1}{B}\sum_{j=1}^{B}h_{j}\;-\;(1-\pi)\cdot\frac{1}{B}\sum_{j=1}^{B}\hat{h}_{j}$
8:   (b) Gradient penalty on raw score (WGAN–GP style):
9:   Sample $\{\alpha_{j}\}_{j=1}^{B}\sim\mathrm{Uniform}(0,1)$;  $\tilde{x}_{j}=\alpha_{j}x_{j}+(1-\alpha_{j})\hat{x}_{j}$
$\mathrm{GP}\;=\;\frac{\lambda_{\mathrm{gp}}}{B}\sum_{j=1}^{B}\Big(\,\|\nabla_{\tilde{x}_{j}}\tilde{h}_{\theta}(\tilde{x}_{j})\|_{2}-1\Big)^{2}$
10:   (c) Critic loss & update (maximize $\mathcal{L}_{\mathrm{BOLT}}$):
$\mathcal{L}_{\theta}\;=\;-\,{\mathcal{L}_{\mathrm{BOLT}}}(\varphi,\theta)+\mathrm{GP}\quad;\quad\theta\leftarrow\theta-\eta_{\mathrm{d}}\,\nabla_{\theta}\mathcal{L}_{\theta}$
11:  end for
12:  Generator update (minimize $\mathcal{L}_{\mathrm{BOLT}}$):
13:  Sample $\{z^{\prime}_{j}\}_{j=1}^{B}\sim P_{Z}$;  $\hat{x}^{\prime}_{j}\leftarrow g_{\varphi}(z^{\prime}_{j})$;  $\hat{h}^{\prime}_{j}\leftarrow\sigma\big(\tilde{h}_{\theta}(\hat{x}^{\prime}_{j})\big)$
$\mathcal{L}_{\varphi}\;=\;-\,(1-\pi)\cdot\frac{1}{B}\sum_{j=1}^{B}\hat{h}^{\prime}_{j}\quad;\quad\varphi\leftarrow\varphi-\eta_{\mathrm{g}}\,\nabla_{\varphi}\mathcal{L}_{\varphi}$
14:end for
F.8 Hyperparameter Sweeps and Selection
We sweep $\lambda_{\mathrm{GP}}\in\{1,5,10,20\}$ separately for each method and dataset, and we report the best FID per method within a fixed compute budget. Other knobs such as $\beta_{1}$ and $n_{\mathrm{d}}$ have a comparatively minor effect in our ranges (e.g., within $\pm 1$ FID on CIFAR‑10).
F.9 Evaluation Metrics
Fréchet Inception Distance (FID).
We use FID as the primary quantitative measure of image‑generation quality. FID computes the Fréchet distance between multivariate Gaussian embeddings of real and generated images in the feature space of a pre‑trained Inception‑V3 network (Heusel et al.,, 2017). Lower values indicate that the generated distribution is closer to the real distribution in terms of both perceptual and statistical similarity. Unless noted otherwise, all FID scores are computed on 50k generated samples using the official Inception statistics for each dataset.
Evaluation Protocol.
Unless stated otherwise, we evaluate every checkpoint with 50k samples and the official Inception statistics. Snapshot selection and the choice of $\lambda_{\mathrm{GP}}$ follow the settings detailed in Sec. S7.
F.10 Logging and Diagnostics
We monitor diagnostics that directly reflect Lipschitz control and training health. In particular, we track the empirical distribution of $\|\nabla_{\tilde{x}}\tilde{h}_{\theta}(\tilde{x})\|_{2}$ on the interpolants used for the penalty, the value of the penalty term itself, and the ratio $\mathbb{E}[(\|\nabla\|-1)^{2}]/\mathbb{E}[\text{critic term}]$. Healthy runs concentrate gradient norms near $1$; persistent mass well above $1$ suggests increasing $\lambda_{\mathrm{GP}}$ or applying the penalty more frequently, whereas mass well below $1$ suggests reducing $\lambda_{\mathrm{GP}}$ or switching to a one‑sided penalty.
F.11 Determinism and Seeds
For exact reproducibility we fix a global seed, propagate per‑worker dataloader seeds, and set torch.backends.cudnn.deterministic=True and benchmark=False. We seed Python, NumPy, and PyTorch random number generators. We release configuration files, grid‑reconstruction scripts, and commit hashes that match the figures and tables.
F.12 Failure Modes and Stability Notes
The non‑Lipschitz BOLT variant ($\lambda_{\mathrm{GP}}=0$) frequently diverges or produces very high FID, which is consistent with optimizing a TV‑strength objective. Combining strong spectral normalization with a large gradient‑penalty coefficient can over‑regularize the critic; in such cases we prefer to rely primarily on a single Lipschitz control. We also avoid BatchNorm in the critic. If the critic saturates, reducing the discriminator learning rate or increasing $n_{\mathrm{d}}$ usually restores stable training.
The training code is available in our anonymous repository.

### v2

@@
 Abstract
-We introduce BOLT-GAN, a simple yet effective modification of the WGAN framework inspired by the Bayes optimal learning threshold (BOLT).
-We show that with a Lipschitz continuous discriminator, BOLT-GAN implicitly minimizes a different metric distance than the Earth-Mover (Wasserstein) distance, and also has better training stability.
-Empirical evaluations on four standard image-generation benchmarks, (CIFAR-10, CelebA-64, LSUN Bedroom-64, and LSUN Church‐64) show that BOLT-GAN consistently outperforms WGAN, achieving 10-60% lower Fréchet Inception Distance (FID).
-Our results suggest that BOLT is a broadly applicable principle for enhancing GAN training.
-The software is available in our anonymous repository.
+We introduce BOLT-GAN, a novel framework for stable GAN training using the Bayes optimal learning threshold (BOLT).
+The discriminator is trained via the BOLT loss under a standard 1-Lipschitz constraint.
+This guides the generator to maximize the Bayes error of the discrimination task.
+We show that the training objective in this case represents a class of metrics on probability measures controlled by a 1-Lipschitz discriminator minimizing an integral probability metric that is upper-bounded by Wasserstein-1 distance.
+Across four standard image-generation benchmarks, BOLT-GAN improves FID and precision/recall over benchmark GAN frameworks under identical architectures and training budgets.
+Our experimental findings further confirm the advantage of linking the GAN training objective to a min-max Bayes error criterion.
+Implementation software is available in our anonymous repository.
 1 Introduction
-The Bayes optimal learning threshold (BOLT) (Naeini et al.,, 2025) is a recently proposed framework for discriminative learning.
-In its basic form, BOLT develops a computational approach to estimate the Bayes error rate, i.e., the lowest classification error achievable by a representation, by sampling an arbitrary learning model.
-This introduces the BOLT loss, whose minimization drives the model toward the Bayes optimal classifier.
-This work adapts the BOLT framework to generative modeling by training the discriminator of a generative adversarial network (GAN) via the BOLT loss.
-We treat the GAN discriminator as a binary classifier and train it using the BOLT framework.
-Under this framework, the BOLT objective recovers the well-known Wasserstein GAN (WGAN) (Arjovsky et al.,, 2017) formulation, up to a minor adjustment: the discriminator outputs are constrained to a bounded range.
-This perspective provides both a Bayesian interpretation of adversarial learning and a pathway to improve training stability.
-Building on this foundation, we propose the BOLT-GAN, which inherits WGAN-style convergence guarantees while driving discriminator performance toward Bayes optimality.
+The Bayes optimal learning threshold (BOLT) (Naeini et al., 2025) is a recently proposed framework for discriminative learning.
+In its basic form, BOLT provides a computational approach to upper-bound (and tightly approximate) the Bayes error rate (BER) by sampling an arbitrary bounded learning model.
+Minimizing the resulting BOLT loss, aligns the model with a Bayes optimal, i.e., maximum-a-posteriori (MAP), classifier.
+Generative adversarial networks (GANs) learn data distribution through a two-player game between a generator and a discriminator.
+Despite their success, vanilla GANs trained with the cross-entropy loss and Jensen-Shannon divergence are notoriously unstable, often suffering from vanishing gradients and mode collapse.
+Since the discriminator’s role is inherently a binary classification task (real versus fake), it is natural to ask whether a Bayes-error-motivated discriminator objective can improve generative training dynamics.
+In this work, we embed the BOLT loss inside the adversarial game: we train the discriminator with the BOLT objective and train the generator to maximize the induced Bayes error rate of the discrimination task.
+We refer to this adversarial formulation as a maximum Bayes error (Max-BER) game.
+Our analysis further shows that the Max-BER game induces a discrepancy between the data distribution and the generator distribution that is controlled by the discriminator.
+With unregularized bounded-discriminator, this game reduces under balanced priors to minimizing the total variation (TV), a strong metric.
+This can yield poor gradient propagation and severe numerical instability.
+By restricting the discriminator class to bounded 1-Lipschitz functions, we obtain a weaker but smoother metric: an integral probability metric (IPM) that is upper-bounded by the Wasserstein-1 distance.
+We refer to this constrained min-max formulation as BOLT-GAN.
+Unlike existing GAN formulations, BOLT-GAN provides both a Bayesian interpretation of adversarial training and a pathway to improved stability.
 1.1 Background and Related Work
 Bayes error estimation
-The Bayes error rate is a central concept in statistical learning, characterizing the minimum error of a classification task (Bishop,, 2006).
-Classical approaches estimate this error rate via distance-based measures such as Bhattacharyya or Mahalanobis distances (Bhattacharyya,, 1943; Mahalanobis,, 1936).
-These methods often fail in high dimensions (Devroye et al.,, 1996).
-More recent methods employ nonparametric or model-based estimators, such as $k$-nearest neighbors (Moon et al.,, 2017) or $\epsilon$-ball techniques (Noshad et al.,, 2019).
-The BOLT framework (Naeini et al.,, 2025) provides a unifying alternative, deriving a bound on Bayes error from classifier outputs and introducing a loss that aligns learning with approximation of the Bayes optimal classifier.
+The Bayes error rate (BER) characterizes the minimum achievable classification error for a given data-generating process (Bishop, 2006).
+Classical BER bounds and estimators are often expressed via distance-like quantities between class-conditional distributions (e.g., Bhattacharyya- or Mahalanobis-type measures), but they can be inaccurate in high dimensions and under model mismatch (Devroye et al., 1996).
+Modern nonparametric approaches estimate BER using nearest-neighbor and local neighborhood constructions (Moon et al., 2017; Noshad et al., 2019).
+The Bayes optimal learning threshold (BOLT) framework (Naeini et al., 2025) provides a universal upper bound on BER based on bounded discriminator outputs and induces a surrogate loss (the BOLT loss) whose minimization aligns a model with the Bayes-optimal, i.e.
+maximum-a-posteriori (MAP), classifier.
 GAN
-GAN (Goodfellow et al.,, 2014) consists of a generator, which synthesizes data samples from a latent variable, and a discriminator, which distinguishes between real and generated samples.
-It is trained via a min-max game between the generator and discriminator.
-While powerful, the original formulation based on Jensen–Shannon divergence is unstable (Goodfellow et al.,, 2014; Arjovsky et al.,, 2017).
-Numerous alternatives have been proposed: Wasserstein GAN (WGAN) (Arjovsky et al.,, 2017) leverages the Lipschitz continuity of the discriminator to improve gradient backpropagation.
-This is implemented by approaches such as WGAN-GP (Gulrajani et al.,, 2017) which enforce the Lipschitz constraint via gradient penalty.
-Other variants include $f$-GAN (Nowozin et al.,, 2016), which generalizes the objective to any $f$-divergence, LS-GAN (Mao et al.,, 2017), BEGAN (Berthelot et al.,, 2017), and techniques such as spectral normalization (Miyato et al.,, 2018) or progressive growing (Karras et al.,, 2018).
-These lines of work aim to stabilize training, but do not try to achieve statistical optimality.
-Statistical and Bayesian Perspectives
-The GAN has been studied from many statistical or Bayesian viewpoints.
-Bayesian GANs (Saatchi and Wilson,, 2017; Ye and Zhu,, 2018) introduce uncertainty quantification by estimating posterior distribution.
-Others, such as (Zhang et al.,, 2021; Arora et al.,, 2017), analyze GAN training through the lens of generalization and statistical distances.
-Our paper adds to this line of study by describing the GAN discriminator explicitly as a model-free Bayes-aligned classifier.
-In doing so, we provide a Bayesian perspective for adversarial learning and introduce a learning objective function, the bolt loss, for learning a representation that minimize the classification error.
+A GAN trains a generator and a discriminator via a min-max game (Goodfellow et al., 2014).
+The original GAN training objective corresponds to optimizing the Jensen-Shannon (JS) divergence between the generative and data distributions.
+This basic approach is known to suffer from gradient saturation and instability, especially when the model and data supports have limited overlap (Goodfellow et al., 2014; Arjovsky et al., 2017).
+Alternatives replace JS with other metrics and discrepancies, e.g.
+$f$-divergence (Nowozin et al., 2016), least-squares losses (Mao et al., 2017), and auto-encoding reconstruction-loss (Berthelot et al., 2017).
+Among these designs, Wasserstein GAN (WGAN) shows substantial training stability (Arjovsky et al., 2017).
+WGAN replaces JS divergence with the Wasserstein-1 distance $W_{1}$, yielding a smoother training signal and improved stability.
+Stable GAN Training
+A central component in WGAN training is enforcing the 1-Lipschitz constraint on the discriminator function.
+This is commonly addressed via gradient penalty (Gulrajani et al., 2017) or spectral normalization (Miyato et al., 2018).
+Other stabilization approaches include training heuristics such as progressive growing (Karras et al., 2018).
+Bayesian GAN formulations place priors over network parameters and sample from posterior distributions to quantify uncertainty (Saatchi and Wilson, 2017; Ye and Zhu, 2018).
+Several studies analyze generalization and equilibrium behavior via statistical distances for learning dynamics (Arora et al., 2017; Zhang et al., 2021).
+Our work complements these lines of work by directly training the discriminator for Bayes optimal discrimination via the BOLT loss.
 1.2 Main Contributions
-The key contributions of this work are as follows: ($i$) We introduce BOLT-GAN, an adversarial framework in which the discriminator is trained with the BOLT loss.
-The generator is implicitly trained to maximize the Bayes error of discrimination task.
-The inner maximization identifies the Bayes-aligned discriminator with the largest real-fake gap, and the outer minimization drives the generator to reduce this gap, effectively pushing the Bayes error toward random guessing.
-($ii$) We characterize convergence of the BOLT-GAN.
-We show that, under balanced class priors, the BOLT objective is directly linked to total variation (TV) between the data and generator distributions.
-This finding shed lights on the numerical instability of the plain BOLT-GAN training algorithm.
-(iiiiii) To address this instability issue, we propose Lipschitz BOLT-GAN, which enforces a 1-Lipschitz constraint on the discriminator.
-We show that under this constraint, the learning objective is bounded above Wasserstein distance.
-This indicates that this framework guides the generator toward the data distribution using a weaker notion of distance.
-This leads to stabilization of the gradient backpropagation.
-($iv$) We validate our results empirically.
-On CIFAR-10, CelebA-64, and LSUN, Lipschitz BOLT-GAN showing significant improvements.
+The key contributions of this work are as follows.
+(i) Bayes-optimal GAN training via Max-BER: we deploy the BOLT framework to formulate Bayes-optimal GAN training, where the generator is trained to maximize the Bayes error rate of the discrimination task.
+We first show the Bayes optimality of BOLT (Theorem 2) and characterize the bias and variance of the BOLT plug-in estimation of the BER (Theorem 3).
+We then conclude that under the Max-BER game, the generator and data distributions match in total variation (Theorem 4).
+(ii) BOLT-GAN framework: We further show that, when the discriminator is restricted to the class of Lipschitz models, the BOLT-trained GAN minimizes the distance between the generator and data distributions with respect to an alternative IPM that is bounded by the Wasserstein-1 distance (Theorem 5).
+The resulting BOLT-GAN trains the generator toward the true data distribution using a weaker notion of distance, leading to more efficient gradient backpropagation and hence better stability.
+(iii) Empirical validation: We validated BOLT-GAN through experiments with the CIFAR-10, CelebA-64, LSUN Bedroom-64, and LSUN Church-64 datasets.
+We report FID, precision, and recall, and compare against WGAN and HINGE-GAN baselines (Miyato et al., 2018).
 Notation
-Throughout the paper, we adapt the following notation: expectations of of a random variable $X$ with respect to a distribution $P$ is denoted by $\mathbb{E}_{X\sim P}[\cdot]$.
-The distribution $P=P_{C}$ is defined as the conditional data distribution given the class C.
-The set $\left\{1,\ldots,n\right\}$ is abbreviated as $[n]$.
-2 Background and Preliminaries
-We restrict our analysis to a binary classification task with classes $C_{1}$ and $C_{2}$, although extensions to multi-class problems are possible.
-Consider a binary classification task with classes $C_{1}$ and $C_{2}$ occurring with prior probabilities $q_{1}$ and $q_{2}=1-q_{1}$, respectively.
-The data samples are drawn from the space $\mathcal{X}\subseteq\mathbb{R}^{d}$, for some integer $d$, with class-conditional distributions $P_{1}$ and $P_{2}$ for the classes $C_{1}$ and $C_{2}$, respectively.
-For this setting, the Bayes error rate (BER), denoted by $\varepsilon_{\mathrm{Bayes}}$, is defined as the minimum error probability achievable by a measurable classifier $f:\mathcal{X}\mapsto\{C_{1},C_{2}\}$.
-The BER is determined from the data distribution as (Bishop,, 2006)
-$\varepsilon_{\mathrm{Bayes}}$ $=1-\int_{\mathcal{X}}\max_{\lambda}\left[q_{\lambda}P_{\lambda}(x)\right]\,\mathrm{d}x.$ (1)
-and is achieved by the maximum a posteriori (MAP) classifier, which assigns a sample $x$ to the class $\lambda\in\{1,2\}$ that maximizes the joint distribution $q_{\lambda}P_{\lambda}(x)$.
-2.1 BOLT Loss for Classification
-An explicit computation of the BER(1) requires full knowledge of the data distribution, i.e.,  both the class prior $q_{\lambda}$ and the class conditional feature distribution $P_{\lambda}$ , As these are unavailable, the BER must be bounded and/or estimated ; see (Devroye et al.,, 1996; Fukunaga,, 2013; Noshad et al.,, 2019) for some sample studies.
-The recent study by Naeini et al., (2025) develops a new framework to estimate an upper bound on the Bayes error.
-The bound is universal in the sense that it can be computed by sampling from an arbitrary measurable and bounded function, e.g., a neural network.
-It is further tight, as it can achieve the Bayes error when optimized over a set of measurable functions.
-Motivated by these observations, Naeini et al., (2025) introduces the BOLT loss, which utilizes the proposed bound to approximate the optimal Bayes classifier via a parameterized learning model.
-Theorem 1 (Binary Bayes error (Naeini et al.,, 2025)).
-Let $h:\mathcal{X}\to[-1,0]$ be a measurable function. Then, the Bayes error of the binary classification satisfies
-$\varepsilon_{\mathrm{Bayes}}\leq q_{2}+q_{2}\mathbb{E}_{X\sim P_{2}}\left[h(X)\right]-\hskip-1.42262ptq_{1}\mathbb{E}_{X\sim P_{1}}\left[h(X)\right].$ (2)
-Naeini et al., (2025) show that the bound (2) is tight when optimized over $h$.
-This enables a computational means to estimate $\varepsilon_{\mathrm{Bayes}}$ by sampling from output of classifier.
-In particular if $h$ is a parameterized predictor model $h_{\theta}(\cdot)$, e.g., a neural network with weights $\theta$, the bound can be estimated and minimized.
+On a measurable space $(\mathcal{X},\mathcal{F})$, probability measures are shown uppercase, e.g., $P$, and their corresponding densities are denoted lowercase, e.g., $p$.
+Expectation with respect to probability measure $P$ is shown by $\mathbb{E}_{X\sim P}[\cdot]$.
+The set $\{1,\dots,n\}$ is abbreviated as $[n]$.
+2 Preliminaries
+We consider binary classification with classes $C\in\{1,2\}$ and data space $\mathcal{X}\subseteq\mathbb{R}^{d}$.
+Let $q_{i}:=\Pr(C=i)$ denote the class priors ($q_{2}=1-q_{1}$), and let $P_{i}$ denote the class-conditional distribution of $X$ given $C=i$.
+When needed, assume $P_{i}$ admits a density $p_{i}$ with respect to a common dominating measure $\mu$ on $\mathcal{X}$.
+The Bayes error rate (BER), denoted by $\varepsilon_{\rm Bayes}$, is the minimum error probability achievable by any measurable classifier $f:\mathcal{X}\to\{C_{1},C_{2}\}$, which is achieved by the maximum a posteriori (MAP) rule (Bishop, 2006).
+In terms of densities, the BER is
+$\varepsilon_{\rm Bayes}=1-\int_{\mathcal{X}}\max_{i\in\{1,2\}}\left\{q_{i}\,p_{i}(x)\right\}\,d\mu(x),$ (1)
+and the MAP classifier assigns sample $x$ to
+$\widehat{C}(x)\in\operatorname*{arg\,max}_{i\in\{1,2\}}q_{i}\,p_{i}(x).$ (2)
+2.1 BOLT Framework
+With unknown data distribution, the BER can be approximated (Devroye et al., 1996; Fukunaga, 2013; Noshad et al., 2019).
+Naeini et al.
+(2025) have developed an empirical estimate of universal upper bound on BER that can be computed by sampling an arbitrary measurable and bounded function (e.g., a neural network).
+The resulting surrogate objective given by this upper bound is called the BOLT risk.
+The following theorem presents the universal bound for binary classification.
+Theorem 1 (Binary Bayes error (Naeini et al., 2025)).
+Let $h:\mathcal{X}\to[0,1]$ be a measurable function.
+Then, the BER is bounded as  [^1] The result in Naeini et al.
+(2025) is given for $\tilde{h}:\mathcal{X}\to[-1,0]$.
+We represent the result after the reparameterization $h=1+\tilde{h}$.
+$\varepsilon_{\rm Bayes}\leq q_{1}+q_{2}\,\mathbb{E}_{X\sim P_{2}}[h(X)]-q_{1}\,\mathbb{E}_{X\sim P_{1}}[h(X)].$ (3)
+Naeini et al.
+(2025) show that the bound (3) is tight when optimized over $h$.
+This enables a computational means to estimate $\varepsilon_{\rm Bayes}$ by sampling the output of a bounded computational model $h$ with parameters $\theta$.
 This transforms the upper bound in Theorem 1 into a training objective referred to as the Bayes optimal learning threshold (BOLT).
-While Naeini et al., (2025) establish the BOLT loss for a general multi-class problem, in this work we focus on the binary case, relevant to the discrimination component of the GANs.
-Training via BOLT
-Consider a training dataset $\mathcal{D}=\left\{(x_{i},\lambda_{i}):i\in[n]\right\}$, where $\lambda_{i}\in\{1,2\}$ denotes the class index from which $x_{i}$ is drawn.
-Assume that the samples in $\mathcal{D}$ are drawn i.i.d.
-from the data distribution.
-BOLT trains a parameterized model $h_{\theta}:\mathcal{X}\rightarrow[-1,0]$ on the dataset $\mathcal{D}$ to estimate the Bayes error rate by optimizing BER empirical loss over $\theta$.
-Remark 1.
-Note that the output of the model $h_{\theta}$ is restricted to be within the interval [1,0][-1,0].
-This is not a significant restriction, as it corresponds to a simple negation of the softmax output of the classifier.
-For a given sample $(x,\lambda)$$\in\mathcal{D}$, the BOLT loss is defined as
-$\ell_{\rm{BOLT}}(h_{\theta}(x),\lambda)=(-1)^{\lambda}h_{\theta}(x).$ (3)
-To train the model $h_{\theta}$ on $\mathcal{D}$, where the empirical risk
-$\hat{R}(\theta)=\frac{1}{n}\sum_{i\in[n]}\ell_{\text{BOLT}}(h_{\theta}(x_{i}),\lambda_{i}),$ (4)
-This minimization can be done by mini-batch stochastic gradient descent or other standard training method.
-The connection, between the BOLT-trained model and the Bayes error rate, can be understood using the results of Naeini et al., (2025): from Theorem 1, we can write
-$\varepsilon_{\rm{Bayes}}\leq q_{2}+\mathbb{E}\left[\ell_{\rm{BOLT}}(h_{\theta}(X),\lambda)\right],$ (5)
-where the expectation is taken with respect to the data distribution.
-Note that the prior $q_{2}$ is efficiently estimated from the data.
-By universal approximation (Cybenko,, 1989) and the tightness result of Naeini et al., (2025), one can conclude that the trained model in this case efficiently estimates the Bayes error.
+While Naeini et al.
+(2025) establish the BOLT loss for a general multi-class problem, we focus on the binary case relevant to the discrimination component of GANs.
+BOLT loss
+Considering a labeled dataset $\mathcal{D}=\{(x_{i},C_{i}):i\in[n]\}$, where $C_{i}\in\{1,2\}$.
+BOLT trains the model $h_{\theta}:\mathcal{X}\to[0,1]$ by minimizing an empirical surrogate of the bound in Theorem 1.
+This can be seen as an empirical risk minimization problem, where each $(x,C)\in\mathcal{D}$ is a (feature,label) sample and the BOLT loss is given by
+$\ell_{\rm BOLT}(h_{\theta}(x),C)=(-1)^{C}\,h_{\theta}(x).$ (4)
+The bound in Theorem 1 can be directly written in terms of BOLT risk as
+$\varepsilon_{\rm Bayes}\leq q_{1}+\mathbb{E}\bigl[\ell_{\rm BOLT}(h_{\theta}(X),C)\bigr],$ (5)
+where the expectation is over $(X,C)$ generated as $\Pr(C=i)=q_{i}$ and $X\mid\{C=i\}\sim P_{i}$, for $i\in\{1,2\}$.
 2.2 Generative Adversarial Networks
-A GAN consists of two learning models: a generator $g_{\phi}$ that generates data samples from the latent variable [^1] Typically, $Z$ is drawn from a multivariate Gaussian distribution.
-$Z\sim P_{Z}$ to generate samples of an estimator of $\mathcal{X}$, and a discriminator $h_{\theta}$ that distinguishes the generated sample $g_{\phi}(Z)$ from the true data sample $X\sim P_{\rm data}$.
-To keep the notation tractable, we denote the induced distribution, i.e., distribution of $g_{\phi}(Z)$, by $P_{\rm g}$, which is determined by $P_{Z}$ and the generator model $g_{\phi}$.
-From a learning perspective, the discriminator solves a binary classification problem: it classifies the input as $C=$ real or $C=$ fake, corresponding to $\lambda=1$ or $\lambda=2$.
-The conventional approach is to train $h_{\theta}$ such that it estimates the posterior class probability.
-This leads to the vanilla GAN design proposed by Goodfellow et al., (2014).
-An alternative approach is to use the Kantorovich-Rubinstein duality (Villani et al.,, 2009) to implicitly minimize the Wasserstein distance between the data and induced distributions.
-This approach leads to the Wasserstein GAN design (Arjovsky et al.,, 2017).
-We briefly explain these two approaches in the sequel.
+A GAN consists of two functions: a generator $g:\mathcal{Z}\to\mathcal{X}$ that maps a latent variable $Z\sim P_{Z}$ to a sample in data space $\mathcal{X}$, and a discriminator $h:\mathcal{X}\to[0,1]$ that distinguishes generated samples $g(Z)$ from true data samples $X\sim P_{\rm data}$.
+We denote the induced distribution of $g(Z)$ by $P_{\rm g}$.
 Vanilla GAN
-The vanilla training approach, proposed in the initial study by Goodfellow et al., (2014), treats the discrimination as a classic binary classification problem.
-In this sense, it sets $h_{\theta}(x)=P(C=\texttt{real}|x)$ and uses the cross-entropy loss to train it.
-[^2] In practice, $h_{\theta}$ is a network with sigmoid-activated output.
-This approach implicitly implements the maximum likelihood training of the generator (Arjovsky et al.,, 2017).
+The vanilla training approach, proposed in the initial study by Goodfellow et al.
+(2014), treats the discrimination as a classic binary classification problem.
+In this sense, it sets $h(x)=P(C=\texttt{real}|x)$ and uses the cross-entropy loss to train it.
+This approach implicitly implements the maximum likelihood training of the generator (Arjovsky et al., 2017).
 The risk function, i.e., the expected loss used for training, in this case is given by
-$R_{\rm ML}(\phi,\theta)=\mathbb{E}\left[\mathrm{CE}(Y,\lambda)\right],$ (6)
-where $\mathrm{CE}$ denotes the cross-entropy loss function, $Y$ is the output of the discriminator, i.e., $Y=h_{\theta}(X)$, for an input $X\in\mathcal{X}$ to the discriminator, and $\lambda\in\left\{1,2\right\}$ is the label of $X$ with $\lambda=1$ and $\lambda=2$ denoting real and fake inputs respectively.
-Where $\lambda$ is sampled from a prior $\pi=\Pr\left\{\lambda=1\right\}$ for some $\pi\in(0,1)$, [^3] Typically, $\pi=0.5$ is considered.
-and $X$ is sampled conditionally from $P_{1}=P_{\rm data}$ when $\lambda=1$ and $P_{2}=P_{\rm g}$ when $\lambda=2$.
-Noting that $P_{2}=P_{\rm g}$ is the distribution of $g_{\phi}(Z)$ computed from the latent sample $Z\sim P_{Z}$, we can use the definition of cross-entropy to expand the risk (6) as
-$R_{\rm ML}(\phi,\theta)=$ $-\pi\mathbb{E}_{X\sim P_{\rm{data}}}\left[\log h_{\theta}(X)\right]$ (7)
-$-(1-\pi)\mathbb{E}_{Z\sim P_{Z}}\left[\log(1-h_{\theta}(g_{\phi}(Z)))\right].$
-Defining $\mathcal{L}_{\rm ML}(\phi,\theta)=-R_{\rm ML}(\phi,\theta)$, the generator and discriminator are jointly trained by the following min-max game
-$\min_{\phi}\max_{\theta}\mathcal{L}_{\rm ML}(\phi,\theta),$ (8)
-which is equivalent to the worst-case risk minimization, i.e., fooling the best discriminator.
-It is shown that the solution to this min-max game, when solved for any measurable generator $g$ and discriminator $h$, guarantees the convergence of $P_{\rm g}$ to $P_{\rm data}$ as measured by the convergence of the Jensen–Shannon (JS) divergence to zero(Goodfellow et al.,, 2014, Theorem 1).
+$R_{\rm ML}(g,h)=\mathbb{E}\left[\mathrm{CE}(Y,C)\right],$ (6)
+where $\mathrm{CE}$ denotes the cross-entropy loss function, $Y$ is the output of the discriminator, i.e., $Y=h(X)$, for an input $X\in\mathcal{X}$ to the discriminator, and $C\in\left\{1,2\right\}$ is the label of $X$ with $C=1$ and $C=2$ denoting real and fake inputs respectively.
+Where $C$ is sampled from a prior $\pi=\Pr\left\{C=1\right\}$ for some $\pi\in(0,1)$, [^2] Typically, $\pi=0.5$ is considered.
+and $X$ is sampled conditionally from $P_{1}=P_{\rm data}$ and $P_{2}=P_{\rm g}$.
+Noting that $P_{2}=P_{\rm g}$ is the distribution of $g(Z)$ computed from the latent sample $Z\sim P_{Z}$, we can use the definition of cross-entropy to expand the risk in (6) as
+$R_{\rm ML}(g,h)=$ $-\pi\mathbb{E}_{X\sim P_{\rm{data}}}\left[\log h(X)\right]$ (7)
+$-(1-\pi)\mathbb{E}_{Z\sim P_{Z}}\left[\log(1-h(g(Z)))\right].$
+Defining $\mathcal{L}_{\rm ML}(g,h)=-R_{\rm ML}(g,h)$, the generator learns to sample $P_{\rm data}$ by the following min-max game
+$\min_{g}\max_{h}\;\mathcal{L}_{\rm ML}(g,h),$ (8)
+which is equivalent to  fooling the best discriminator.
+It is shown that the solution to this min-max game, when solved for any measurable generator $g$ and discriminator $h$, guarantees the convergence of $P_{\rm g}$ to $P_{\rm data}$ as measured by the convergence of the Jensen–Shannon (JS) divergence to zero (Goodfellow et al., 2014, Theorem 1).
 Wasserstein GAN
-Arjovsky et al., (2017) propose to train the generator-discriminator pair by minimizing the Wasserstein-$1$ distance, also known as Earth Mover (EM) distance, between $P_{\rm{data}}$ and $P_{\rm g}$.
-Using the Kantorovich-Rubinstein duality, it is shown that the EM distance minimization is equivalent to solving a min-max problem with the following objective (Arjovsky et al.,, 2017, Theorem 3)
-$\mathcal{L}_{\rm EM}(\phi,\theta)=$ $\;\mathbb{E}_{X\sim P_{\rm{data}}}[h_{\theta}(X)]$
-$-\mathbb{E}_{Z\sim P_{Z}}[h_{\theta}(g_{\phi}(Z))].$ (9)
-The training is carried out by solving the same min-max game as in (8) with $\mathcal{L}_{\rm EM}(\phi,\theta)$ as the objective.
-While the WGAN still guarantees the convergence of $P_{\rm g}$ to $P_{\rm data}$, the notion of convergence is weaker than vanilla GAN, in the sense that it only minimizes the EM distance between the two distributions instead of an information divergence.
-Arjovsky et al., (2017) show that the WGAN has better training stability, as compared with the vanilla GAN approach.
-Remark 2.
-Note that in the WGAN, the discriminator model $h_{\theta}$ is not restricted to compute probabilities.
-Intuitively, The discriminator in this case is a real-valued critic rather than a probabilistic classifier.
-The model $h_{\theta}$ is however constrained to be 1-Lipschitz.
-In practice, this constraint can be enforced by approaches such as weight clipping, gradient penalty, or spectral normalization.
-See the supplementary file for additional discussion.
-3 BOLT-GAN Framework
-Here we develop the proposed BOLT-GAN framework for training a GAN model.
-This scheme invokes the BOLT loss to train the generator to fool the Bayes optimal discriminator.
-The convergence in BOLT-GAN is in the sense of convergence in total variation, i.e., stronger than vanilla GAN, for a Lipschitz discriminator it converges in a weaker sense than WGAN.
-This implies a trade-off between the strength of convergence and stability.
-3.1 MAP Estimator via BOLT
-To understand the BOLT-GAN framework, we first present a result which relates a BOLT-trained model to the Bayesian optimal classifier: consider $h_{\theta}:\mathcal{X}\rightarrow[-1,0]$ that is trained for a binary classification task using the BOLT loss.
-Theorem 2 shows how an estimator of the Bayesian optimal classifier, i.e., the MAP estimator, is computed from $h_{\theta}$.
-Theorem 2 (BOLT Estimator of MAP Classifier).
-Let $h^{\star}$ be the minimizer of the upper bound in Theorem 1.
-Assume that the model $h_{\theta}$ approximates $h^{\star}$ in the sense that $\|h^{\star}(x)-h_{\theta}(x)\|_{\infty}\leq\epsilon$.
-Then, the plug-in classifier
-$\widehat{C}_{\theta}(x)=\begin{cases}C_{1}&h_{\theta}(x)\geq-0.5,\\[4.0pt] C_{2}&h_{\theta}(x)<-0.5.\end{cases}$ (10)
-is a plug-in estimator of the MAP classifier, i.e.
-$\widehat{C}(x)=\arg\max_{\lambda\in\{1,2\}}q_{\lambda}P_{\lambda}(x)$ (11)
+Arjovsky et al.
+(2017) propose to train the generator–discriminator pair by minimizing the Wasserstein-$1$ distance (Earth Mover distance) between $P_{\rm data}$ and $P_{\rm g}$.
+Using the Kantorovich–Rubinstein duality, this distance can be written as a supremum over $1$-Lipschitz functions; accordingly, the WGAN “discriminator” is a real-valued critic (rather than a probabilistic classifier) and is constrained to be $1$-Lipschitz in theory.
+In practice, this constraint is typically enforced via weight clipping, gradient penalty, or spectral normalization (see the supplementary material for additional discussion).
+This yields the min–max objective (Arjovsky et al., 2017, Theorem 3)
+$\mathcal{L}_{\rm EM}(g,h)=\mathbb{E}_{X\sim P_{\rm data}}[h(X)]-\mathbb{E}_{Z\sim P_{Z}}[h(g(Z))].$ (9)
+Learning proceeds by solving the same min-max game as in (8) with $\mathcal{L}_{\rm EM}(g,h)$ as the objective.
+While WGAN still guarantees convergence of $P_{\rm g}$ to $P_{\rm data}$, the notion of convergence is in Wasserstein distance (a weaker topology than that induced by common information divergences), which often leads to improved training stability compared with vanilla GAN (Arjovsky et al., 2017).
+3 GAN with Maximum BER
+Discriminator training in GANs is inherently a binary classification problem (real versus fake), which motivates using BOLT.
+However, deploying BOLT inside the adversarial game requires understanding how the Bayes-optimal discriminator is estimated from a BOLT-trained model.
+In the sequel, we characterize this connection by showing that a population BOLT minimizer recovers the Bayes optimal discriminator by thresholding.
+We further derive a finite-sample guarantee for a plug-in BER estimator computed by empirical BOLT risk.
+These results enable us to formulate the GAN training via a min-max problem, in which the generator is trained adversarially to increase the discriminator-induced BER.
+We characterize the metric under which $P_{\rm g}$ converges to $P_{\rm data}$.
+3.1 Bayes Optimal Discriminator via BOLT
+Naeini et al.
+(2025) conjectured that a model trained via empirical BOLT risk can estimate the Bayes optimal classifier, and proposed a heuristic plug-in rule based on thresholding the learned score $h$.
+In the sequel, we formalize this connection.
+Theorem 2 shows that any population minimizer of the BOLT risk is Bayes-aligned and that thresholding at $1/2$ recovers the MAP decision rule.
+Theorem 2 (MAP vs. BOLT optimum).
+Consider the binary classification setup in Section 2 and let $P_{X}$ denote the marginal distribution of $X$. Let
+$h^{\star}\in\operatorname*{arg\,min}_{h:\mathcal{X}\to[0,1]}\;\mathbb{E}\bigl[\ell_{\rm BOLT}(h(X),C)\bigr].$ (10)
+and
+$\eta(x):=\Pr\left(C=1\mid X=x\right)$ (11)
+be the likelihood of $x$. Then, for $P_{X}$-a.e. $x$, we have
+$h^{\star}(x)=\begin{cases}1,&\eta(x)>\tfrac{1}{2},\\ 0,&\eta(x)<\tfrac{1}{2},\\ \text{z},&\eta(x)=\tfrac{1}{2}.\end{cases}$ (12)
+for some $z\in[0,1]$. Consequently, the plug-in classifier
+$\widehat{C}(x)=\begin{cases}1,&h^{\star}(x)\geq\tfrac{1}{2},\\ 2,&h^{\star}(x)<\tfrac{1}{2},\end{cases}$ (13)
+coincides $P_{X}$-a.e. with the MAP classifier
+$C_{\rm MAP}(x)\in\operatorname*{arg\,max}_{k\in\{1,2\}}\Pr\{C=k\mid X=x\}.$ (14)
 Proof.
-The proof is given in the supplementary file. ∎
-Remark 3 (BOLT Bias and Variance).
-Let $U(x)$ denote the likelihood ratio $U(x)=P_{1}(x)/P_{2}(x)$ and let $\widehat{U}$ be its plug-in estimator (see the supplementary file for construction details).
-Assume $\mathbb{E}[\,|\widehat{U}(X)-U(X)|\,]\leq\varepsilon_{0}$.
-Then the BOLT plug-in Bayes-error estimator satisfies
-$\bigl|\mathbb{E}\big[\widehat{\varepsilon}_{\mathrm{BOLT}}\big]-\varepsilon_{\mathrm{Bayes}}\bigr|$ $\;\leq\;q_{1}\,\varepsilon_{0}+\mathcal{O}\!\big(M^{-1/2}\big),$ (12)
-$\mathrm{Var}\!\big(\widehat{\varepsilon}_{\mathrm{BOLT}}\big)$ $\;=\;\mathcal{O}\!\big(M^{-1}\big),$ (13)
-where $M$ is the number of samples used in the plug-in estimator. Additional details are given in the supplementary file.
-Theorem 2 gives an intuitive interpretation to the BOLT loss.
-It indicates that using BOLT, the model learns to mimic the Bayesian optimal classifier.
-From the computational point of view, using BOLT leads to gradients that guide the model to approximate the MAP estimator.
-This finding enables us to use the BOLT loss to train the discriminator of a GAN model.
-We illustrate this approach in the sequel.
-BOLT-GAN
-Let $s_{\theta}:\mathcal{X}\to\mathbb{R}$ denote a model which computes a raw score for input $x$, and let $h_{\theta}$ be a negative sigmoid activation of its output, i.e.
-$h_{\theta}(x)=-\sigma(s_{\theta}(x)).$ (14)
-From Theorem 2, one can conclude that in this case, the model $h_{\theta}$ estimates the Bayesian optimal discriminator.
-Following a similar approach as in vanilla GAN, we can consider sampling from data and generator distributions conditionally with with $P_{1}=P_{\rm data}$ and $P_{2}=P_{\rm g}$ using a prior $\pi=P(\lambda=1)$ for some $\pi\in(0,1)$, where $\lambda=1$ and $\lambda=2$ correspond to real and fake samples, respectively.
-Noting that the generator samples are given by $g_{\phi}(Z)$ for latent samples $Z\sim P_{Z}$, the expected BOLT loss of the discriminator is given by
-$R_{\text{BOLT}}(\phi,\theta)=-\mathcal{L}_{\text{BOLT}}(\phi,\theta),$ (15)
-with $\mathcal{L}_{\text{BOLT}}(\phi,\theta)$ being
-$\mathcal{L}_{\text{BOLT}}(\phi,\theta)=\,$ $\pi\mathbb{E}_{X\sim P_{\text{data}}}[h_{\theta}(X)]$
-$(1-\pi)\mathbb{E}_{Z\sim P_{Z}}[h_{\theta}(g_{\phi}(Z))].$ (16)
-Given this risk function, the generator and discriminator are jointly trained by maximizing the minimum loss. This is equivalent to solving the following min-max game:
-$\min_{\phi}\max_{\theta}\mathcal{L}_{\text{BOLT}}(\phi,\theta).$ (17)
-Considering Theorem 1, one can see that the inner optimization minimizes the Bayes error upper bound. [^4] Note that $\mathcal{L}_{\text{BOLT}}=-R_{\text{BOLT}}$.
-Stability of BOLT-GAN
-The learning objective in BOLT-GAN is similar to that of WGAN, with two main differences: ($i$) BOLT-GAN considers an arbitrary sampling prior $\pi$, as in vanilla GAN, and ($ii$) it does not impose any Lipschitz continuity restriction on the discriminator model.
-The discussions in (Arjovsky et al.,, 2017; Gulrajani et al.,, 2017; Miyato et al.,, 2018) suggest that the latter difference can lead to lack of stability, when we use BOLT-GAN in this vanilla form.
-Our numerical experiments confirm the validity of this conjecture.
-To understand the reason behind this behavior, we study the convergence of the vanilla BOLT-GAN in the sequel.
-Our result indicates that BOLT-GAN implicitly minimizes the total variation (TV), which is a strict notion of distance.
-This finding enables us to modify the BOLT-GAN learning objective to a stable form.
-3.2 Convergence of BOLT-GAN
-Consider the general learning objective
-$\mathcal{L}_{\text{BG}}^{(\pi)}(g,h)=\,$ $\pi\mathbb{E}_{X\sim P_{\text{data}}}[h(X)]$ (18)
-$-(1-\pi)\mathbb{E}_{Z\sim P_{Z}}[h(g(Z))],$
-which is defined for any measurable $h:\mathcal{X}\mapsto[0,1]$ and measurable $g:\mathcal{Z}\mapsto\mathcal{X}$ that maps that latent space $\mathcal{Z}$ to the data space.
-$\mathcal{L}_{\text{BG}}^{(\pi)}$ reduces to $\mathcal{L}_{\text{BOLT}}$, when we replace $g$ and $h$ are parameterized models $g_{\theta}$ and $h_{\theta}$.
-Our goal is to understand the metric, under which the generator distribution $P_{\rm g}$ converges to $P_{\rm data}$, when the min-max game with objective $\mathcal{L}_{\text{BG}}^{(\pi)}$ is solved.
+See Appendix B.∎
+Theorem 2 further enables us to express the MAP classifier in terms of the likelihood ratio and rewrite the Bayes error as a hinge expectation.
+Corollary 1 (Optimal BOLT function and hinge form of $\varepsilon_{\rm Bayes}$).
+Assume $q_{1},q_{2}>0$ and that $P_{1},P_{2}$ admit densities $p_{1},p_{2}$ w.r.t.
+a common dominating measure.
+Define the likelihood ratio $U(x):=p_{1}(x)/p_{2}(x)$ on $\{p_{2}(x)>0\}$ and the prior threshold $\tau:=q_{2}/q_{1}$.
+Then a pointwise optimizer of the BOLT functional over $h:\mathcal{X}\to[0,1]$ is
+$h^{\star}(x)\;=\;\begin{cases}0,&U(x)<\tau,\\[2.0pt] 1,&U(x)\geq\tau,\end{cases}$ (15)
+and, defining the hinge map $t_{0}(u):=[\,q_{2}-q_{1}u\,]_{+}$, one has
+$\mathcal{L}_{\mathrm{BOLT}}(h^{\star})=\varepsilon_{\mathrm{Bayes}}\;=\;q_{2}-\mathbb{E}_{X\sim P_{2}}\!\big[t_{0}(U(X))\big].$ (16)
+Proof.
+For $x$ with $p_{2}(x)>0$, Bayes’ rule gives
+$\eta(x):=\Pr(C=1\mid X=x)=\frac{q_{1}U(x)}{q_{1}U(x)+q_{2}}.$ (17)
+Hence, $\eta(x)\gtrless\tfrac{1}{2}$ iff $U(x)\gtrless\tau$ with $\tau:=q_{2}/q_{1}$ .
+Substituting this relation into Theorem 2 yields (15).
+The hinge representation (16) follows by rewriting the tight BOLT bound at $h^{\star}$ in terms of $U$.
+∎
+Corollary 1 shows that $\varepsilon_{\rm Bayes}$ can be written as an expectation of a hinge of $U$.
+In practice, a discriminator trained via empirical BOLT risk can be used to form a finite-sample likelihood-ratio estimate $\widehat{U}$ (see Appendix B).
+Plugging $\widehat{U}$ into Corollary 1, we can estimate the MAP classifier and BER.
+The following result characterizes the bias and variance of the BER estimator.
+Theorem 3 (BER estimator).
+Let $\widehat{U}:\mathcal{X}\to\mathbb{R}$ be an estimator of the likelihood ratio $U$, and let $\widehat{q}_{1},\widehat{q}_{2}\in(0,1)$ be empirical class priors computed on M samples.
+Define $\widehat{\varepsilon}_{\rm BOLT}$ as the plug-in estimator obtained from the hinge representation in Corollary 1 by sample averaging with $M$ samples.
+Assume $\mathbb{E}_{X\sim P_{2}}\!\big[\,|\widehat{U}(X)-U(X)|\,\big]\leq\varepsilon_{0}$.
+Then, we have
+$\big|\mathbb{E}[\widehat{\varepsilon}_{\rm BOLT}]-\varepsilon_{\rm Bayes}\big|\leq q_{1}\,\varepsilon_{0}+O(M^{-1/2}),$ (18)
+and
+$\mathrm{Var}(\widehat{\varepsilon}_{\rm BOLT})=O(M^{-1}).$ (19)
+Proof.
+See Appendix B.∎
+3.2 Generation with Maximum BER
+The results of Section 3.1 allow us to interpret a BOLT-trained discriminator as an estimator of the Bayes-optimal classifier for the real-fake classification task, and to view its objective as a tight surrogate for the discrimination BER.
+We therefore construct an adversarial objective where, for a fixed generator $g$, the discriminator $h$ is trained via the (population) BOLT risk under prior $\pi$, and the generator is trained to maximize the induced BER (equivalently, minimize the discriminator best achievable performance).
+Intuitively, this pushes $P_{\rm g}$ toward $P_{\rm data}$ by making real and fake samples maximally confusable.
+We characterize this behavior by showing that the resulting maximum BER problem minimizes the TV between $P_{\rm data}$ and $P_{\rm g}$.
+Maximum BER objective
+Consider a GAN with generator $g$ and a bounded discriminator $h:\mathcal{X}\to[0,1]$.
+Let $P_{\rm g}$ denote the distribution of $g(Z)$ for $Z\sim P_{Z}$, and define $\pi=\Pr\{C=1\}$, i.e.
+probability of sampling a real data-point.
+We train $h$ by the empirical BOLT risk, and train $g$ adversarially to maximize the induced discrimination BER.
+Following the BOLT formulation, the discrimination BER is estimated (up to a constant) via
+$\max_{h}\;\mathcal{L}_{\rm BG}^{(\pi)}(g,h),$ (20)
+where $\mathcal{L}_{\rm BG}^{(\pi)}(g,h)$ is defined as [^3] Note that $\mathcal{L}_{\rm BG}^{(\pi)}(g,h)$ is the negative of the (population) BOLT risk, up to the additive constant $\pi$.
+$\mathcal{L}_{\rm BG}^{(\pi)}(g,h)\;:=\;$ $\pi\,\mathbb{E}_{X\sim P_{\rm data}}\!\big[h(X)\big]$
+$-\,(1-\pi)\,\mathbb{E}_{Z\sim P_{Z}}\!\big[h(g(Z))\big].$ (21)
+By Theorem 2, thresholding the BOLT-optimal discriminator yields the Bayes optimal discrimination rule.
+Accordingly, the generator can adversarially fool the Bayes optimal discrimination by solving the Max-BER problem that is given by
+$\min_{g}\max_{h}\;\mathcal{L}_{\text{BG}}^{(\pi)}(g,h).$ (22)
+Generator distribution
+Intuitively, the Max-BER problem leads to learning data distribution.
+We next show this and characterize the metric, under which the generator distribution $P_{\rm g}$ converges to $P_{\rm data}$.
 To this end, let us first define total variation (TV).
 Definition 1 (Total variation norm).
-The total variation between two distributions $P$ and $Q$ defined on a measurable $\mathcal{X}$ is defined as
-$TV(P,Q)=\sup_{A\subseteq\mathcal{X}}|P(A)-Q(A)|,$ (19)
-with the supremum taken over all events $A$ in the $\sigma$-field over $\mathcal{X}$.
-The convergence of TV to zero implies for convergence of other distances, such as including JS and EM. The following lemma gives an alternative representation for TV.
-Lemma 1 (Dudley, 2002a ).
+The total variation between two distributions $P$ and $Q$ defined on a measurable space (X, B) is defined as
+$\mathrm{TV}(P,Q)=\sup_{A\in\cal{B}}|P(A)-Q(A)|,$ (23)
+where $\mathcal{B}$ is the $\sigma$-algebra on $\mathcal{X}$.
+In the sequel, we use the standard variational representation of TV given in the following lemma.
+Lemma 1 (Dudley (2002a)).
 The total variation between $P$ and $Q$ is equivalently given by
-$TV(P,Q)=\sup_{f:\mathcal{X}\to[0,1]}\int f\,\mathrm{d}(P-Q),$ (20)
-where the difference PQP-Q is a signed measure on $\mathcal{X}$.
+$\mathrm{TV}(P,Q)=\sup_{f:\mathcal{X}\to[0,1]}\int f\,\mathrm{d}(P-Q),$ (24)
+where the difference PQP-Q is a signed measure on $\mathcal{B}$.
 The following theorem characterizes the connection between the BOLT-GAN learning objective $\mathcal{L}_{\text{BG}}^{(\pi)}$ and the TV between the generator and the data distributions.
 Theorem 4 (BOLT vs TV).
-Fix the prior $\pi\in(0,1)$. For generator $g$, let $P_{\rm g}$ denote the distribution of $g(Z)$, when $Z\sim P_{Z}$. Define $D^{(\pi)}$ as
-$D^{(\pi)}(g)$ $\triangleq\max_{h}\mathcal{L}_{\text{BG}}^{(\pi)}(g,h),$ (21)
-with the maximum taken over all discriminators, i.e., all $h:\mathcal{X}\mapsto[0,1]$. Then, $D^{(\pi)}$ satisfies the following property
-$D^{(\pi)}(g)+D^{(1-\pi)}(g)$ $\;\geq\;\operatorname{TV}(P_{\mathrm{data}},P_{\rm g}).$ (22)
-with equality when $\pi=0.5$.
+Fix the prior $\pi\in(0,1)$. For generator $g$, let $P_{\rm g}$ denote the distribution of $g(Z)$, when $Z\sim P_{Z}$. Define $\mathcal{D}^{(\pi)}$ as
+$\mathcal{D}^{(\pi)}(g)$ $\triangleq\sup_{h}\mathcal{L}_{\text{BG}}^{(\pi)}(g,h),$ (25)
+with the maximum taken over all bounded discriminators, i.e., all $h:\mathcal{X}\mapsto[0,1]$. Then, $\mathcal{D}^{(\pi)}$ satisfies
+$\mathcal{D}^{(\pi)}(g)+\mathcal{D}^{(1-\pi)}(g)$ $\;\geq\;\operatorname{TV}(P_{\mathrm{data}},P_{\rm g}).$ (26)
+with equality holding when $\pi=0.5$.
 Proof.
-From the definition (18), we have
+From the definition (3.2), we have
 $\mathcal{L}_{\text{BG}}^{(\pi)}(g,h)+\mathcal{L}_{\text{BG}}^{(1-\pi)}(g,h)=$ $\mathbb{E}_{X\sim P_{\text{data}}}[h(X)]$
-$-\mathbb{E}_{Z\sim P_{Z}}[h(g(Z))],$ (23)
-which noting that $g(Z)$ is distributed as $P_{\rm g}$ can be written as
+$-\mathbb{E}_{Z\sim P_{Z}}[h(g(Z))].$ (27)
+Noting that $g(Z)$ has distribution $P_{\rm g}$ , we can write
 $\mathcal{L}_{\text{BG}}^{(\pi)}(g,h)+\mathcal{L}_{\text{BG}}^{(1-\pi)}(g,h)=$ $\mathbb{E}_{X\sim P_{\text{data}}}[h(X)]$
 $-\mathbb{E}_{X\sim P_{\rm g}}[h(X)].$ (28)
-Setting the signed metric $P_{\text{data}}-P_{\rm g}$ to be the difference of densities, we can conclude that
-$\Sigma(g)$ $\triangleq\max_{h:\mathcal{X}\mapsto[0,1]}\left[\mathcal{L}_{\text{BG}}^{(\pi)}(g,h)+\mathcal{L}_{\text{BG}}^{(1-\pi)}(g,h)\right]$ (25)
-$=\max_{h:\mathcal{X}\mapsto[0,1]}\left[\int h(X)\mathrm{d}(P_{\text{data}}-P_{\rm g})\right]$ (26)
+Then
+$\Sigma(g)$ $\triangleq\sup_{h:\mathcal{X}\mapsto[0,1]}\left[\mathcal{L}_{\text{BG}}^{(\pi)}(g,h)+\mathcal{L}_{\text{BG}}^{(1-\pi)}(g,h)\right]$ (29)
+$=\sup_{h:\mathcal{X}\mapsto[0,1]}\left[\int h(X)\mathrm{d}(P_{\text{data}}-P_{\rm g})\right]$ (30)
 $\stackrel{{\scriptstyle\text{(a)}}}{{=}}\mathrm{TV}(P_{\rm data},P_{\rm g}),$ (31)
 where (a) follows from Lemma 1. We next use the fact that
-$\Sigma(g)\leq\max_{h}\left[\mathcal{L}_{\text{BG}}^{(\pi)}(g,h)\right]+\max_{h}\left[\mathcal{L}_{\text{BG}}^{(1-\pi)}(g,h)\right],$ (28)
-and using the definition of $D^{(\pi)}$ to conclude that
+$\Sigma(g)\leq\sup_{h}\left[\mathcal{L}_{\text{BG}}^{(\pi)}(g,h)\right]+\sup_{h}\left[\mathcal{L}_{\text{BG}}^{(1-\pi)}(g,h)\right],$ (32)
+and the definition of $\mathcal{D}^{(\pi)}$ to conclude that
 $\mathrm{TV}(P_{\rm data},P_{\rm g})$ $=\Sigma(g)$ (33)
-$\leq D^{(\pi)}(g)+D^{(1-\pi)}(g)$ (30)
-which proves the inequality. We further note that with $\pi=0.5$
+$\leq\mathcal{D}^{(\pi)}(g)+\mathcal{D}^{(1-\pi)}(g).$ (34)
+This proves the inequality. We next note that at $\pi=0.5$
 $\mathrm{TV}(P_{\rm data},P_{\rm g})$ $=\Sigma(g)$
-$=\max_{h}\left[\mathcal{L}_{\text{BG}}^{(0.5)}(g,h)+\mathcal{L}_{\text{BG}}^{(0.5)}(g,h)\right]$
-$=2D^{(0.5)}(g)$ (31)
-which indicates that equality is achieved with $\pi=0.5$. ∎
-The above result indicates that using BOLT-GAN, the generator distribution converges to the data distribution at least as strong as minimizing the TV between the two distributions.
+$=\sup_{h}\left[\mathcal{L}_{\text{BG}}^{(0.5)}(g,h)+\mathcal{L}_{\text{BG}}^{(0.5)}(g,h)\right]$
+$=2\mathcal{D}^{(0.5)}(g).$ (35)
+This concludes the proof. ∎
+Theorem 4 indicates that the solution to the Max-BER problem leads to $P_{\rm g}$ converging $P_{\rm data}$ in TV.
+In particular, under the balanced prior $\pi=0.5$, we obtain the exact identity $\mathrm{TV}(P_{\rm data},P_{\rm g})=2\mathcal{D}^{(0.5)}(g)$, so the Max-BER problem is equivalent to minimizing TV.
+For general $\pi\in(0,1)$, Theorem 4 relates TV to the pair of games specified by $\mathcal{D}^{(\pi)}(g)$ and $\mathcal{D}^{(1-\pi)}(g)$, and hence the convergence in TV is guaranteed when the worst-case objective converges to zero.
 Corollary 2 (Half–TV guarantee).
-Let $D^{(\pi)}$ be as in Theorem 3. Assume for a given generator $g$, we have
-$\max\left\{D^{(\pi)}(g),D^{(1-\pi)}(g)\right\}\leq\varepsilon$ (32)
+Let $\mathcal{D}^{(\pi)}$ be defined as in Theorem 4. Assume for a given generator $g$, we have
+$\max\left\{\mathcal{D}^{(\pi)}(g),\mathcal{D}^{(1-\pi)}(g)\right\}\leq\varepsilon.$ (36)
 Then, the TV between the generator and data distributions is bounded as
@@
 Using Theorem 4, we have
-$2\max\left\{D^{(\pi)}(g),D^{(1-\pi)}(g)\right\}$ $\geq D^{(\pi)}(g)+D^{(1-\pi)}(g)$ (34)
+$2\max\left\{\mathcal{D}^{(\pi)}(g),\mathcal{D}^{(1-\pi)}(g)\right\}$ $\geq\mathcal{D}^{(\pi)}(g)+\mathcal{D}^{(1-\pi)}(g)$
 $\;\geq\;\operatorname{TV}(P_{\mathrm{data}},P_{\rm g}).$ (38)
+This concludes the proof. ∎
+4 BOLT-GAN Framework
+With an unregularized bounded-discriminator, i.e., the Max-BER setting, the optimal discriminator may be highly non-smooth, and the training of a parameterized model can be unstable.
+Theorem 4 shows that under balanced priors this problem is directly linked to TV, a strong discrepancy that can yield poorly behaved gradients in adversarial optimization.
+For stable training, we propose BOLT-GAN, which solves the Max-BER problem while restricting the discriminator to be 1-Lipschitz.
+More precisely, BOLT-GAN solves the following constrained Max-BER problem
+$\min_{g}\max_{h\in\mathcal{H}_{\rm Lip}}\;\mathcal{L}^{(\pi)}_{\rm BG}(g,h),$ (39)
+where $\mathcal{H}_{\mathrm{Lip}}$ denotes the class of bounded 1-Lipschitz discriminators, i.e.,
+$\mathcal{H}_{\mathrm{Lip}}=\{h:\mathcal{X}\mapsto[0,1]:\,$ $|h(x)-h(y)|\leq\|x-y\|$ (40)
+$\forall\,x,y\in\mathcal{X}\}.$
+We next characterize the convergence of BOLT-GAN.
+4.1 BOLT-GAN Generator Distribution
+Theorem 5 shows that the constrained Max-BER problem solved by BOLT-GAN minimizes an induced IPM that is upper-bounded by the Wasserstein-1 distance between $P_{\rm g}$ and $P_{\rm data}$.
+Before stating the main result, let us define the Wasserstein-1 distance.
+Definition 2 (Wasserstein-1 distance).
+Let $(\mathcal{X},\|\cdot\|)$ be a metric space and let $P,Q$ be probability measures on $\mathcal{X}$ with finite first moment. The Wasserstein-1 distance is
+$W_{1}(P,Q)\triangleq\inf_{\gamma\in\Pi(P,Q)}\mathbb{E}_{(X,Y)\sim\gamma}\big[\|X-Y\|\big],$ (41)
+where $\Pi(P,Q)$ is the set of couplings of $P$ and $Q$.
+Computing the Wasserstein-1 distance between two general probability measures is intractable in closed form.
+The Kantorovich-Rubinstein duality result Villani (2009) gives an alternative formulation which enables estimating this distance by sampling.
+Lemma 2 (Kantorovich-Rubinstein duality Villani (2009)).
+Let $P,Q$ be probability measures on $(\mathcal{X},\|\cdot\|)$ with finite first moment. Then
+$W_{1}(P,Q)=\sup_{f:\,\mathrm{Lip}(f)\leq 1}\left\{\mathbb{E}_{X\sim P}[f(X)]-\mathbb{E}_{X\sim Q}[f(X)]\right\}.$
+We next present the Theorem 5, which characterizes the converging behavior of the generator when trained by BOLT-GAN.
+Theorem 5 (BOLT-GAN Convergence).
+Fix $0<\pi\leq 0.5$, for the generator $g$, define
+$\mathcal{D}^{(\pi)}_{\mathrm{Lip}}(g)\triangleq\max_{h\in\mathcal{H}_{\mathrm{Lip}}}\mathcal{L}_{\rm BG}^{(\pi)}(g,h),$ (42)
+with $\mathcal{L}_{\rm BG}^{(\pi)}(g,h)$ defined in (3.2). Then, for any prior $0<\pi\leq 0.5$, we have
+$\mathcal{D}^{(\pi)}_{\mathrm{Lip}}(g)\leq W_{1}\big(P_{\mathrm{data}},P_{\rm g}\big).$ (43)
+where $W_{1}$ denotes the Wasserstein–$1$ distance.
+Proof.
+Starting with (3.2) in the proof of Theorem 4, we can write
+$\Sigma_{\mathrm{Lip}}(g)\triangleq\sup_{h\in\mathcal{H}_{\mathrm{Lip}}}\left[\mathcal{L}_{\text{BG}}^{(\pi)}(g,h)+\mathcal{L}_{\text{BG}}^{(1-\pi)}(g,h)\right]$ (44)
+$=\sup_{h\in\mathcal{H}_{\mathrm{Lip}}}\Bigl\{\mathbb{E}_{X\sim P_{\mathrm{data}}}\!\big[h(X)\big]-\mathbb{E}_{X\sim P_{\rm g}}\!\big[h(X)\big]\Bigr\}.$ (45)
+In the Appendix C, we show that for $0<\pi\leq 0.5$,
+$\mathcal{D}^{(\pi)}_{\mathrm{Lip}}(g)\leq\Sigma_{\mathrm{Lip}}(g).$ (46)
 We can hence write
-$\varepsilon$ $\geq\max\left\{D^{(\pi)}(g),D^{(1-\pi)}(g)\right\}$ (36)
-$\;\geq\;\frac{1}{2}\operatorname{TV}\!\bigl(P_{\mathrm{data}},P_{\rm g}\bigr).$ (37)
-which concludes the proof. ∎
-The above result suggests that the BOLT-GAN objective encourages the generated distribution $P_{g}$ to approach the true data distribution $P_{\mathrm{data}}$ under the total variation (TV) metric.
-3.3 Regularized BOLT-GAN via Lipschitz
-Lipschitz critics are known to stabilize adversarial training and provide informative, non-saturating gradients Arjovsky et al., (2017); Gulrajani et al., (2017); Miyato et al., (2018).
-Motivated by this, we enforce $1$-Lipschitz constraint on the discriminator.
-Intuitively, this can lead to better gradients while training the generator.
-This can reduce gradient explosion or vanishing, as well as the variance of stochastic updates flowing through $\nabla_{\phi}\,h_{\theta}(g_{\phi}(z))$.
-In this section, we show that under this modification, the BOLT-GAN implies a weaker notion of convergence and better stability.
-Lipschitz BOLT-GAN
-The Lipschitz BOLT-GAN trains the generator and discriminator by solving the min-max game (17) , subject to Lipschitz continuity of the discriminator. More precisely, it solves
-$\min_{\phi}\max_{h\in\mathcal{H}_{\mathrm{Lip}}}\mathcal{L}_{\text{BOLT}}(\phi,\theta),$ (38)
-where $\mathcal{H}_{\mathrm{Lip}}$ is the set of functions $h_{\theta}:\mathcal{X}\mapsto[-1,0]$ and
-$\mathcal{H}_{\mathrm{Lip}}=\{h:\mathcal{X}\mapsto[0,1]:\,$ $|h(x)-h(y)|\leq\|x-y\|$ (39)
-$\forall\,x,y\in\mathcal{X}\},$
-The Lipschitz continuity further imposes a smoothness regularization on the BOLT-GAN discriminator.
-For simplicity, we drop the subscript $\theta$ in $h_{\theta}$, stating the following theorem. 4.
-Theorem 4 (Convergence of Lipchitz BOLT-GAN).
-Fix $0<\pi\leq 0.5$, and define the bounded 1-Lipschitz class of discriminator models as
-$\mathcal{H}_{\mathrm{Lip}}=\{h:\mathcal{X}\mapsto[0,1]:\,$ $|h(x)-h(y)|\leq\|x-y\|$ (40)
-$\forall\,x,y\in\mathcal{X}\},$
-for some norm $\|\cdot\|$ defined on data space $\mathcal{X}$. For the generator model $g$, define
-$D^{(\pi)}_{\mathrm{Lip}}(g)\triangleq\max_{h\in\mathcal{H}_{\mathrm{Lip}}}L^{(\pi)}_{\mathrm{BG}}(g,h),$ (41)
-with $L^{(\pi)}_{\mathrm{BG}}(g,h)$ being defined in (18). Then, for any prior $0<\pi\leq 0.5$, we have
-$D^{(\pi)}_{\mathrm{Lip}}(g)\leq W_{1}\big(P_{\mathrm{data}},P_{\rm g}\big).$ (42)
-where $W1$ is the Lipchitz Wasserstein‑1 measure defined in (46) below.
-Proof.
-Starting from (3.2) in the proof of Theorem 3, we can write
-$\Sigma_{\mathrm{Lip}}(g)\triangleq\max_{h\in\mathcal{H}_{\mathrm{Lip}}}\left[\mathcal{L}_{\text{BG}}^{(\pi)}(g,h)+\mathcal{L}_{\text{BG}}^{(1-\pi)}(g,h)\right]$ (43)
-$=\max_{h\in\mathcal{H}_{\mathrm{Lip}}}\Bigl\{\mathbb{E}_{X\sim P_{\mathrm{data}}}\!\big[h(X)\big]-\mathbb{E}_{X\sim P_{\rm g}}\!\big[h(X)\big]\Bigr\}.$ (44)
-In the supplementary file, we show that for $0<\pi\leq 0.5$,
-$D^{(\pi)}_{\mathrm{Lip}}(g)\leq\max_{h\in\mathcal{H}_{\mathrm{Lip}}}\left[\mathcal{L}_{\text{BG}}^{(\pi)}(g,h)+\mathcal{L}_{\text{BG}}^{(1-\pi)}(g,h)\right]$ (45)
-and therefore
-$D^{(\pi)}_{\mathrm{Lip}}(g)\leq\max_{h\in\mathcal{H}_{\mathrm{Lip}}}\Bigl\{\mathbb{E}_{X\sim P_{\mathrm{data}}}\!\big[h(X)\big]-\mathbb{E}_{X\sim P_{\rm g}}\!\big[h(X)\big]\Bigr\}.$
-The Kantorovich-Rubinstein duality (Villani et al.,, 2009) gives
-$W_{1}(P_{\mathrm{data}},P_{g})=$ (46)
-$\sup_{h\in\mathcal{H}_{\mathrm{Lip}}^{+}}\Bigl\{\mathbb{E}_{X\sim P_{\mathrm{data}}}\!\big[h(X)\big]-\mathbb{E}_{X\sim P_{\rm g}}\!\big[h(X)\big]\Bigr\},$
-where $\mathcal{H}_{\mathrm{Lip}}^{+}$ denotes the set of all 1-Lipschitz functions on $\mathcal{X}$
-$\mathcal{H}_{\mathrm{Lip}}^{+}=\{h:|h(x)-h(y)|\leq\|x-y\|\hskip 2.84526pt\forall\,x,y\in\mathcal{X}\}.$ (47)
-Noting that $\mathcal{H}_{\mathrm{Lip}}\subset\mathcal{H}_{\mathrm{Lip}}^{+}$, we can write
+$\mathcal{D}^{(\pi)}_{\mathrm{Lip}}(g)\leq\sup_{h\in\mathcal{H}_{\mathrm{Lip}}}\Bigl\{\mathbb{E}_{X\sim P_{\mathrm{data}}}\!\big[h(X)\big]-\mathbb{E}_{X\sim P_{\rm g}}\!\big[h(X)\big]\Bigr\}.$
+Let $\mathcal{H}_{\mathrm{Lip}}^{+}$ denotes the set of 1-Lipschitz functions, i.e.,
+$\mathcal{H}_{\mathrm{Lip}}^{+}=\{h:|h(x)-h(y)|\leq\|x-y\|\hskip 2.84526pt\forall\,x,y\in\mathcal{X}\}.$
+Noting that $\mathcal{H}_{\mathrm{Lip}}\subset\mathcal{H}_{\mathrm{Lip}}^{+}$, we can use the Kantorovich-Rubinstein duality to write
 $W_{1}(P_{\mathrm{data}},P_{\rm g})$
 $\geq\sup_{h\in\mathcal{H}_{\mathrm{Lip}}}\Bigl\{\mathbb{E}_{X\sim P_{\mathrm{data}}}\!\big[h(X)\big]-\mathbb{E}_{X\sim P_{\rm g}}\!\big[h(X)\big]\Bigr\},$ (47)
-$\geq D^{(\pi)}_{\mathrm{Lip}}(g).$ (49)
-This concludes the proof. ∎
-The above result indicates an interesting dichotomy.
-While the WGAN implicitly minimizes the Earth-Mover (Wasserstein-1) distance between $P_{\mathrm{data}}$ and $P_{g}$, the BOLT-GAN objective can be viewed as implicitly minimizing the total variation (TV) distance.
-The Lipschitz BOLT-GAN, on the other hand, bridges these two perspectives by enforcing Lipschitz continuity, thereby promoting a smoother optimization landscape.
-This intuition may be suggestive of the relative advantages of these three methods.
-Lipschitz enforcement.
-We impose Lipschitzness on $h_{\theta}$ by adding a gradient penalty (GP, $\lambda_{\mathrm{GP}}{=}10$) (Gulrajani et al.,, 2017): for $\hat{x}=\epsilon x+(1-\epsilon)g_{\phi}(z)$ with $\epsilon\!\sim\!\mathcal{U}(0,1)$, we add $\lambda_{\mathrm{GP}}\,\mathbb{E}_{\hat{x}}\!\big(\|\nabla_{\hat{x}}h_{\theta}(\hat{x})\|_{2}-1\big)^{2}$ to the loss.
-Alternatives (spectral normalization, weight clipping) and implementation details are discussed in the supplement.
+$\geq\mathcal{D}^{(\pi)}_{\mathrm{Lip}}(g),$ (48)
+which concludes the proof. ∎
+We note that $\mathcal{D}^{(\pi)}_{\mathrm{Lip}}(g)$ is an IPM between $P_{g}$ and $P_{\rm{data}}$, and it is minimized by BOLT-GAN as indicated in (39).
+Theorem 5 implies that convergence in the Wasserstein-1 distance suffices to guarantee convergence in the BOLT-GAN IPM, while the converse does not necessarily hold.
+Our numerical experiments further show that this weaker distance metric leads to more stable training.
+4.2 Computational Lipschitz Enforcement
+In practice, we parameterize the discriminator as $h_{\theta}(X)=\sigma(f_{\theta}(x))$ for some score function $f_{\theta}(x)$, where $\sigma$ denotes the sigmoid function.
+The 1-Lipschitz constraint is computationally enforced on the score $f_{\theta}$ using a gradient penalty (GP) (Gulrajani et al., 2017): for $\hat{X}=\varepsilon X+(1-\varepsilon)g(Z)$ with $\varepsilon\sim{\rm Unif}(0,1)$, we add
+$\lambda_{\rm GP}\,\mathbb{E}\Bigl(\|\nabla_{\hat{X}}f_{\theta}(\hat{X})\|_{2}-1\Bigr)^{2}$ (49)
+to the discriminator loss for some regularizer $\lambda_{\rm GP}$.
+Since the sigmoid function $\sigma$ is Lipschitz, this enforcement implies that $h$ is also Lipschitz.
+[^4] Indeed, $h$ is Lipschitz with $\mathrm{Lip}(h)\leq\mathrm{Lip}(\sigma)\,\mathrm{Lip}(f_{\theta})$.
 5 Experiments and Results
-To establish the benefits of the lipchitz BOLT-GAN design, we conduct experiments on standard image generation tasks using the MNIST (LeCun et al.,, 1998) dataset.
-and CelebA (Liu et al.,, 2015).
-We use the Fréchet inception distance (FID) (Heusel et al.,, 2017) to assess the quality of the generated images over training epochs and provide comparisons to the WGAN.
+To evaluate the proposed BOLT-GAN, we conduct unconditional image-generation experiments on four standard benchmarks: CIFAR-10, CelebA-64, LSUN Bedroom-64, and LSUN Church-64.
+We compare against (i) WGAN, since it uses the same 1-Lipschitz enforcement mechanism as our implementation, and (ii) a HINGE-GAN, a widely used stability-focused baseline.
+Unless otherwise stated, BOLT-GAN and WGAN use the same generator and discriminator architectures and optimizer settings, i.e., the only difference is the training objective.
+In the experiments, we report the Fréchet inception distance (FID), as well as Improved Precision and Recall (Kynkäänniemi et al., 2019) to quantify sample fidelity and mode coverage.
 5.1 Experimental Setup
-We adopt a residual DCGAN backbone (Gulrajani et al.,, 2017): the generator upsamples a set of 128-D Gaussian latent images to $64{\times}64$ RGB via residual blocks; the critic symmetrically downsamples these images with global-sum pooling and a final linear head for classification.
-BatchNorm/ReLU activation is used in the generator.
-LeakyReLU in the critic, with a Lipschitz constraint enforced through a gradient penalty ($\lambda_{GD}=10$).
-We evaluate performance on CIFAR-10, CelebA-64, and LSUN Bedroom/Church-64.
-For training we use Adam ($\alpha=2{\times}10^{-4},\beta_{1}=0.5,\beta_{2}=0.999$), with batch size 64, for 20 epochs.
-FID (Heusel et al.,, 2017) is computed every 5 epochs on 10,000 samples using torch-fidelity (Seitzer,, 2020), averaged over three seeds.
-Additional architectural and hardware details appear in the supplementary.
-4.2 Quantitative and Convergence Results
-Table 3 shows that BOLT–GAN consistently surpasses WGAN, achieving 10–60% lower FID across all benchmarks.
-As illustrated in Figure 1, it also converges faster and exhibits smoother training dynamics with reduced variance and fewer oscillations.
-Complementary qualitative results in Figure 2 highlight sharper textures and fewer artifacts in samples generated by BOLT–GAN after 100 epochs on CIFAR-10.
-We also trained BOLT–GAN without Lipschitz regularization.
-This TV-based variant quickly diverges (FID >>300–400), confirming that the gradient-penalty term is essential for stability.
-Table 1: FID (↓) after 20 epochs for the non-Lipschitz BOLT–GAN ($\lambda_{\mathrm{GP}}=0$) compared with the Lipschitz version.
-Dataset Non-Lipschitz ($\lambda_{\mathrm{GP}}=0$) Lipschitz ($\lambda_{\mathrm{GP}}=10$)
-CIFAR-10 $\approx$350 (unstable) 44.2 ± 1.2
-CelebA-64 $\approx$300 9.2 ± 0.4
-LSUN Bedroom-64 Diverged (>>400) 40.3 ± 2.3
-4.3 Ablation Studies
 […]
-Table 2: FID (↓) of WGAN-GP and BOLT-GAN-GP on CIFAR-10 across different values of $\lambda_{\mathrm{GP}}$.
-Method $\lambda_{\mathrm{GP}}=1$ 5 10 20
-WGAN-GP 60.8 59.6 60.0 61.1
-BOLT-GAN-GP 44.3 43.6 44.2 46.7
-Table 3: FID ($\downarrow$) after 20 epochs.
-Dataset WGAN BOLT GAN Relative $\Delta$ (%)
-CIFAR-10 $60.0\!\pm\!1.8$ $44.2\!\pm\!1.2$ 26.3-26.3
-CelebA-64 $10.3\!\pm\!0.5$ $9.2\!\pm\!0.4$ 10.7-10.7
-LSUN Bedroom-64 $102.5\!\pm\!2.9$ $40.3\!\pm\!2.3$ 60.7-60.7
-LSUN Church-64 $43.5\!\pm\!1.4$ $14.8\!\pm\!0.6$ 65.9-65.9
-Table 4: Best FID ($\downarrow$) after 20 epochs for different $\lambda_{\text{GP}}$. Bold marks the default used in all other experiments.
-Dataset $\lambda_{\text{GP}}$
-1 5 10 20
-[…]
-CelebA-64 $10.5\!\pm\!0.5$ $9.4\!\pm\!0.4$ $\mathbf{9.2\!\pm\!0.4}$ $10.0\!\pm\!0.5$
-Table 5: CIFAR‑10 FID at 20 ep with balanced vs. imbalanced prior.
-Configuration FID $\Delta$ vs. $\pi=0.50$
-Balanced ($\pi=0.50$) $44.2$
-Imbalanced ($\pi=0.45$) $\mathbf{43.1}$  $2.4\%$
-020204040606080801001005050100100150150EpochFID (lower is better)WGAN (mean & 95% CI)BOLT GAN(mean & 95% CI)
-[…]
+020204040606080801001005050100100150150EpochFID (lower is better)WGAN-GP (mean & 95% CI)BOLT-GAN(mean & 95% CI)
+Figure 1: CIFAR-10 FID vs. epochs (mean and 95% confidence interval over 3 seeds), showing consistently improved convergence of BOLT-GAN relative to WGAN-GP.
 Figure 2: CIFAR-10 samples at 100 epochs.
@@
 […]
+5.2 Quantitative Results
+Figure 1 shows the CIFAR-10 FID trajectory over training, where BOLT-GAN consistently achieves lower FID than WGAN-GP across epochs.
+Table 1 reports FID after 20 epochs.
+BOLT-GAN improves substantially over WGAN across all datasets and is competitive with (and in several cases slightly better than) HINGE-GAN under the same compute budget.
+To quantify mode coverage, Table 2 reports Precision and Recall (higher is better).
+Across datasets where we computed the metric, BOLT-GAN improves both precision and recall relative to WGAN, indicating better quality and coverage.
+Figure 2 shows CIFAR-10 samples after 100 epochs, where BOLT-GAN yields sharper images with fewer artifacts compared to WGAN.
+[…]
+Dataset HINGE-GAN BOLT-GAN WGAN
+CIFAR-10 $6.9\pm 0.8 44$.2 $\pm$ 0.6 $60.0\pm 1.0$
+LSUN Church-64 $19.7\pm 0.5$ 14.8 $\pm$ 0.4 $43.5\pm 0.9$
+LSUN Bedroom-64 $28.8\pm 0.6$ 28.4 $\pm$ 0.5 $102.5\pm 2.0$
+CelebA-64 $19.6\pm 0.5$ 9.2 $\pm$ 0.3 $10.3\pm 0.4$
+Table 2: Improved Precision/Recall ($\uparrow$) with $n=2000$, $k=3$ (higher is better). Best values per dataset/epoch are bolded.
+Dataset Epochs Method Precision Recall
+CIFAR-10 20 BOLT-GAN 0.763 0.352
+20 WGAN 0.714 0.183
+100 BOLT-GAN 0.812 0.469
+100 WGAN 0.764 0.102
+LSUN Church-64 20 BOLT-GAN 0.750 0.411
+20 WGAN 0.681 0.176
+100 BOLT-GAN 0.792 0.502
+100 WGAN 0.731 0.098
+LSUN Bedroom-64 20 BOLT-GAN 0.598 0.450
+20 WGAN 0.388 0.098
+CelebA-64 20 BOLT-GAN 0.840 0.548
+20 WGAN 0.672 0.104
+Table 3 compares Lipschitz discriminator ($\lambda_{\rm GP}=10$) with an unconstrained one ($\lambda_{\rm GP}=0$).
+The results show that removing Lipschitz enforcement, i.e., setting $\lambda_{\rm{GP}}=0$, leads to severe numerical instability, while training via a Lipschitz-constrained discriminator remains stable.
+Table 3: FID ($\downarrow$) after 20 epochs for Max-BER (unregularized; $\lambda_{\rm GP}=0$) compared with the proposed BOLT-GAN (Lipschitz; $\lambda_{\rm GP}=10$).
+Dataset   Max-BER BOLT-GAN
+CIFAR-10 $352\pm 48$ $44.2\pm 1.2$
+CelebA-64 $276\pm 35$ $9.2\pm 0.4$
+LSUN Church-64 $468\pm 56$ $31.6\pm 0.7$
+5.3 Ablation Studies
+Label smoothing can be accomplished in BOLT-GAN by changing the prior $\pi$.
+Table 4 shows the FID at 20 and 100 epochs for various choices of prior $\pi$.
+As the results show, a modest prior imbalance can slightly improve FID on CIFAR-10.
+[…]
+Prior $\pi$ FID @ 20 epochs FID @ 100 epochs
+$0.50$ $44.2\pm 0.6$ $36.6\pm 0.8$
+$0.45$ $45.3\pm 0.7$ $\mathbf{36.2\pm 0.7}$
+$0.40$ $48.5\pm 1.0$ $41.0\pm 1.1$
+$0.35$ $51.4\pm 1.3$ $44.8\pm 1.4$
+Table 5: FID ($\downarrow$) of WGAN and BOLT-GAN on CIFAR-10 across different values of $\lambda_{\rm GP}$.
+$\lambda_{\rm GP}$ 1 5 10 20
+WGAN 60.8 59.6 60.0 61.1
+BOLT-GAN 44.3 43.6 44.2 46.7
+We next test the robustness to gradient-penalty strength, we sweep $\lambda_{\rm GP}\in\{1,5,10,20\}$ for 20 epochs and report FID averaged over N seeds; see Table 5.
+The results show that performance stays within roughly $\pm 1$ FID across the range.
+Figure 3 further reports FID trajectories across several GP strengths, showing only mild sensitivity over $\lambda_{\rm GP}\in\{1,5,10,20\}$ and motivating our default choice $\lambda_{\rm GP}=10$.
 6 Conclusion
-This paper has presented BOLT–GAN, a streamlined modification of the WGAN framework inspired by the Bayes optimal learning threshold (BOLT).
-Through extensive experiments on four diverse image‐generation benchmarks, namely CIFAR-10, CelebA-64, LSUN Bedroom-64, and LSUN Church‐64, we have demonstrated that BOLT–GAN consistently produces smoother training dynamics and lowers FID scores by 10-60% compared to the standard WGAN.
-These gains hold across both low‐ and high‐resolution settings, underscoring the robustness and generality of our approach.
-The current study introduces two extreme cases of BOLT-GAN, namely the vanilla and Lipschitz BOLT-GAN.
-While the former minimizes the total variation, the latter restricts a metric that is weaker than the Wasserstein distance.
-Exploring constraints under which BOLT-GAN framework restricts alternative measures remains an interesting direction for future research.
+We proposed BOLT-GAN, a Bayes-error-motivated objective for stable GAN training that trains a bounded 1-Lipschitz discriminator with the BOLT loss.
+The discriminator in this case estimates a Bayes optimal binary classifier for the real-fake classification task, and the generator learns the data distribution adversarially by maximizing the discrimination BER.
+We showed that the Max-BER problem solved by BOLT-GAN minimized an IPM between the generator and data distributions that is upper-bounded by the Wasserstein-1 distance.
+In absence of Lipschitz constraint at the discriminator, the Max-BER problem minimizes the TV between the two distributions, leading to empirical instability.
+Experiments demonstrate that BOLT-GAN improves FID and Improved Precision/Recall over the state-of-the-art frameworks, such as WGAN and HINGE-GAN, under identical architectures and training budgets.
+Understanding how alternative constraints at the discriminator impact the convergence behavior of the Max-BER approach for training is an interesting direction for future research.
 Appendix A Definitions and Preliminaries
@@
 The total variation can be computed by equivalent forms. The following lemma gives two equivalent expressions for it.
-Lemma 2 (Equivalent forms of total variation (Dudley, 2002b, )).
+Lemma 3 (Equivalent forms of total variation (Dudley, 2002b)).
 For probability measures $P$ and $Q$ on $(\mathcal{X},\mathcal{F})$,
@@
 Hence $\mathrm{Lip}(f)\leq 1$ exactly when $f$ is $1$-Lipschitz.
-Lemma 3 (Kantorovich–Rubinstein duality for $W_{1}$(Villani et al.,, 2009)).
+Lemma 4 (Kantorovich–Rubinstein duality for $W_{1}$(Villani and others, 2009)).
 On a Polish metric space $(\mathcal{X},d)$, for probability measures $P,Q$ with finite first moments,
@@
 In general, the total variation is a stronger notion of distance in the probability space. This can be observed from the following result.
-Lemma 4 (Bounding $W_{1}$ by TV on bounded spaces (Villani et al.,, 2009)).
+Lemma 5 (Bounding $W_{1}$ by TV on bounded spaces (Villani and others, 2009)).
 If $(\mathcal{X},d)$ has finite diameter $D\vcentcolon=\sup_{x,y}d(x,y)<\infty$, then for all probability measures $P$ and $Q$,
@@
 Definition 6 (BOLT functional).
-For a measurable $h:\mathcal{X}\to[-1,0]$,
-$\mathcal{L}_{\mathrm{BOLT}}(h)\;=\;q_{2}-\Big(q_{1}\,\mathbb{E}_{X\sim P_{1}}[h(X)]-q_{2}\,\mathbb{E}_{X\sim P_{2}}[h(X)]\Big).$ (57)
-The universal bound of Naeini et al. (2025, Thm. 1) implies that for any such $h$, $\varepsilon_{\mathrm{Bayes}}\leq\mathcal{L}_{\mathrm{BOLT}}(h)$.
-Risk and per-example BOLT loss
-Let $Y\in\{1,2\}$, $X\mid(Y=i)\sim P_{i}$, and let $h:\mathcal{X}\to[-1,0]$ be measurable.
-Define $s(y)\vcentcolon=\mathbbm{1}\{y=2\}-\mathbbm{1}\{y=1\}\in\{-1,+1\}$ (or equivalently, $s(y)=(-1)^{y}$), .
-We further define the per-example BOLT loss and the population risk:
-$\ell_{\mathrm{BOLT}}(z,y)$ $\;\vcentcolon=\;s(y)\,z,$ (58)
-$R(h)$ $\;\vcentcolon=\;\mathbb{E}\!\big[\ell_{\mathrm{BOLT}}(h(X),Y)\big]\;=\;-\,q_{1}\,\mathbb{E}_{X\sim P_{1}}[h(X)]\;+\;q_{2}\,\mathbb{E}_{X\sim P_{2}}[h(X)].$ (59)
-With the BOLT functional from Def. 5,
-$\mathcal{L}_{\mathrm{BOLT}}(h)$ $=q_{2}\;-\;\Big(q_{1}\,\mathbb{E}_{X\sim P_{1}}[h(X)]\;-\;q_{2}\,\mathbb{E}_{X\sim P_{2}}[h(X)]\Big)$ (60)
-$=q_{2}\;+\;R(h),$ (61)
-so the minimizers of $R(h)$ coincide with the minimizers of $\mathcal{L}_{\mathrm{BOLT}}(h)$.
+For a measurable $h:\mathcal{X}\to[0,1]$, define
+$\mathcal{L}_{\mathrm{BOLT}}(h)\;\coloneqq\;q_{1}\;+\;q_{2}\,\mathbb{E}_{X\sim P_{2}}[h(X)]\;-\;q_{1}\,\mathbb{E}_{X\sim P_{1}}[h(X)].$ (57)
+The universal bound implies that for any such $h$,
+$\varepsilon_{\rm Bayes}\;\leq\;\mathcal{L}_{\mathrm{BOLT}}(h).$ (58)
+Risk and per-example BOLT loss.
+Let $Y\in\{1,2\}$, $X\mid(Y=i)\sim P_{i}$, and let $h:\mathcal{X}\to[0,1]$ be measurable. Define
+$s(y)\coloneqq\mathbf{1}\{y=2\}-\mathbf{1}\{y=1\}\in\{-1,+1\},$ (59)
+and the per-example loss and population risk by
+$\ell_{\rm BOLT}(z,y)$ $\coloneqq s(y)\,z,$ (60)
+$R(h)$ $\coloneqq\mathbb{E}\!\left[\ell_{\rm BOLT}(h(X),Y)\right]=-q_{1}\,\mathbb{E}_{X\sim P_{1}}[h(X)]+q_{2}\,\mathbb{E}_{X\sim P_{2}}[h(X)].$ (61)
+With (57),
+$\mathcal{L}_{\mathrm{BOLT}}(h)=q_{1}+R(h),$ (62)
+so minimizers of $R(h)$ coincide with minimizers of $\mathcal{L}_{\mathrm{BOLT}}(h)$.
 Definition 7 (Prior‑weighted GAN/BOLT–GAN functional).
@@
 In this section we restate the result on the BOLT-assisted estimator of MAP classifier given in the main paper and provide a complete proof.
-Theorem 5 (BOLT Estimator of MAP Classifier (Theorem 2 from main paper)).
-Let $(X,Y)$ describe a binary classification problem with $Y\in\{1,2\}$, priors $q_{1}=\Pr\{Y=1\}$ and $q_{2}=\Pr\{Y=2\}$, class‑conditional laws $P_{1}$ and $P_{2}$, and posterior
-$\eta(x)\;\vcentcolon=\;\Pr\{Y=1\mid X=x\}.$ (64)
-Let $h^{\star}\in\arg\min_{h:\,\mathcal{X}\to[-1,0]}\mathcal{L}_{\mathrm{BOLT}}(h)$ (or equivalently, $h^{\star}\in\arg\min R(h)$).
-Then, $h^{\star}$ is pointwise optimal almost surely and takes only the endpoint values:
-$h^{\star}(x)\;=\;\begin{cases}0,&\eta(x)>\tfrac{1}{2},\\[2.0pt] -1,&\eta(x)<\tfrac{1}{2},\\[2.0pt] \text{any }z\in[-1,0],&\eta(x)=\tfrac{1}{2}.\end{cases}$ (65)
-Consequently, the plug‑in classifier at threshold 0.5-0.5, i.e.
-$\widehat{C}(x)\;=\;1\{\,h^{\star}(x)\geq-0.5\,\},$ (66)
-coincides almost everywhere with the MAP decision rule $C_{\mathrm{MAP}}(x)=1\{\eta(x)\geq\tfrac{1}{2}\}$.
-Moreover, if a model $h_{\theta}$ satisfies $\|h_{\theta}-h^{\star}\|_{\infty}\leq\epsilon$, then the plug‑in decision
-$\widehat{C}_{\theta}(x)\;=\;\begin{cases}C_{1},&h_{\theta}(x)\geq-0.5,\\[2.0pt] C_{2},&h_{\theta}(x)<-0.5,\end{cases}$ (67)
-is a plug‑in estimator of the MAP classifier.
+Theorem 6 (BOLT estimator of MAP classifier).
+Let $(X,Y)$ describe a binary classification problem with $Y\in\{1,2\}$, priors $q_{1}=\Pr\{Y=1\}$ and $q_{2}=\Pr\{Y=2\}=1-q_{1}$, class-conditional laws $X\mid(Y=i)\sim P_{i}$, and posterior
+$\eta(x)\coloneqq\Pr\{Y=1\mid X=x\}.$ (65)
+Consider the BOLT upper bound in the $[0,1]$ normalization (as in the main paper): for measurable $h:\mathcal{X}\to[0,1]$ define
+$\mathcal{B}(h)\;\coloneqq\;q_{1}\;+\;q_{2}\,\mathbb{E}_{X\sim P_{2}}[h(X)]\;-\;q_{1}\,\mathbb{E}_{X\sim P_{1}}[h(X)].$ (66)
+Let $h^{\star}\in\arg\min_{h:\mathcal{X}\to[0,1]}\mathcal{B}(h)$. Then $h^{\star}$ is pointwise optimal almost surely and satisfies, for $(P_{1}+P_{2})$-a.e. $x$,
+$h^{\star}(x)=\begin{cases}1,&\eta(x)>\tfrac{1}{2},\\ 0,&\eta(x)<\tfrac{1}{2},\\ \text{any }z\in[0,1],&\eta(x)=\tfrac{1}{2}.\end{cases}$ (67)
+Consequently, the plug-in classifier at threshold $1/2$,
+$\widehat{C}(x)\;=\;\mathbf{1}\!\left\{h^{\star}(x)\geq\tfrac{1}{2}\right\},$ (68)
+coincides almost everywhere with the MAP decision rule $\mathbf{1}\{\eta(x)\geq\tfrac{1}{2}\}$.
+Moreover, if a learned model $h_{\theta}$ satisfies $\|h_{\theta}-h^{\star}\|_{\infty}\leq\varepsilon$ for some $\varepsilon<\tfrac{1}{2}$, then the plug-in decision
+$\widehat{C}_{\theta}(x)\;=\;\mathbf{1}\!\left\{h_{\theta}(x)\geq\tfrac{1}{2}\right\}$ (69)
+agrees with $\widehat{C}(x)$ (and hence with MAP) except possibly on the decision boundary set $\{x:\eta(x)=\tfrac{1}{2}\}$.
 Proof.
-Fix $x\in\mathcal{X}$ and abbreviate $\eta=\eta(x)$. The conditional risk of predicting a value $z\in[-1,0]$ is
-$r_{\eta}(z)$ $\;\vcentcolon=\;\mathbb{E}\!\big[\ell_{\mathrm{BOLT}}(z,Y)\mid X=x\big]\;$ (68)
-$=\;\eta\,\ell_{\mathrm{BOLT}}(z,1)+(1-\eta)\,\ell_{\mathrm{BOLT}}(z,2)=$ (69)
-$\;\eta(-z)+(1-\eta)z\;=\;(1-2\eta)\,z.$ (70)
-To minimize $r_{\eta}(z)$ over $z\in[-1,0]$, we can say
-  1. 1.
-    If $\eta>\tfrac{1}{2}$, then $1-2\eta<0$. Thus, $r_{\eta}(z)$ is strictly decreasing in $z$. The minimum over [1,0][-1,0] in this case is attained at the maximal $z=0$.
-  2. 2.
-    If $\eta<\tfrac{1}{2}$, then $1-2\eta>0$. Thus, $r_{\eta}(z)$ is strictly increasing in $z$. The minimum therefore occurs at the minimal z=1z=-1.
-  3. 3.
-    If $\eta=\tfrac{1}{2}$, then $r_{\eta}(z)\equiv 0$ on [1,0][-1,0], and hence any $z\in[-1,0]$ is optimal.
+The additive constant $q_{1}$ does not affect minimizers, so minimizing $\mathcal{B}(h)$ over $h$ is equivalent to minimizing
+$R(h)\;\coloneqq\;q_{2}\,\mathbb{E}_{X\sim P_{2}}[h(X)]\;-\;q_{1}\,\mathbb{E}_{X\sim P_{1}}[h(X)].$ (70)
+Fix $x\in\mathcal{X}$ and abbreviate $\eta=\eta(x)$. The conditional risk of predicting a value $z\in[0,1]$ is
+$r_{\eta}(z)$ $\coloneqq\mathbb{E}\!\left[\,\ell(z,Y)\mid X=x\,\right]$ (71)
+$=\eta\,\ell(z,1)+(1-\eta)\,\ell(z,2),$
+where (consistent with the main text) $\ell(z,1)=-z$ and $\ell(z,2)=+z$. Hence
+$r_{\eta}(z)$ $=\eta(-z)+(1-\eta)z$
+$=(1-2\eta)\,z.$ (72)
+To minimize $r_{\eta}(z)$ over $z\in[0,1]$:
+  • If $\eta>\tfrac{1}{2}$ then $1-2\eta<0$ and $r_{\eta}(z)$ is decreasing in $z$, so the minimum is at $z=1$.
+  • If $\eta<\tfrac{1}{2}$ then $1-2\eta>0$ and $r_{\eta}(z)$ is increasing in $z$, so the minimum is at $z=0$.
+  • If $\eta=\tfrac{1}{2}$ then $r_{\eta}(z)\equiv 0$ and any $z\in[0,1]$ is optimal.
 This yields the pointwise characterization (67).
-Since $\widehat{C}(x)=1\{h^{\star}(x)\geq-0.5\}$ equals $1\{\eta(x)\geq\tfrac{1}{2}\}$ by (65), the plug‑in rule (66) coincides with MAP almost everywhere.
-The final claim for $h_{\theta}$ follows the fact that $\|h_{\theta}-h^{\star}\|_{\infty}\leq\epsilon$ leaves the decision boundary at 0.5-0.5 unchanged whenever $h^{\star}\in\{-1,0\}$.
+For the plug-in rule, when $\eta(x)\neq\tfrac{1}{2}$ we have $h^{\star}(x)\in\{0,1\}$, and
+$h^{\star}(x)\geq\tfrac{1}{2}\;\Longleftrightarrow\;h^{\star}(x)=1\;\Longleftrightarrow\;\eta(x)>\tfrac{1}{2},$ (73)
+so (68) agrees with MAP almost everywhere (ties only when $\eta(x)=\tfrac{1}{2}$).
+Finally, if $\|h_{\theta}-h^{\star}\|_{\infty}\leq\varepsilon<\tfrac{1}{2}$ and $h^{\star}(x)\in\{0,1\}$, then $h_{\theta}(x)$ cannot cross the threshold $1/2$, so (69) matches (68) away from the boundary set $\{\eta=\tfrac{1}{2}\}$.
 ∎
-Remark 4 (Alternative normalization).
-If one parametrizes $h:\mathcal{X}\to[0,1]$ (e.g., the bounded score after sigmoid activation), the same argument applied to the loss $\tilde{\ell}(z,y)=(-1)^{y}(2z-1)$ yields $h^{\star}(x)=1\{\eta(x)>\tfrac{1}{2}\}$ and a threshold at +0.5+0.5.
-The two normalizations are related by an affine map, which adjusts the threshold accordingly.
-Corollary 2 (Bayes consistency of the plug‑in classifier).
-Assume $\Pr\!\big(\eta(X)=\tfrac{1}{2}\big)=0$ under the data distribution of $(X,Y)$. Let $\widehat{h}_{n}:\mathcal{X}\to[-1,0]$ be risk consistent for $\ell_{\mathrm{BOLT}}$, i.e.
-$R(\widehat{h}_{n})\;\xrightarrow[n\to\infty]{}\;\inf_{h}R(h)\;=\;R(h^{\star}),$ (71)
-with $R$ as in (59). Define $\widehat{C}_{n}(x)=1\{\widehat{h}_{n}(x)\geq-0.5\}$. Then the 011 risk converges to the Bayes risk:
-$\Pr\!\big(\widehat{C}_{n}(X)\neq Y\big)\;\xrightarrow[n\to\infty]{}\;\Pr\!\big(C_{\mathrm{MAP}}(X)\neq Y\big).$ (72)
+Corollary 3 (Bayes consistency of the plug-in classifier).
+Assume
+$\Pr\!\left(\eta(X)=\tfrac{1}{2}\right)=0,$ (74)
+under the distribution of $(X,Y)$, where $\eta(x)=\Pr\{Y=1\mid X=x\}$. Let $\widehat{h}_{n}:\mathcal{X}\to[0,1]$ be risk-consistent for the BOLT loss $\ell(z,1)=-z$ and $\ell(z,2)=+z$, i.e.
+$R(\widehat{h}_{n})\xrightarrow[n\to\infty]{}\inf_{h:\mathcal{X}\to[0,1]}R(h)=R(h^{\star}),$ (75)
+where
+$R(h)\coloneqq\mathbb{E}\!\left[\ell\!\left(h(X),Y\right)\right]=-q_{1}\,\mathbb{E}_{X\sim P_{1}}[h(X)]+q_{2}\,\mathbb{E}_{X\sim P_{2}}[h(X)].$ (76)
+Define the plug-in classifier at threshold $1/2$,
+$\widehat{C}_{n}(x)\coloneqq\mathbf{1}\!\left\{\widehat{h}_{n}(x)\geq\tfrac{1}{2}\right\}.$ (77)
+Then the 011 risk converges to the Bayes risk:
+$\Pr\!\left(\widehat{C}_{n}(X)\neq Y\right)\xrightarrow[n\to\infty]{}\Pr\!\left(C_{\rm MAP}(X)\neq Y\right).$ (78)
 Proof.
-Write the conditional excess risk at $x$ as
-$r_{\eta(x)}\big(\widehat{h}_{n}(x)\big)-r_{\eta(x)}\big(h^{\star}(x)\big)\;=\;|1-2\eta(x)|\cdot\big|\,\widehat{h}_{n}(x)-h^{\star}(x)\,\big|.$ (73)
-Then, by (71), we can write
-$R(\widehat{h}_{n})-R(h^{\star})=\mathbb{E}\!\big[|1-2\eta(X)|\,|\widehat{h}_{n}(X)-h^{\star}(X)|\big]\to 0.$ (74)
-Fix any $\delta\in(0,1)$ and split the misclassification event as follows:
-$\mathcal{E}_{n}\;\vcentcolon=\;\{\widehat{C}_{n}(X)\neq C_{\mathrm{MAP}}(X)\}\;=\;\big(\mathcal{E}_{n}\cap\{|1-2\eta(X)|\geq\delta\}\big)\;\cup\;\{|1-2\eta(X)|<\delta\}.$ (75)
-For the case $|1-2\eta(X)|\geq\delta$, a decision mismatch implies $|\widehat{h}_{n}(X)-h^{\star}(X)|\geq\tfrac{1}{2}$ because $h^{\star}\in\{-1,0\}$ and the plug‑in threshold is 0.5-0.5. Therefore,
-$R(\widehat{h}_{n})-R(h^{\star})$ $\;\geq\;\mathbb{E}\!\big[\,|1-2\eta(X)|\,\tfrac{1}{2}\,\mathbf{1}_{\mathcal{E}_{n}\cap\{|1-2\eta(X)|\geq\delta\}}\,\big]$ (76)
-$\;\geq\;\tfrac{\delta}{2}\,\Pr\!\big(\mathcal{E}_{n}\cap\{|1-2\eta(X)|\geq\delta\}\big).$ (77)
-Letting $n\to\infty$ gives $\Pr(\mathcal{E}_{n}\cap\{|1-2\eta(X)|\geq\delta\})\to 0$.
-Moreover, by assumption, $\Pr(|1-2\eta(X)|<\delta)\to 0$ as $\delta\downarrow 0$.
-We hence first send $n\to\infty$, and then $\delta\downarrow 0$ in (75) to obtain (72).
+Let $h^{\star}$ be a population minimizer from Theorem 6 (your $[0,1]$ version), so $h^{\star}(x)\in\{0,1\}$ for $\eta(x)\neq\tfrac{1}{2}$.
+Fix $x\in\mathcal{X}$ and abbreviate $\eta=\eta(x)$.
+The conditional risk of predicting $z\in[0,1]$ is
+$r_{\eta}(z)\coloneqq\mathbb{E}\!\left[\ell(z,Y)\mid X=x\right]=\eta(-z)+(1-\eta)z=(1-2\eta)\,z.$ (79)
+As in Theorem 6, $z^{\star}\in\arg\min_{z\in[0,1]}r_{\eta}(z)$ satisfies $z^{\star}=1$ if $\eta>\tfrac{1}{2}$, $z^{\star}=0$ if $\eta<\tfrac{1}{2}$, and any $z^{\star}\in[0,1]$ if $\eta=\tfrac{1}{2}$.
+Hence, for any $\widehat{z}\in[0,1]$,
+$r_{\eta}(\widehat{z})-r_{\eta}(z^{\star})=|1-2\eta|\cdot|\widehat{z}-z^{\star}|.$ (80)
+Taking expectation over $X$ gives
+$R(\widehat{h}_{n})-R(h^{\star})=\mathbb{E}\!\left[\,|1-2\eta(X)|\,\big|\widehat{h}_{n}(X)-h^{\star}(X)\big|\,\right]\xrightarrow[n\to\infty]{}0.$ (81)
+Let $E_{n}\coloneqq\{\widehat{C}_{n}(X)\neq C_{\rm MAP}(X)\}$. For any $\delta\in(0,1)$,
+$E_{n}\subseteq\Big(E_{n}\cap\{|1-2\eta(X)|\geq\delta\}\Big)\ \cup\ \{|1-2\eta(X)|<\delta\}.$ (82)
+On the event $\{|1-2\eta(X)|\geq\delta\}$, a decision mismatch (threshold $1/2$) implies $\big|\widehat{h}_{n}(X)-h^{\star}(X)\big|\geq\tfrac{1}{2}$ because $h^{\star}(X)\in\{0,1\}$ there. Therefore,
+$R(\widehat{h}_{n})-R(h^{\star})$ $\geq\mathbb{E}\!\left[\,|1-2\eta(X)|\cdot\big|\widehat{h}_{n}(X)-h^{\star}(X)\big|\cdot\mathbf{1}_{E_{n}\cap\{|1-2\eta(X)|\geq\delta\}}\,\right]$
+$\geq\frac{\delta}{2}\,\Pr\!\left(E_{n}\cap\{|1-2\eta(X)|\geq\delta\}\right).$ (83)
+Sending $n\to\infty$ and using (81) yields
+$\Pr\!\left(E_{n}\cap\{|1-2\eta(X)|\geq\delta\}\right)\xrightarrow[n\to\infty]{}0.$ (84)
+Moreover, by (74), $\Pr(|1-2\eta(X)|<\delta)\to 0$ as $\delta\downarrow 0$.
+Combining this with (82) proves that $\Pr(E_{n})\to 0$, i.e.
+$\widehat{C}_{n}(X)$ agrees with $C_{\rm MAP}(X)$ with probability tending to $1$.
+This implies the 011 risk convergence (78).
 ∎
@@
 Theorem 7 (Optimal bounding function and tightness).
-Let $c(x)\vcentcolon=q_{1}p_{1}(x)-q_{2}p_{2}(x)$. A pointwise maximizer of c(x)h(x)c(x)\,h(x) over $h(x)\in[-1,0]$ is
-$h^{\star}(x)\;=\;\begin{cases}-1,&U(x)<\tau,\\[2.0pt] 0,&U(x)\geq\tau,\end{cases}$ (79)
-and $\mathcal{L}_{\mathrm{BOLT}}(h^{\star})=\varepsilon_{\mathrm{Bayes}}$.
+Assume $q_{1},q_{2}>0$ and that $P_{1},P_{2}$ admit densities $p_{1},p_{2}$ w.r.t. a common dominating measure. Define
+$U(x)\coloneqq\frac{p_{1}(x)}{p_{2}(x)}\quad\text{on }\{p_{2}(x)>0\},\qquad\tau\coloneqq\frac{q_{2}}{q_{1}},$ (86)
+and let
+$c(x)\vcentcolon=q_{1}p_{1}(x)-q_{2}p_{2}(x).$ (87)
+A pointwise maximizer of c(x)h(x)c(x)\,h(x) over $h(x)\in[0,1]$ is
+$h^{\star}(x)\;=\;\begin{cases}0,&U(x)<\tau,\\[2.0pt] 1,&U(x)\geq\tau,\end{cases}$ (88)
+and the BOLT bound is tight:
+$\mathcal{L}_{\mathrm{BOLT}}(h^{\star})=\varepsilon_{\mathrm{Bayes}},$ (89)
+where (consistent with Def. 6 )
+$\mathcal{L}_{\mathrm{BOLT}}(h)\;\coloneqq\;q_{1}\;+\;q_{2}\,\mathbb{E}_{X\sim P_{2}}[h(X)]\;-\;q_{1}\,\mathbb{E}_{X\sim P_{1}}[h(X)].$ (90)
 Proof.
-We start by showing the endpoint optimality.
-For fixed $x$, the map $h\mapsto c(x)h$ is affine on [1,0][-1,0], hence maximized at an endpoint chosen by the sign of $c(x)$, i.e., $h^{\star}(x)=-1$ if $c(x)<0$ and $h^{\star}(x)=0$ otherwise.
-Since $c(x)<0$ holds if and only if $U(x)<\tau$, we can conclude (79).
-We next show the tightness of the result. Using the definition of $\mathcal{L}_{\mathrm{BOLT}}$ (Def. 5, Sec. S1) and substituting $h^{\star}$, we have
-$\mathcal{L}_{\mathrm{BOLT}}(h^{\star})$ $=q_{2}\;-\;\Big(q_{1}\,\mathbb{E}_{X\sim P_{1}}[h^{\star}(X)]-q_{2}\,\mathbb{E}_{X\sim P_{2}}[h^{\star}(X)]\Big)$ (80)
-$=q_{2}\;-\;\int\big(q_{1}p_{1}(x)-q_{2}p_{2}(x)\big)\,h^{\star}(x)\,dx$ (81)
-$=q_{2}\;-\;\int\big(q_{2}p_{2}(x)-q_{1}p_{1}(x)\big)\,1\{U(x)<\tau\}\,dx$ (82)
-$=q_{2}\;-\;\int\big[\,q_{2}p_{2}(x)-q_{1}p_{1}(x)\,\big]_{+}\,dx$ (83)
+Endpoint optimality.
+For fixed $x$, the map $h\mapsto c(x)h$ is affine on $[0,1]$, hence maximized at an endpoint chosen by the sign of $c(x)$: $h^{\star}(x)=1$ if $c(x)\geq 0$ and $h^{\star}(x)=0$ if $c(x)<0$.
+Since $c(x)<0$ holds if and only if $q_{1}p_{1}(x)<q_{2}p_{2}(x)$, i.e.
+$U(x)<\tau$, we obtain (88).
+Tightness. Using the definition (90) and substituting $h^{\star}$, we have
+$\mathcal{L}_{\mathrm{BOLT}}(h^{\star})$ $=q_{1}\;+\;q_{2}\,\mathbb{E}_{X\sim P_{2}}[h^{\star}(X)]\;-\;q_{1}\,\mathbb{E}_{X\sim P_{1}}[h^{\star}(X)]$ (91)
+$=q_{1}\;+\;\int\big(q_{2}p_{2}(x)-q_{1}p_{1}(x)\big)\,h^{\star}(x)\,dx$ (92)
+$=q_{1}\;+\;\int\big(q_{2}p_{2}(x)-q_{1}p_{1}(x)\big)\,\mathbf{1}\{U(x)\geq\tau\}\,dx$ (93)
+$=q_{1}\;-\;\int\big(q_{1}p_{1}(x)-q_{2}p_{2}(x)\big)\,\mathbf{1}\{U(x)\geq\tau\}\,dx$ (94)
+$=q_{1}\;-\;\int\big[\,q_{1}p_{1}(x)-q_{2}p_{2}(x)\,\big]_{+}\,dx$ (95)
+$=q_{1}\;-\;\int\Big(\max\{q_{1}p_{1}(x),q_{2}p_{2}(x)\}-q_{2}p_{2}(x)\Big)\,dx$ (96)
+$=q_{1}\;-\;\int\max\{q_{1}p_{1}(x),q_{2}p_{2}(x)\}\,dx\;+\;\int q_{2}p_{2}(x)\,dx$ (97)
 $=1\;-\;\int\max\{q_{1}p_{1}(x),q_{2}p_{2}(x)\}\,dx\;=\;\varepsilon_{\mathrm{Bayes}}.$ (98)
-Here, we have used the fact that a $1\{a>0\}=[a]_{+} a$nd $q_{2}=\int q_{2}p_{2}$. This completes the proof. ∎
+Here we used [a]+=a $_{+}=a\,\mathbf{1}\{a>0\}$ and the identity $b-a]_{+}=\max\{b,a\}-a$ (with $b=q_{1}p_{1}(x)$, $a=q_{2}p_{2}(x)$), and also $\int q_{2}p_{2}(x)\,dx=q_{2}$.
+This completes the proof.
+∎
 The above result gives an interesting connection between the Bayes error and the likelihood ratio.
-Corollary 3 (Bayes error as a hinge of likelihood ratio).
-With $h^{\star}$ in (79), define the hinge map as
+Corollary 4 (Bayes error as a hinge of the likelihood ratio).
+Define the hinge map
 $t_{0}(u)\;\vcentcolon=\;[\,q_{2}-q_{1}u\,]_{+}.$ (99)
-The Bayes error rate is then given by
+Then the Bayes error rate is given by
 $\varepsilon_{\mathrm{Bayes}}=q_{2}-\mathbb{E}_{X\sim P_{2}}\!\Big[t_{0}\big(U(X)\big)\Big].$ (100)
 Proof.
-Starting from (81) and using $h^{\star}=-1\{U<\tau\}$,
-$q_{1}\mathbb{E}_{P_{1}}[h^{\star}]-q_{2}\mathbb{E}_{P_{2}}[h^{\star}]$ $=\int\!\big(q_{1}p_{1}-q_{2}p_{2}\big)h^{\star}\,dx=\int\!\big(q_{2}p_{2}-q_{1}p_{1}\big)\,1\{U<\tau\}\,dxlabel{eq:S3-swap}$ (87)
-$=\int\!\big[\,q_{2}p_{2}-q_{1}p_{1}\,\big]_{+}\,dx=\mathbb{E}_{X\sim P_{2}}\!\big[\,q_{2}-q_{1}U(X)\,\big]_{+},$ (88)
-where the last equality is obtain from dividing by as $p_{2}(x)$ on $\{p_{2}>0\}$ (the complement contributes 0).
-Plugging into the expression for $\mathcal{L}_{\mathrm{BOLT}}(h^{\star})$, we obtain (86).
-∎
+Recall
+$\varepsilon_{\mathrm{Bayes}}=1-\int\max\{q_{1}p_{1}(x),q_{2}p_{2}(x)\}\,dx=\int\min\{q_{1}p_{1}(x),q_{2}p_{2}(x)\}\,dx.$ (101)
+Using $\min\{a,b\}=b-[\,b-a\,]_{+}$ with $a=q_{1}p_{1}(x)$ and $b=q_{2}p_{2}(x)$, we get
+$\varepsilon_{\mathrm{Bayes}}$ $=\int q_{2}p_{2}(x)\,dx-\int[\,q_{2}p_{2}(x)-q_{1}p_{1}(x)\,]_{+}\,dx$
+$=q_{2}-\int p_{2}(x)\,[\,q_{2}-q_{1}U(x)\,]_{+}\,dx=q_{2}-\mathbb{E}_{X\sim P_{2}}\!\Big[[\,q_{2}-q_{1}U(X)\,]_{+}\Big],$ (102)
+which is (100). ∎
 Plug-in estimator (empirical context).
-Suppose we observe $M_{1}$ samples from class $C_{1}$ and $M_{2}$ from class $C_{2}$ (with total $M=M_{1}+M_{2}$), and set empirical priors $\widehat{q}_{i}\vcentcolon=M_{i}/M$.
+Suppose we observe $M_{1}$ samples from class $C_{1}$ and $M_{2}$ from class $C_{2}$ (with total $M=M_{1}+M_{2}$), and set empirical priors
+$\widehat{q}_{i}\vcentcolon=\frac{M_{i}}{M},\qquad i\in\{1,2\}.$ (103)
 Let $\widehat{U}$ estimate $U$, and let $X^{(2)}_{1},\ldots,X^{(2)}_{M_{2}}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}P_{2}$ be the $M_{2}$ samples of class-$C_{2}$ used to estimate the expectation over $P_{2}$ in (100).
@@
 (104)
-Note $0\leq t_{0}(u)\leq q_{2}\leq 1$ and $0\leq\widehat{t}_{0}(u)\leq\widehat{q}_{2}\leq 1$, so all summands are bounded in $[0,1]$.
+Note that $0\leq t_{0}(u)\leq q_{2}\leq 1$ and $0\leq\widehat{t}_{0}(u)\leq\widehat{q}_{2}\leq 1$, so all summands are bounded in $[0,1]$.
 Lemma 6 (Lipschitz property of the hinge).
@@
 Theorem 8 (Bias and variance of the binary BOLT plug-in estimator).
-Assume i.i.d. draws and bounded outputs in $[0,1]$. Then
-$\big|\,\mathbb{E}[\widehat{\varepsilon}_{\mathrm{BOLT}}]-\varepsilon_{\mathrm{Bayes}}\,\big|$ $\;\leq\;q_{1}\,\mathbb{E}\!\big[\,|\widehat{U}(X)-U(X)|\,\big]\;+\;O(M^{-1/2}),$ (91)
+Assume i.i.d. draws and bounded hinge summands in $[0,1]$. Then
+$\big|\,\mathbb{E}[\widehat{\varepsilon}_{\mathrm{BOLT}}]-\varepsilon_{\mathrm{Bayes}}\,\big|$ $\;\leq\;q_{1}\,\mathbb{E}_{X\sim P_{2}}\!\big[\,|\widehat{U}(X)-U(X)|\,\big]\;+\;O(M^{-1/2}),$ (106)
 $\mathrm{Var}\!\big(\widehat{\varepsilon}_{\mathrm{BOLT}}\big)$ $\;=\;O(M^{-1}).$ (107)
-In particular, if $\mathbb{E}[\,|\widehat{U}(X)-U(X)|\,]\leq\varepsilon_{0}$, then
+In particular, if $\mathbb{E}_{X\sim P_{2}}\!\big[\,|\widehat{U}(X)-U(X)|\,\big]\leq\varepsilon_{0}$, then
 $\big|\,\mathbb{E}[\widehat{\varepsilon}_{\mathrm{BOLT}}]-\varepsilon_{\mathrm{Bayes}}\,\big|\;\leq\;q_{1}\,\varepsilon_{0}\;+\;O(M^{-1/2}).$ (108)
 Proof.
-Bias. By (86)-(89) and adding and subtracting $\mathbb{E}_{X\sim P_{2}}\!\big[t_{0}(\widehat{U}(X))\big]$, we have:
-$\mathbb{E}[\widehat{\varepsilon}_{\mathrm{BOLT}}]-\varepsilon_{\mathrm{Bayes}}=\underbrace{\mathbb{E}_{X\sim P_{2}}\!\big[t_{0}(U(X))-t_{0}(\widehat{U}(X))\big]}_{\text{plug-in error}}+\underbrace{\mathbb{E}\!\Big[t_{0}(\widehat{U}(X))-\tfrac{1}{M_{2}}\sum_{i=1}^{M_{2}}\widehat{t}_{0}\big(\widehat{U}(X_{i})\big)\Big]}_{\text{sampling/priors}}.$ (94)
-By Lemma 5, $t_{0}$ is $q_{1}$-Lipschitz, so $\big|\mathbb{E}_{P_{2}}[t_{0}(U)-t_{0}(\widehat{U})]\big|\leq q_{1}\,\mathbb{E}_{P_{2}}[|\widehat{U}-U|]$.
-For the sampling term, write $Z_{i}=\widehat{t}_{0}(\widehat{U}(X_{i}))\in[0,1]$; Hoeffding gives $\mathbb{E}\big|\tfrac{1}{M_{2}}\sum Z_{i}-\mathbb{E}Z_{1}\big|=O(M_{2}^{-1/2})$.
+Bias. By (104) and adding/subtracting $\mathbb{E}_{X\sim P_{2}}\!\big[t_{0}(\widehat{U}(X))\big]$, we have
+$\mathbb{E}[\widehat{\varepsilon}_{\mathrm{BOLT}}]-\varepsilon_{\mathrm{Bayes}}$ $=\underbrace{\mathbb{E}_{X\sim P_{2}}\!\big[t_{0}(U(X))-t_{0}(\widehat{U}(X))\big]}_{\text{plug-in error}}+\underbrace{\mathbb{E}\!\Bigg[\mathbb{E}_{X\sim P_{2}}\!\big[t_{0}(\widehat{U}(X))\big]-\frac{1}{M_{2}}\sum_{i=1}^{M_{2}}\widehat{t}_{0}\big(\widehat{U}(X^{(2)}_{i})\big)\Bigg]}_{\text{sampling/priors}}.$ (109)
+By Lemma 6, $t_{0}$ is $q_{1}$-Lipschitz, so
+$\Big|\mathbb{E}_{X\sim P_{2}}\!\big[t_{0}(U(X))-t_{0}(\widehat{U}(X))\big]\Big|\leq q_{1}\,\mathbb{E}_{X\sim P_{2}}\!\big[|\widehat{U}(X)-U(X)|\big].$ (110)
+For the sampling/priors term, write $Z_{i}=\widehat{t}_{0}(\widehat{U}(X^{(2)}_{i}))\in[0,1]$.
+Hoeffding implies $\mathbb{E}\big|\tfrac{1}{M_{2}}\sum_{i=1}^{M_{2}}Z_{i}-\mathbb{E}Z_{1}\big|=O(M_{2}^{-1/2})$.
 The empirical priors $\widehat{q}_{i}=M_{i}/M$ add a zero-mean fluctuation with variance $O(M^{-1})$, hence $O(M^{-1/2})$ in expectation.
-With fixed class proportions ($M_{2}=\Theta(M)$), these contributions are of the same order, yielding (91).
+With fixed class proportions ($M_{2}=\Theta(M)$), these contributions yield (106).
 Variance.
 Using $\widehat{\varepsilon}_{\mathrm{BOLT}}=\widehat{q}_{2}-\tfrac{1}{M_{2}}\sum_{i=1}^{M_{2}}Z_{i}$ with bounded $Z_{i}$, $\mathrm{Var}\!\big(\tfrac{1}{M_{2}}\sum_{i}Z_{i}\big)=O(M_{2}^{-1})$ and $\mathrm{Var}(\widehat{q}_{2})=O(M^{-1})$.
-Any covariance between these two bounded averages is $O(M^{-1})$ and is dominated by the same order; hence (92).
+Any covariance between these two bounded averages is $O(M^{-1})$, hence (107).
 ∎
@@
 $\varepsilon_{0}\;=\;O\!\big(W^{-\gamma}+N^{-1/2}\big),$
-where $W$ is a width/capacity proxy, $N$ is the number of training samples used to learn $\widehat{U}$, and $\gamma>0$ depends on the target smoothness and the approximating family (Lu et al.,, 2020; Shalev-Shwartz and Ben-David,, 2014).
+where $W$ is a width/capacity proxy, $N$ is the number of training samples used to learn $\widehat{U}$, and $\gamma>0$ depends on the target smoothness and the approximating family (Lu et al., 2020; Shalev-Shwartz and Ben-David, 2014).
 Appendix C Further Results on BOLT-GAN
@@
 $\mathcal{H}_{\mathrm{Lip}}$ $\;\vcentcolon=\;\big\{\,h:\mathcal{X}\!\to\![0,1]\;:\;|h(x)-h(y)|\leq\|x-y\|\ \forall\,x,y\in\mathcal{X}\,\big\},$ (113)
-$D^{(\pi)}_{\mathrm{Lip}}(g)$ $\;\vcentcolon=\;\sup_{h\in\mathcal{H}_{\mathrm{Lip}}}\mathbb{E}_{X\sim P_{\mathrm{data}}}\!\big[h(X)\big]-(1-\pi)\,\mathbb{E}_{X\sim P_{g}}\!\big[h(X)\big],$ (98)
+$D^{(\pi)}_{\mathrm{Lip}}(g)$ $\;\vcentcolon=\;\pi\sup_{h\in\mathcal{H}_{\mathrm{Lip}}}\mathbb{E}_{X\sim P_{\mathrm{data}}}\!\big[h(X)\big]-(1-\pi)\,\mathbb{E}_{X\sim P_{g}}\!\big[h(X)\big],$ (114)
 $\Sigma_{\mathrm{Lip}}(g)$ $\;\vcentcolon=\;\sup_{h\in\mathcal{H}_{\mathrm{Lip}}}\Big(\mathbb{E}_{X\sim P_{\mathrm{data}}}\!\big[h(X)\big]-\mathbb{E}_{X\sim P_{g}}\!\big[h(X)\big]\Big).$ (115)
@@
 Symmetry for $\pi>\tfrac{1}{2}$.
-The case $\pi>\tfrac{1}{2}$ follows by exchanging labels (replacing $\pi$ with $1-\pi$, i.e., swaping $P_{\mathrm{data}}$ and $P_{g}$).
+The case $\pi>\tfrac{1}{2}$ follows by exchanging labels (replacing $\pi$ with $1-\pi$, i.e., swapping $P_{\mathrm{data}}$ and $P_{g}$).
 Thus the conclusions above extend to all $\pi\in(0,1]$ by symmetry.
@@
 In our code, the gradient penalty is applied to the raw critic output $\tilde{h}_{\theta}$ (pre‑activation), and the bounded score $h_{\theta}=\sigma(\tilde{h}_{\theta})\in[0,1]$ is used inside the BOLT objective, as summarized in Algorithm 1.
-(Gulrajani et al.,, 2017; Miyato et al.,, 2018, Sec.
+(Gulrajani et al., 2017; Miyato et al., 2018, Sec.
 3).
 D.1 Preliminaries on Lipschitz continuity
-Lemma 9 (Basic Lipschitz calculus (Miyato et al.,, 2018)).
+Lemma 10 (Basic Lipschitz calculus (Miyato et al., 2018)).
 Let $f$, $g$, $f_{1}$, and $f_{2}$ be Lipschitz. Then
@@
 $\mathrm{Lip}(f_{1}+f_{2})$ $\leq\mathrm{Lip}(f_{1})+\mathrm{Lip}(f_{2}).$ (119)
-Lemma 10 (Layerwise Lipschitz constants in common architectures (Miyato et al.,, 2018; Anil et al.,, 2019)).
+Lemma 11 (Layerwise Lipschitz constants in common architectures (Miyato et al., 2018; Anil et al., 2019)).
 Assume all vector spaces are equipped with the Euclidean norm $\|\cdot\|_{2}$, and for a function $f$ we write $\mathrm{Lip}(f)$ for its global Lipschitz constant with respect to $\|\cdot\|_{2}$.
@@
 For convolutions, $\sigma(W)$ is the operator norm of the induced linear map and can be estimated by alternating conv/transpose‑conv passes.
-(Miyato et al.,, 2018).
+(Miyato et al., 2018).
 Guarantee.
@@
 This yields strong stability at low computational cost and is widely adopted in GAN critics.
-(Miyato et al.,, 2018).
+(Miyato et al., 2018).
 Residual blocks and skips.
@@
 Gradient penalty encourages $\|\nabla_{x}\tilde{h}_{\theta}(x)\|_{2}\approx 1$ on a chosen sampling distribution.
-The most common choice is the interpolation penalty (Gulrajani et al.,, 2017): draw $(x_{\mathrm{real}},x_{\mathrm{fake}})$, form $\hat{x}=\alpha x_{\mathrm{real}}+(1-\alpha)x_{\mathrm{fake}}$ with $\alpha\sim\mathrm{Unif}[0,1]$, and add
+The most common choice is the interpolation penalty (Gulrajani et al., 2017): draw $(x_{\mathrm{real}},x_{\mathrm{fake}})$, form $\hat{x}=\alpha x_{\mathrm{real}}+(1-\alpha)x_{\mathrm{fake}}$ with $\alpha\sim\mathrm{Unif}[0,1]$, and add
 $\lambda\,\mathbb{E}_{\hat{x}}\!\big(\,\|\nabla_{\hat{x}}\tilde{h}_{\theta}(\hat{x})\|_{2}-1\,\big)^{2}$ (120)
@@
 It does not by itself prove a global $1$-Lipschitz bound, but it aligns with the KR optimal‑critic condition and markedly improves training stability.
-(Gulrajani et al.,, 2017).
+(Gulrajani et al., 2017).
 Implementation notes.
@@
 Method.
-Clip each parameter after every update: $w\leftarrow\mathrm{clip}(w,\,-\mathcal{\mathcal{}}{L}_{\mathrm{wc}},\,\mathcal{\mathcal{}}{L}_{\mathrm{wc}})$ (Arjovsky et al.,, 2017).
+Clip each parameter after every update: $w\leftarrow\mathrm{clip}(w,\,-\mathcal{\mathcal{}}{L}_{\mathrm{wc}},\,\mathcal{\mathcal{}}{L}_{\mathrm{wc}})$ (Arjovsky et al., 2017).
 This loosely bounds per‑layer operator norms via element‑wise bounds.
@@
 In practice we prefer SN or GP.
-(Arjovsky et al.,, 2017; Miyato et al.,, 2018).
+(Arjovsky et al., 2017; Miyato et al., 2018).
 D.5 Other gradient penalties and related methods
   • One‑sided GP (Lipschitz penalty).
-Penalize only $\|\nabla\tilde{h}_{\theta}\|_{2}>1$: $\lambda\,\mathbb{E}[\max(0,\|\nabla\tilde{h}_{\theta}\|-1)^{2}]$, avoiding a push toward <1<1 gradients that can reduce capacity (Petzka et al.,, 2018).
+Penalize only $\|\nabla\tilde{h}_{\theta}\|_{2}>1$: $\lambda\,\mathbb{E}[\max(0,\|\nabla\tilde{h}_{\theta}\|-1)^{2}]$, avoiding a push toward <1<1 gradients that can reduce capacity (Petzka et al., 2018).
   • Zero‑centered penalties $R_{1}/R_{2}$.
-Apply $\lambda\,\mathbb{E}_{x\sim P_{\mathrm{data}}}\|\nabla_{x}\tilde{h}_{\theta}(x)\|_{2}^{2}$ ($R_{1}$) or $\lambda\,\mathbb{E}_{x\sim P_{g}}\|\nabla_{x}\tilde{h}_{\theta}(x)\|_{2}^{2}$ ($R_{2}$) to improve convergence empirically—even though they do not strictly enforce $1$-Lipschitzness (Mescheder et al.,, 2018).
+Apply $\lambda\,\mathbb{E}_{x\sim P_{\mathrm{data}}}\|\nabla_{x}\tilde{h}_{\theta}(x)\|_{2}^{2}$ ($R_{1}$) or $\lambda\,\mathbb{E}_{x\sim P_{g}}\|\nabla_{x}\tilde{h}_{\theta}(x)\|_{2}^{2}$ ($R_{2}$) to improve convergence empirically—even though they do not strictly enforce $1$-Lipschitzness (Mescheder et al., 2018).
   • DRAGAN (local penalties).
-Sample $\hat{x}=x_{\mathrm{real}}+\delta$ with small noise and penalize $\|\nabla_{\hat{x}}\tilde{h}_{\theta}(\hat{x})\|_{2}$ near the data manifold to promote local smoothness (Kodali et al.,, 2017).
-  • Orthogonal/Parseval constraints. Constrain $W^{\top}W\approx I$ (Parseval/Björck) to shrink $\|W\|_{2}$ toward $1$ at higher compute cost (Cissé et al.,, 2017).
+Sample $\hat{x}=x_{\mathrm{real}}+\delta$ with small noise and penalize $\|\nabla_{\hat{x}}\tilde{h}_{\theta}(\hat{x})\|_{2}$ near the data manifold to promote local smoothness (Kodali et al., 2017).
+  • Orthogonal/Parseval constraints. Constrain $W^{\top}W\approx I$ (Parseval/Björck) to shrink $\|W\|_{2}$ toward $1$ at higher compute cost (Cissé et al., 2017).
   • Lipschitz activations.
-Using activations with known Lipschitz constants (e.g., GroupSort) can yield provably $1$-Lipschitz networks when combined with spectral control of linear layers (Anil et al.,, 2019).
+Using activations with known Lipschitz constants (e.g., GroupSort) can yield provably $1$-Lipschitz networks when combined with spectral control of linear layers (Anil et al., 2019).
   • Optimizer‑level gradient clipping. Stabilizes updates but does not impose a function‑level Lipschitz bound; . We emphasize that this is different from SN or GP.
@@
 Appendix E Additional Experimental Results
-E.1 CelebA–64 $\lambda_{\mathrm{GP}}$ Sensitivity for BOLT–GAN
-For completeness, Figure 4 presents the CelebA–64 $\lambda_{\mathrm{GP}}$ sweep mirroring the CIFAR-10 analysis in the main text.
-The stability/performance trends are consistent: $\lambda_{\mathrm{GP}}\!\in\![5,10]$ is best, with variations within $\pm 1$ FID.
+E.1 CelebA–64 $\lambda_{\rm{GP}}$ Sensitivity for BOLT–GAN
+For completeness, Figure 4 presents the CelebA–64 $\lambda_{\rm{GP}}$ sweep mirroring the CIFAR-10 analysis in the main text.
+The stability/performance trends are consistent: $\lambda_{\rm{GP}}\!\in\![5,10]$ is best, with variations within $\pm 1$ FID.
 […]
@@
 E.2 Qualitative Results
-We provide qualitative samples to complement the quantitative results in the main paper: each panel below shows an uncurated $4\times 4$ grid of random images generated by BOLT–GAN–GP after 20 training epochs under the same setup as our primary experiments (same architectures, optimizer, and $\lambda_{\mathrm{GP}}$).
+We provide qualitative samples to complement the quantitative results in the main paper: each panel below shows an uncurated $4\times 4$ grid of random images generated by BOLT–GAN–GP after 20 training epochs under the same setup as our primary experiments (same architectures, optimizer, and $\lambda_{\rm{GP}}$).
 Figure 5: Qualitative samples (first experiment of the main paper).
-BOLT-GAN-GP after 20 epochs with $\lambda_{\mathrm{GP}}{=}10$.
+BOLT-GAN-GP after 20 epochs with $\lambda_{\rm{GP}}{=}10$.
 Top-left: CIFAR-10; top-right: CelebA-64; bottom-left: LSUN Bedroom-64; bottom-right: LSUN Church-64.
@@
 $\mathcal{L}_{\mathrm{BOLT}}(\varphi,\theta)=\pi\,\frac{1}{B}\sum_{j=1}^{B}h_{j}\;-\;(1-\pi)\,\frac{1}{B}\sum_{j=1}^{B}\hat{h}_{j}.$
-For reference, we also evaluate a non‑Lipschitz baseline by setting the gradient‑penalty coefficient to zero $(\lambda_{\mathrm{GP}}=0)$.
-As expected, this variant is numerically unstable and yields high FID; see Sec.
-S5 for details.
 F.5 Lipschitz Enforcement
@@
 Concretely, we sample interpolants $\tilde{x}=\alpha x+(1-\alpha)\hat{x}$ with $\alpha\sim\mathrm{Unif}[0,1]$ and add
-$\mathrm{GP}\;=\;\frac{\lambda_{\mathrm{GP}}}{B}\sum_{j=1}^{B}\Big(\,\|\nabla_{\tilde{x}_{j}}\tilde{h}_{\theta}(\tilde{x}_{j})\|_{2}-1\Big)^{2}$
+$\mathrm{GP}\;=\;\frac{\lambda_{\rm{GP}}}{B}\sum_{j=1}^{B}\Big(\,\|\nabla_{\tilde{x}_{j}}\tilde{h}_{\theta}(\tilde{x}_{j})\|_{2}-1\Big)^{2}$
 to the critic objective.
-We set $\lambda_{\mathrm{GP}}=10$ as our default because this value is the de facto choice introduced by WGAN‑GP and widely adopted in subsequent implementations; it offers a robust trade‑off between stability and capacity in practice (Gulrajani et al.,, 2017; Mescheder et al.,, 2018).
-In Sec.
-S5 we also report a small sweep over $\lambda_{\mathrm{GP}}\in\{1,5,10,20\}$ that confirms the default is near‑optimal under our settings.
+We set $\lambda_{\rm{GP}}=10$ as our default because this value is the de facto choice introduced by WGAN‑GP and widely adopted in subsequent implementations; it offers a robust trade‑off between stability and capacity in practice (Gulrajani et al., 2017; Mescheder et al., 2018).
+In Appendix F we also report a small sweep over $\lambda_{\rm{GP}}\in\{1,5,10,20\}$ that confirms the default is near‑optimal under our settings.
 F.6 Optimization and Training Schedule
@@
 F.8 Hyperparameter Sweeps and Selection
-We sweep $\lambda_{\mathrm{GP}}\in\{1,5,10,20\}$ separately for each method and dataset, and we report the best FID per method within a fixed compute budget.
+We sweep $\lambda_{\rm{GP}}\in\{1,5,10,20\}$ separately for each method and dataset, and we report the best FID per method within a fixed compute budget.
 Other knobs such as $\beta_{1}$ and $n_{\mathrm{d}}$ have a comparatively minor effect in our ranges (e.g., within $\pm 1$ FID on CIFAR‑10).
@@
 We use FID as the primary quantitative measure of image‑generation quality.
-FID computes the Fréchet distance between multivariate Gaussian embeddings of real and generated images in the feature space of a pre‑trained Inception‑V3 network (Heusel et al.,, 2017).
+FID computes the Fréchet distance between multivariate Gaussian embeddings of real and generated images in the feature space of a pre‑trained Inception‑V3 network (Heusel et al., 2017).
 Lower values indicate that the generated distribution is closer to the real distribution in terms of both perceptual and statistical similarity.
-Unless noted otherwise, all FID scores are computed on 50k generated samples using the official Inception statistics for each dataset.
+Unless noted otherwise, all FID scores are computed on 10k generated samples using the official Inception statistics for each dataset.
 Evaluation Protocol.
-Unless stated otherwise, we evaluate every checkpoint with 50k samples and the official Inception statistics.
-Snapshot selection and the choice of $\lambda_{\mathrm{GP}}$ follow the settings detailed in Sec.
-S7.
+Unless stated otherwise, we evaluate every checkpoint with 10k samples and the official Inception statistics.
+Snapshot selection and the choice of $\lambda_{\rm{GP}}$ follow the settings detailed in Appendix F.
 F.10 Logging and Diagnostics
@@
 In particular, we track the empirical distribution of $\|\nabla_{\tilde{x}}\tilde{h}_{\theta}(\tilde{x})\|_{2}$ on the interpolants used for the penalty, the value of the penalty term itself, and the ratio $\mathbb{E}[(\|\nabla\|-1)^{2}]/\mathbb{E}[\text{critic term}]$.
-Healthy runs concentrate gradient norms near $1$; persistent mass well above $1$ suggests increasing $\lambda_{\mathrm{GP}}$ or applying the penalty more frequently, whereas mass well below $1$ suggests reducing $\lambda_{\mathrm{GP}}$ or switching to a one‑sided penalty.
+Healthy runs concentrate gradient norms near $1$; persistent mass well above $1$ suggests increasing $\lambda_{\rm{GP}}$ or applying the penalty more frequently, whereas mass well below $1$ suggests reducing $\lambda_{\rm{GP}}$ or switching to a one‑sided penalty.
 F.11 Determinism and Seeds
@@
 F.12 Failure Modes and Stability Notes
-The non‑Lipschitz BOLT variant ($\lambda_{\mathrm{GP}}=0$) frequently diverges or produces very high FID, which is consistent with optimizing a TV‑strength objective.
+The non‑Lipschitz BOLT variant ($\lambda_{\rm{GP}}=0$) frequently diverges or produces very high FID, which is consistent with optimizing a TV‑strength objective.
 Combining strong spectral normalization with a large gradient‑penalty coefficient can over‑regularize the critic; in such cases we prefer to rely primarily on a single Lipschitz control.